### v2

# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

Abstract.
We extend Ratner’s theorem on equidistribution of individual orbits of unipotent flows on finite volume homogeneous spaces of Lie groups to trajectories of non-contracting curves definable in a polynomially bounded o-minimal structure.
To be precise, let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous curve whose coordinate functions are definable in a polynomially bounded o-minimal structure; for example, rational functions. Suppose that $\varphi$ is non-contracting; that is, for any linearly independent vectors $v_{1},\ldots,v_{k}$ in $\mathbb{R}^{n}$, $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\not\to 0$ as $t\to\infty$. Then, there exists a unique closed connected subgroup $H_{\varphi}\subseteq\text{SL}(n,\mathbb{R})$ of smallest dimension such that $H_{\varphi}$ is generated by unipotent one-parameter subgroups and such that $\varphi(t)H_{\varphi}\to g_{0}H_{\varphi}$ in $\text{SL}(n,\mathbb{R})/H_{\varphi}$ as $t\to\infty$, for some $g_{0}\in\text{SL}(n,\mathbb{R})$.
Let $\mathcal{G}\subseteq\text{SL}(n,\mathbb{R})$ be a closed subgroup and let $\Gamma\subseteq\mathcal{G}$ be a lattice. Suppose that $\varphi([0,\infty))\subseteq\mathcal{G}$. Then $H_{\varphi}\subseteq\mathcal{G}$, and for any $x\in\mathcal{G}/\Gamma$, the trajectory $\{\varphi(t)x:t\in[0,T]\}$ is equidistributed with respect to the measure $g_{0}\mu_{Lx}$ as $T\to\infty$, where $L\subseteq\mathcal{G}$ is a closed subgroup such that $\overline{H_{\varphi}x}=Lx$ and $Lx$ admits a unique $L$-invariant probability measure, denoted by $\mu_{Lx}$.
A crucial new ingredient is the proof that for any finite-dimensional representation $V$ of $\text{SL}(n,\mathbb{R})$, there exist $T_{0}>0$, $C>0$, and $\alpha>0$ such that for any $v\in V$, the map $t\mapsto\lVert\varphi(t)v\rVert$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
1. Introduction
[…]
More generally, for a continuous curve $\varphi:[0,\infty)\to\mathcal{G}$, $x\in X$, and $T>0$, consider the probability measure $\mu_{T,\varphi,x}$ on $X$ defined by:
[…]
In [Shah1994LimitDO], it was shown that the equidistribution of unipotent flows due to Ratner generalizes to polynomial curves. Namely, if $\mathcal{G}\subseteq\text{SL}(n,\mathbb{R})$ and $\varphi:\mathbb{R}\to\mathcal{G}$ is a map whose coordinate functions are polynonmials, then the measures $\mu_{T,\varphi,x}$ converge to $\varphi(0)\mu_{Hx}$ as $T\to\infty$, where $H\subseteq\mathcal{G}$ is the smallest closed subgroup such that $\{\varphi(0)^{-1}\varphi(t):t\in\mathbb{R}\}\subseteq H$ and the orbit $Hx$ is periodic.
More recently, Peterzil and Starchenko [Peterzil2018OminimalFO] studied such questions in the following setting: Let $\mathcal{G}$ be a closed subgroup of the group of the upper-triangular unipotent matrices in $\text{SL}(n,\mathbb{R})$ and let $\Gamma\subseteq\mathcal{G}$ be a (cocompact) lattice. They showed that if $S\subseteq\mathcal{G}$ is definable in an o-minimal structure, then there exists a definable set $S^{\prime}\subseteq\mathcal{G}$ whose image in $\mathcal{G}/\Gamma$ is the closure of the image of $S$ in $\mathcal{G}/\Gamma$. Moreover, they proved the following [Peterzil2018OminimalFO, Theorem 1.6]: Let $\varphi:[0,\infty)\to\mathcal{G}$ be a curve definable in a polynomially bounded o-minimal structure, and let $x\in\mathcal{G}/\Gamma$ such that $\{\varphi(t)x:t\geq 0\}$ is dense in $G/\Gamma$. Then, $\mu_{T,\varphi,x}$ converges to the $\mathcal{G}$-invariant probability on $\mathcal{G}/\Gamma$ as $T\to\infty$.
Our goal in this paper is to generalize the above equidistribution results for curves definable in polynomially bounded o-minimal structures taking values in linear Lie groups, under the assumption that an additional non-contraction condition is satisfied.
1.1. O-minimal curves
We now provide an overview of the notion of an o-minimal structure. In this paper, o-minimal structures refer to o-minimal structures on the real field. We have provided a more detailed summary of the required definitions, results, and references in Appendix B.
The notion of o-minimality was introduced in [generalization_tarski_seind_dries] as a generalization of semi-algebraic geometry. The semi-algebraic sets in $\mathbb{R}^{d}$ are formed by taking finite unions and intersections of algebraic sets $\{x\in\mathbb{R}^{d}:P(x)=0\}$ and of sets of the form $\{x\in\mathbb{R}^{d}:P(x)>0\}$, where $P$ is a polynomial in $d$ variables. These semi-algebraic sets form the smallest o-minimal structure, and our results are already new in the semi-algebraic setup.
More generally, an o-minimal structure is composed of sets in $\mathbb{R}^{d}$ for each $d\in\mathbb{N}$ such that several axioms are satisfied. The collection of semi-algebraic sets satisfy those axioms and form the smallest o-minimal structure. In a general o-minimal structure, the collection of sets are referred to as the definable sets, and a function is called a definable function if its graph is a definable subset. The definable sets and functions share many desirable properties, see Appendix B. Particularly, the collection of definable functions is both rich and flexible: for example, in any o-minimal structure, all rational functions are definable; these are functions of the form $f(t)=\frac{P(t)}{Q(t)}$, on the domain where $Q(t)\neq 0$, and where $P$ and $Q$ are polynomials. Moreover, if a definable function is injective, its inverse is also definable, and the composition of definable functions remains definable.
An o-minimal structure is called polynomially bounded, if for every definable function $f:\mathbb{R}\to\mathbb{R}$, there exists $r\in\mathbb{R}$ such that $f(x)=O(x^{r})$, as $x\to\infty$. The semi-algebraic o-minimal structure is polynomially bounded. Larger polynomially bounded o-minimal structures include as definable functions the analytic functions restricted to compact sets, and the power functions $t\mapsto t^{r}$, for $r\in\mathbb{R}$; see [generalization_tarski_seind_dries, Miller_Expansions_of_real_field].
We note the following useful fact.
Proposition 1.1.
[miller_exp_hard_avoid]: If $f:[0,\infty)\to\mathbb{R}$ is definable in a polynomially bounded o-minimal structure, then either $f$ is eventually constantly zero, or there exists $r\in\mathbb{R}$ such that $\lim_{t\to\infty}f(t)/t^{r}=c\neq 0$. In the latter case, we define
(1.2) $\deg(f):=r.$
Remark 1.2.
Our equidistribution results will be stated only for polynomially bounded o-minimal structures. As noted in [Peterzil2018OminimalFO], this setup is optimal in the following sense. Notice that the measures on the circle defined by
[…]
do not converge as $T\to\infty$. By [miller_exp_hard_avoid], if an o-minimal structure is not polynomially bounded, then $x\mapsto e^{x}$ is definable, and as a consequence, $x\mapsto\log(x)$ will also be definable.
We will denote by $\mathbf{G}$ an affine algebraic group defined over $\mathbb{R}$, and in what follows, we let $G:=\mathbf{G}(\mathbb{R})$ be the real points of $\mathbf{G}$.
Definition 1.3.
Let $\mathbb{R}[G]$ be the coordinate ring of real polynomial functions on $G$. We say that a curve $\varphi:[0,\infty)\to G$ is definable in an o-minimal structure, if the map $t\mapsto f(\varphi(t))$ on $[0,\infty)$ is definable, for all $f\in\mathbb{R}[G]$.
We observe that if $V$ is a finite-dimensional rational representation, and $\varphi:[0,\infty)\to G$ is definable in $\mathscr{S}$, then the curve $t\mapsto\varphi(t)\cdot v$ is definable in $\mathscr{S}$ for all $v\in V$. So, if $G$ is embedded in $\text{SL}(n,\mathbb{R})\subseteq\mathrm{M}(n,\mathbb{R})$, and
(1.3) $\varphi(t)=[\varphi_{i,j}(t)]_{1\leq i,j\leq n},\ \forall t\in[0,\infty),$
then $\varphi$ is definable in $\mathscr{S}$ if and only if each $\varphi_{i,j}:[0,\infty)\to\mathbb{R}$ is definable in $\mathscr{S}$.
1.1.1. The non-contraction property
We now introduce the non-contraction property, which narrows down the polynomially bounded o-minimal curves to a class of curves whose orbits on rational linear representations of $G$ have the $(C,\alpha)$-good growth property introduced by Kleinbock and Margulis [Klein_Marg_flows_and_dioph].
Definition 1.4.
We say that a definable curve $\varphi:[0,\infty)\to G$ is non-contracting for $G$ if for all rational finite-dimensional representations $V$ of $G$ defined over $\mathbb{R}$, it holds that $\lim_{t\to\infty}\varphi(t)\cdot v\neq 0$ for all $v\in V\setminus\{0\}$.
We note that for any definable curve $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ and any $v\in\mathbb{R}^{n}$, $\lim_{t\to\infty}\varphi(t)\cdot v$ exists in $V\cup\{\infty\}$ by the monotonicity theorem B.5.
Remark 1.5.
Suppose that $G$ is a unipotent group and $\varphi:[0,\infty)\to G$ is a definable curve. Then $\varphi$ is non-contracting. This is so because, for any algebraic action of an algebraic unipotent group on a finite-dimensional vector space, every orbit of a nonzero vector is Zariski closed, and in particular, the closure of the orbit does not contain the origin. See [Birks71, Theorem 12.1].
We now give a practical criterion that allows one to verify the non-contraction property for curves in $\text{SL}(n,\mathbb{R})$, see also Proposition 2.22. Consider the exterior representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
$g\cdot(v_{1}\wedge v_{2}\wedge\cdots\wedge v_{k}):=gv_{1}\wedge gv_{2}\wedge\cdots\wedge gv_{k}.$
Proposition 1.6.
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve definable in an o-minimal structure. Then $\varphi$ is non-contracting if and only if for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$, it holds that
(1.4) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\neq 0.$
1.2. The main equidistribution result
To state our main result, we need to introduce an important subgroup associated with our curve. Let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve definable in an o-minimal structure. In Theorem 2.5, we will show that there exists a smallest closed connected subgroup $H_{\varphi}\subseteq G$ generated by unipotent one-parameter subgroups such that the image of $\varphi([0,\infty))$ in $G/H_{\varphi}$ is bounded. The subgroup $H_{\varphi}$ will be called the hull of $\varphi$. We also establish the existence of a bounded definable curve $\beta:[0,\infty)\to G$, which will be called a correcting curve, such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t$. Since $\beta$ is bounded and definable, $\lim_{t\to\infty}\beta(t)=\beta_{\infty}\in G$ by the monotonicity theorem, see Theorem B.5.
Now suppose that $\varphi$ takes its values in a connected unimodular Lie subgroup $\mathcal{G}\subseteq G$ and the o-minimal structure is polynomially bounded. Then $H_{\varphi}\subseteq\mathcal{G}$, see Proposition 5.13, and hence $\beta$ takes its values in $\mathcal{G}$. Suppose that $\Gamma$ is a lattice in $\mathcal{G}$. Then, by Ratner’s theorem [Ratner91a], since $H_{\varphi}$ is generated by one-parameter unipotent subgroups, for any $x_{0}\in\mathcal{G}/\Gamma$ there exists a closed connected subgroup $L\subseteq\mathcal{G}$ such that $Lx_{0}$ is a periodic orbit, and $\overline{H_{\varphi}x_{0}}=Lx_{0}$.
Theorem 1.7.
Let $\mathcal{G}\subseteq G$ be a connected Lie subgroup and let $\Gamma\subseteq\mathcal{G}$ be a lattice. Let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve definable in a polynomially bounded o-minimal structure such that $\varphi\left([0,\infty)\right)\subseteq\mathcal{G}$. Let $x_{0}\in\mathcal{G}/\Gamma$. Then,
$\lim_{T\to\infty}\mu_{T,\varphi,x_{0}}=\beta_{\infty}^{-1}\mu_{Lx_{0}},$
in the weak-$\ast$ topology, where $Lx_{0}=\overline{H_{\varphi}x_{0}}$ is periodic, and $\mu_{Lx_{0}}$ is the $L$-invariant probability measure on $Lx_{0}$.
As mentioned above, Theorem 1.7 generalizes Shah’s result on polynomial curves [Shah1994LimitDO], and generalizes the result of Peterzil and Starchenko on dense curves in nilmanifolds [Peterzil2018OminimalFO].
Remark 1.8.
The non-contraction assumption in Theorem 1.7 is essential. Indeed, if $\varphi$ is a contracting curve in $\text{SL}(n,\mathbb{R})$, then, there exists $x_{0}\in\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that $\lim_{t\to\infty}\varphi(t)x_{0}=\infty$. To see this, we note that if $\varphi$ is contracting, then by Proposition 1.6, there exist linearly independent vectors $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$ such that $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})\to 0$ as $t\to\infty$. Let $g_{0}\in\text{SL}(n,\mathbb{R})$ satisfy $v_{1}\wedge\cdots\wedge v_{k}=g_{0}\cdot(e_{1}\wedge\cdots\wedge e_{k})$. Since the orbit $\text{SL}(n,\mathbb{Z})\cdot(e_{1}\wedge\cdots\wedge e_{k})$ is discrete and does not contain 00, given a compact set $C\subseteq\text{SL}(n,\mathbb{R})$, it holds that $\varphi(t)g_{0}\notin C\text{SL}(n,\mathbb{Z})$ for all large $t$.
In the next result, proved in Section 7, we give an example of a definable curve in $\text{SL}(n+1,\mathbb{R})$ whose hull is $\text{SL}(n+1,\mathbb{R})$. As a consequence, every trajectory of such a curve in the space of unimodular lattice in $\mathbb{Z}^{n+1}$ gets equidistributed (Corollary 1.10).
Proposition 1.9.
Let $f_{0},...,f_{n}$, and $h_{1},\ldots,h_{n}$ be continuous real valued functions on $[0,\infty)$ definable in a polynomially bounded o-minimal structure. For $t\geq 0$, define
(1.5) $\varphi(t):=\begin{bmatrix}f_{0}(t)&f_{1}(t)&\cdots&f_{n}(t)\\ &h_{1}(t)^{-1}&&\\ &&\ddots&\\ &&&h_{n}(t)^{-1}\end{bmatrix}$
Suppose that the functions $h_{i}$’s and $f_{i}$’s defining $\varphi$ satisfy the following conditions:
  1. (1)
    $f_{0}=h_{1}\cdots h_{n}$, $\deg f_{0}=n$, and $\deg h_{1}\geq\cdots\geq\deg h_{n}>0$.
  2. (2)
    $\deg(f_{0}+g)\geq n$ for any $g$ in the linear span of $f_{1},\ldots,f_{n}$.
  3. (3)
    For any $i\in\{1,\ldots,n\}$, $\deg(f_{i}+g)>n-i$ for any $g$ in the linear span of $f_{i+1},\ldots,f_{n}$.
Then the hull of $\varphi$ is $H_{\varphi}=\text{SL}(n+1,\mathbb{R})$.
For example, let $f_{0}(t)=t^{n}$. Suppose $f_{1},\dots,f_{n}$ are rational functions such that $\deg f_{1}>\cdots>\deg f_{n}$, and $\deg f_{i}\neq n$ and $\deg f_{i}>n-i$ for all $1\leq i\leq n$. Let $h_{i}(t)=t^{r_{i}}$ for some $r_{1}\geq\cdots\geq r_{n}>0$ such that $\sum_{i=1}^{n}r_{i}=n$. Then, these functions satisfy the conditions of Proposition 1.9.
Using Theorem 1.7, we obtain the following:
Corollary 1.10.
Let $\mathcal{G}$ be a Lie group and $\Gamma$ be a lattice in $\mathcal{G}$. Suppose that $\rho:\text{SL}(n+1,\mathbb{R})\to\mathcal{G}$ is a continuous homomorphism. Let $\varphi:[0,\infty)\to\text{SL}(n+1,\mathbb{R})$ be a curve satisfying the conditions of Proposition 1.9. Then, for any $x\in\mathcal{G}/\Gamma$, and any bounded continuous function $f$ on $\mathcal{G}/\Gamma$,
[…]
where $\mu_{X}$ is the homogeneous probability measure on the homogeneous space $\overline{\rho(\text{SL}(n+1,\mathbb{R}))x}$.
1.3. Growth of o-minimal functions
We now describe the $(C,\alpha)$-good property, introduced by Kleibock and Margulis [Klein_Marg_flows_and_dioph], concerning the growth of certain families of functions definable in polynomially bounded o-minimal structures. This property is well known for polynomials [Dani1993LimitDO, 4.1 Lemma], but is new in the o-minimal setting. It is crucial for the techniques used in this paper.
Definition 1.11.
Let $C$ and $\alpha$ be positive constants. We say that a function $f:[0,\infty)\to\mathbb{R}$ is $(C,\alpha)$-good on $[T_{0},\infty)\subseteq[0,\infty)$, if for every $\epsilon>0$, and for every bounded interval $I\subseteq[T_{0},\infty)$, it holds that:
(1.6) |$|\{t\in I:|f(t)|\leq\epsilon\}|\leq C\left(\frac{\epsilon}{\|f\|_{I}}\right)^{\alpha}\cdot|I|.$
For the following statement, we say that a subset $\mathscr{F}$ of a finite-dimensional vector space $\mathcal{V}$ over $\mathbb{R}$ is a definable cone, if $\mathscr{F}$ is a cone in the vector space, and $\mathscr{F}$ is a definable set when identifying $\mathcal{V}\cong\mathbb{R}^{n}$ by choosing a basis.
Theorem 1.12.
($(C,\alpha)$-good property). Let $\mathcal{V}$ be a finite-dimensional real vector space spanned by functions $f:[0,\infty)\to\mathbb{R}$, each definable in the same polynomially bounded o-minimal structure. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\setminus\{0\}$ it holds that
$\lim_{t\to\infty}f(t)\neq 0.$
Then, there exist $C>0$, $\alpha>0$ and $T_{0}\geq 1$ such that for any $f\in\mathscr{F}\setminus\{0\}$, $f$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
We obtain the following statement as a direct corollary of Theorem 1.12.
Proposition 1.13.
[…]
$\mathcal{V}:=\text{Span}_{\mathbb{R}}\{t\mapsto\|\psi(t)\textbf{v}\|^{2},\ t\geq 0:\textbf{v}\in\mathbb{R}^{m}\}$
is finite dimensional, and :=$\mathscr{F}:=\{t\mapsto\|\psi(t)\textbf{v}\|^{2},\ t\geq 0:\textbf{v}\in\mathbb{R}^{m}\}$ is a closed definable cone.
1.4. Proof ideas and structure of the paper
Our main theorem will be proved by the following strategy employed in many works on equidistribution in homogeneous spaces, for example [Dani1993LimitDO, Mozes1995OnTS, Shah1994LimitDO]:
  1. (1)
    Proving that there is no escape of mass (Section 4) – that is, any limiting measure $\mu$ of the probability measures $\mu_{T,\varphi,x}$ is a probability measure.
  2. (2)
    Showing that any limiting measure $\mu$ of the measures $\mu_{T,\varphi,x}$ is invariant under a group generated by nontrivial unipotent one-parameter subgroups (Section 5).
  3. (3)
    Studying the ergodic decomposition of a limiting measure $\mu$ (Section 6). By Ratner’s theorem [Ratner91a], every ergodic component appearing in the ergodic decomposition of $\mu$ is homogeneous. The theorem is proved once it is verified that $\mu$ is equal to exactly one uniquely determined ergodic component. For this step, we will use the linearization technique of [Dani1993LimitDO] to prove avoidance of singular sets that contain ergodic components of various types.
To establish unipotent invariance, we will use an idea similar to the one used for polynomial curves [Shah1994LimitDO]. This idea generalizes to the o-minimal setting via the notion of the Peterzil-Steinhorn group (see Section 5), which was also used in [Peterzil2018OminimalFO].
The $(C,\alpha)$-good property will be used to prove the non-escape of mass, which is avoidance of the point at infinity, and to apply the linearization technique.
1.5. Notation and conventions
We will denote the Zariski closure of a subset $A$ of $\mathbb{R}^{N}$ by $\text{zcl}(A)$. For a linear Lie group $H$, we denote by $H_{u}$ the subgroup of $H$ generated by all the one-parameter unipotent subgroups contained in $H$. We note that if $H\subseteq\text{GL}(n,\mathbb{R})$, then $H_{u}$ equals the connected component of the identity in $\text{zcl}(H_{u})$.
Acknowledgement.
We thank Chris Miller for helpful discussions. We thank anonymous referees for their careful reading and valuable suggestions which helped us improve the readability of this article.
2. On the non-contracting property and the hull
This section is divided into two main parts. In the first part, we prove Proposition 1.6, which gives a convenient criterion for verifying the non-contraction property. The second part of the section concerns the existence of the hull and the correcting curve, and discusses their behavior under homomorphisms.
2.1. The criterion for the non-contracting property
Our main goal here is to prove Proposition 1.6, namely to reduce the verification of the non-contracting property for any vector in any representation to decomposable vectors in the exterior of the standard representation of $\text{SL}(n,\mathbb{R})$. Our proof follows from the observations in [shah2023equidistribution, Section 2], which uses Kempf’s numerical criteria on geometric invariant theory [Kem78].
We recall the following preliminaries. Let ${S}$ denote the full diagonal subgroup of $\text{SL}(n,\mathbb{R})$. The group of the unipotent upper triangular matrices in $\text{SL}(n,\mathbb{R})$, denoted by $N$, is a maximal unipotent subgroup of $\text{SL}(n,\mathbb{R})$ normalized by ${S}$. Let $X^{\ast}({S})$ denote the group of algebraic homomorphisms from ${S}$ to $\mathbb{R}^{\ast}$ defined over $\mathbb{R}$. Then $X^{\ast}({S})$ is a free abelian group with $(n-1)$ generators, and we treat it as an additive group. For each $i\in\{1,\ldots,n-1\}$, let $\alpha_{i}\in X^{\ast}({S})$ be defined by $\alpha_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{i}/t_{i+1}$. Then $\Delta=\{\alpha_{i}:1\leq i\leq n-1\}$ is the set of simple roots on ${S}$ corresponding to the choice of the maximal unipotent subgroup $N$. For each $11\leq i\leq n-$ 1, let $\mu_{i}\in X^{\ast}({S})$ be defined by $\mu_{i}(\text{diag}(t_{1},\ldots,t_{n}))=t_{1}\cdots t_{i}$. Then $\{\mu_{i}:1\leq i\leq n-1\}$ is the set of fundamental characters (weights) of ${S}$ with respect to our choice of simple roots. For each $11\leq i\leq n-$ 1, let $W_{i}$ denote the $i$-th exterior of the standard representation of $\text{SL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$, called a fundamental representation of $\text{SL}(n,\mathbb{R})$. So $W_{i}=\wedge^{i}\mathbb{R}^{n}$, and for any $v_{1}\wedge\cdots\wedge v_{i}\in W_{i}$ and $g\in\text{SL}(n,\mathbb{R})$, $g\cdot(v_{1}\wedge\cdots\wedge v_{i})=gv_{1}\wedge\cdots\wedge gv_{i}$. Let $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$. Then $w_{i}$ is fixed by $N$, and ${S}$ acts on the line $\mathbb{R}w_{i}$ via the fundamental character $\mu_{i}$.
A dominant integral character (weight) is a non-negative integral combination of fundamental characters. Suppose $\chi=m_{1}\mu_{1}+\cdots+m_{n-1}\mu_{n-1}$, where $m_{i}\in\mathbb{Z}_{\geq 0}$. For each $i$, let $W_{i}^{\otimes m_{i}}$ denote the tensor product of $m_{i}$-copies of $W_{i}$, and $w_{i}^{m_{i}}:=w_{i}\otimes\cdots\otimes w_{i}\in W_{i}^{\otimes m_{i}}$. Consider the tensor product representation of $\text{SL}(n,\mathbb{R})$ on $W_{\chi}=W_{1}^{\otimes m_{1}}\otimes\cdots\otimes W_{n-1}^{\otimes m_{n-1}}$, and let $w_{\chi}=w_{1}^{m_{1}}\otimes\cdots\otimes w_{n-1}^{m_{n-1}}$. Then $w_{\chi}$ is fixed by $N$, and ${S}$ acts on the line $\mathbb{R}w_{\chi}$ via the character $\chi$. We endow $W_{\chi}$ with a $\text{SO}(n,\mathbb{R})$-invariant norm.
Proof of Proposition 1.6.
Let $V$ be a finite dimensional representation of $G:=\text{SL}(n,\mathbb{R})$ endowed with some norm. Suppose $v\in V$ is a nonzero vector such that $\varphi(t)\cdot v\to 0$ as $t\to\infty$. Then $\overline{G\cdot v}\ni 0$, so $v$ is a $G$-unstable vector in $V$. So, by [shah2023equidistribution, Remark 2.3], there exist $g_{0}\in G$, a dominant integral character $\chi\in X^{\ast}({S})$, and constants $\beta>0$ and $C>0$ such that for any $g\in G$,
(2.1) $\lVert gg_{0}\cdot w_{\chi}\rVert\leq C\lVert g\cdot v\rVert^{\beta}.$
Therefore, $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{\chi}=0$. Let $A={S}^{0}$, the component of the identity in ${S}$. Let $K=\text{SO}(n,\mathbb{R})$. Then by Iwasawa decomposition, $G=KAN$. For each $t$, we express $\varphi(t)g_{0}=k_{t}a_{t}n_{t}$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$. Since the norm on $W_{\chi}$ is $K$-invariant and $N$ fixes $w_{\chi}$,
$\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert=\lVert a_{t}\cdot w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
By (2.1), $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{\chi}=0$. Therefore, $\lim_{t\to\infty}\chi(a_{t})=0$. Now
$\chi(a_{t})=\mu_{1}(a_{t})^{m_{1}}\cdots\mu_{n-1}(a_{t})^{m_{n-1}}.$
Since $m_{i}\geq 0$, after passing to a subsequence, we can pick $i\in\{1,\ldots,n-1\}$ and a sequence $t_{l}\to\infty$ such that $\lim_{l\to\infty}\mu_{i}(a_{t_{l}})=0$. Therefore, since $N$ fixes $w_{i}$, and the norm is $K$-invariant,
$\lVert\varphi(t_{l})g_{0}\cdot w_{i}\rVert=\lVert{a_{t_{l}}\cdot w_{i}}\rVert=\mu_{i}(a_{t_{l}})\lVert w_{i}\rVert\to 0\text{, as $l\to\infty$.}$
Since the coordinate functions of $t\to\varphi(t)g_{0}\cdot w_{i}$ are definable in an o-minimal structure, we get $\lim_{t\to\infty}\varphi(t)g_{0}\cdot w_{i}=0$. Let $v_{j}=g_{0}e_{j}$ for each $j$. Since $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$, $\varphi(t)\cdot(g_{0}\cdot w_{i})=\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{i})\to 0$ as $t\to\infty$. This contradicts the assumption that the curve is non-contracting. ∎
We record the following statement. It will be used in the proof of Lemma 7.1, which is in turn needed for proving Proposition 1.9.
Lemma 2.1.
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a map such that
(2.2) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and linearly independent vectors $v_{1},...,v_{k}\in\mathbb{R}^{n}$. Then for any finite-dimensional rational representation $V$ and any $v\in V$, if $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$, then $\text{SL}(n,\mathbb{R})\cdot v$ is Zariski closed.
Proof.
Suppose that $V$ is a finite-dimensional rational representation of $G=\text{SL}(n,\mathbb{R})$ and let $v\in V$ such that $G\cdot v$ is not Zariski closed. So there exists a finite-dimensional rational representation $W$ of $G$ and a $G$-equivariant polynomial map $P:V\to W$ such that $P(v)\neq 0$ and $\overline{G\cdot P(v)}\ni 0$, see Lemma A.1. Therefore, by [shah2023equidistribution, Remark 2.3], there exist a $g_{0}\in G$, a nontrivial dominant integral character $\chi\in X^{\ast}({S})$, and constants $\beta>0$ and $C>0$ such that for any $g\in G$,
$\lVert gg_{0}\cdot w_{\chi}\rVert\leq C\lVert g\cdot P(v)\rVert^{\beta}=C\lVert P(g\cdot v)\rVert^{\beta},$
cf. [shah2023equidistribution, Corollary 2.5]. Since $P$ is continuous, given any $R>0$, there exists $C_{R}>0$ such that for any $g\in G$,
(2.3) if gvR, then $\text{if $\|g\cdot v\|\leq R$, then }\lVert gg_{0}\cdot w_{\chi}\rVert\leq C_{R}\lVert g\cdot v\rVert^{\beta}.$
Now suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a map satisfying the diverging property (2.2) and assume for contradiction that $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$. Then, by (2.3), we get that $\{\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert:t>0\}$ is bounded. As in the proof of Proposition 1.6, by Iwasawa decomposition, for all $t>0$, $\varphi(t)g_{0}=k_{t}a_{t}n_{t}$, where $k_{t}\in K$, $a_{t}\in A$ and $n_{t}\in N$, and
$\lVert\varphi(t)g_{0}\cdot w_{\chi}\rVert=\lVert a_{t}\cdot w_{\chi}\rVert=\chi(a_{t})\lVert w_{\chi}\rVert.$
We claim that $\lim_{t\to\infty}\chi(a_{t})=\infty,$ which is a contradiction. Indeed, we have
$\chi(a_{t})=\mu_{1}(a_{t})^{m_{1}}\cdots\mu_{n-1}(a_{t})^{m_{n-1}},$
for some $(m_{1},\ldots,m_{n-1})\in\mathbb{Z}_{\geq 0}$ and $m_{i}>0$ for some $i$. Now
$\lVert\varphi(t)g_{0}\cdot w_{i}\rVert=\lVert{a_{t}\cdot w_{i}}\rVert=\mu_{i}(a_{t})\lVert w_{i}\rVert,$
where $w_{i}=e_{1}\wedge\cdots\wedge e_{i}$ and $11\leq i\leq n-$ 1. And by (2.2), we have $\lim_{t\to\infty}\lVert\varphi(t)g_{0}\cdot w_{i}\rVert=\infty$. So, $\lim_{t\to\infty}\mu_{i}(a_{t})\to\infty$. ∎
2.2. On the hull and the correcting curve
In what follows, $G:=\mathbf{G}(\mathbb{R})$ is the real points of an algebraic group $\mathbf{G}$ over $\mathbb{R}$. To define the hull and obtain its useful properties, we will need the following notion.
Definition 2.2.
We say that a Zariski closed subgroup $H$ is observable in $G$ if there is a finite-dimensional rational representation $V$ over $\mathbb{R}$ and $v\in V$ such that $H=\{g\in G:g\cdot v=v\}$.
The notion of observable groups was introduced [Mostow_et_al_observable_groups]. More precisely, [Mostow_et_al_observable_groups] defines observable groups as the algebraic subgroups of algebraic groups for which every finite-dimensional rational representation extends to a finite-dimensional rational representation of the ambient group. The above definition turns out to be equivalent to the extension property [Mostow_et_al_observable_groups, Theorem 8] (see [Observable_groups_over_fields, Theorem 9] for this statement for non-algebraically closed fields). The following known result will be useful, see e.g. [Observable_group_grosshans, Corollary 2.8].
Lemma 2.3.
Let $H$ be an algebraic subgroup of an algebraic group $G$ over $\mathbb{R}$. If the radical of $H$ is unipotent, then $H$ is an observable subgroup of $G$.
Remark 2.4.
Suppose that $L$ is a closed subgroup of $\text{SL}(n,\mathbb{R})$ generated by unipotent one-parameter subgroups. Then $L=\text{zcl}(L)^{0}$ and the radical of $\text{zcl}(L)$ is unipotent (see [Shah1991, Lemma 2.9]). Hence, by Lemma 2.3, $\text{zcl}(L)$ is observable in $\text{SL}(n,\mathbb{R})$.
In what follows, we say that a curve $\varphi([0,\infty))\subseteq G$ is bounded modulo a closed subgroup $H\subseteq G$ if there exists a compact subset $K\subseteq G$ such that $\varphi([0,\infty))\subseteq KH$.
We fix an o-minimal structure $\mathscr{S}$ for the rest of the following two subsections. We stress that the assumption of polynomial boundedness on $\mathscr{S}$ is unnecessary here.
Here is our main theorem in this section.
Theorem 2.5.
Let $\varphi:[0,\infty)\to G$ be a continuous, non-contracting curve definable in $\mathscr{S}$. Then, there exists a unique closed connected (in the Hausdorff topology) subgroup $H_{\varphi}\subseteq G$ of the smallest dimension such that $H_{\varphi}$ is generated by one-parameter unipotent subgroup and such that $\varphi$ is bounded modulo $H_{\varphi}$. Moreover, the following properties hold:
  1. (1)
    The subgroup $H_{\varphi}$ is minimal in the following sense: if $H^{\prime}$ is an observable group such that $\varphi$ is bounded modulo $H^{\prime}$, then $H_{\varphi}\subseteq H^{\prime}$. Moreover, if $V$ is a rational representation, and $v\in V$ is such that $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$, then $\varphi$ is bounded modulo the isotropy group of $v$ and $H_{\varphi}\cdot v=\{v\}$.
  2. (2)
    There is a bounded, continuous curve $\beta:[0,\infty)\to G$ definable in $\mathscr{S}$ such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$.
  3. (3)
    If $V$ is a rational finite-dimensional representation of $G$ over $\mathbb{R}$, then for any $v\in V$, either $H_{\varphi}\cdot v=\{v\}$ or $\lim_{t\to\infty}\beta(t)\varphi(t)\cdot v=\infty$.
In view of the above statement, we can now define the hull.
Definition 2.6.
Let $\varphi:[0,\infty)\to G$ be a continuous non-contracting curve definable in $\mathscr{S}$. We define the hull $H_{\varphi}$ to be the smallest connected closed subgroup generated by unipotent elements such that $\varphi$ is bounded modulo $H_{\varphi}$, and we call a curve $\beta:[0,\infty)\to G$ a correcting curve if $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$.
Remark 2.7.
In view of Remark 2.4, $H_{\varphi}=\text{zcl}(H_{\varphi})^{0}$ and $\text{zcl}(H_{\varphi})$ is the observable subgroup of smallest dimension modulo which $\varphi$ is bounded.
Remark 2.8.
We note that if $\varphi$ is as in the above theorem and $\beta$ is a definable correcting curve, then the hull of $\psi(t):=\beta(t)\varphi(t)$ is the same as the hull of $\varphi$, and the correcting curve of $\psi$ can be chosen to be the constant identity matrix. This follows from the uniqueness of the hull.
Remark 2.9.
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a definable curve whose entry functions $\varphi_{i,j}$ generate an algebra which consists of functions that are either constant or diverging to $\infty$, e.g., when $t\mapsto\varphi_{i,j}(t)$ are polynomials for all $i,j$. Then, by Theorem 2.5, $\varphi(0)^{-1}\varphi([0,\infty))\subseteq H_{\varphi}$. A correcting curve is given by the constant curve $\beta(t)=\varphi(0)^{-1}$ for all $t\geq 0$.
Before proving Theorem 2.5, we establish that the hull is intrinsically defined.
Lemma 2.10.
Let $G_{1}$ and $G_{2}$ be the real points of algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a rational homomorphism. Suppose that $\varphi:[0,\infty)\to G_{1}$ is a non-contracting curve definable in $\mathscr{S}$. Then $f(H_{\varphi})=H_{f\circ\varphi}$.
Proof.
We start by showing that $H_{f\circ\varphi}\subseteq f(H_{\varphi})$. Let $\beta:[0,\infty)\to G_{1}$ be a correcting curve of $\varphi$. Since $H_{\varphi}$ is generated by unipotents, we get that $f(H_{\varphi})$ is generated by unipotents. Moreover, the curve $f\circ\varphi$ is bounded modulo $f(H_{\varphi})$. In fact, $t\mapsto f(\beta(t))$ is bounded, and for all $t\geq 0$:
$f(\beta(t)\varphi(t))=f(\beta(t))f(\varphi(t))\in f(H_{\varphi}).$
Thus, by minimality of the hull, we conclude that $H_{f\circ\varphi}\subseteq f(H_{\varphi})$.
We proceed to show the other inclusion. Since $\text{zcl}(H_{f\circ\varphi})$ is observable, we choose a rational representation $\rho:G_{2}\to\text{GL}(V)$ and a $v\in V$ such that
$\text{zcl}(H_{f\circ\varphi})=\{g\in G_{2}:\rho(g)v=v\}.$
Since $f\circ\varphi$ is bounded modulo $H_{f\circ\varphi}$, the trajectory,
$\rho((f\circ\varphi)(t))v=(\rho\circ f)(\varphi(t))v,~{}t\geq 0,$
is bounded in $V$. Now, $\eta:=\rho\circ f:G_{1}\to\text{GL}(V)$ is a representation of $G_{1}$ such that $t\mapsto\eta(\varphi(t))v,~{}t\geq 0$ is bounded. By Theorem 2.5(1), we obtain that $\rho(f(H_{\varphi}))v=\eta(H_{\varphi})v=\{v\}$. That is, $f(H_{\varphi})$ is contained in the isotropy group of $v$ which is equal to $\text{zcl}(H_{f\circ\varphi})$. Since $f(H_{\varphi})$ is generated by unipotent one-parameter subgroups,
$f(H_{\varphi})\subseteq\text{zcl}(H_{f\circ\varphi})^{0}=H_{f\circ\varphi}.$
2.3. Proving Theorem 2.5
We begin with a “curve correcting” lemma.
Lemma 2.11.
Consider an algebraic action of $G$ on the $\mathbb{R}$ points of an affine or projective variety $\mathcal{Z}$ over $\mathbb{R}$. Let $\varphi:[0,\infty)\to G$ be a continuous curve definable in $\mathscr{S}$. Suppose that $x\in\mathcal{Z}$ and $g_{0}\in G$ are such that
$\lim_{t\to\infty}\varphi(t)\cdot x=g_{0}\cdot x.$
Then, there exists a continuous $\mathscr{S}$-definable curve $\delta:[0,\infty)\to G$ such that
$\lim_{t\to\infty}\delta(t)=e\text{ and }\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\ \forall t\geq 0.$
Proof.
Identify $G$ as a closed subgroup of $\text{SL}(n,\mathbb{R})$. Consider the following definable set:
$\mathcal{D}$ :=$:=\{(t,g):t>0,\ g\in G,\|g-I_{n}\|<1\text{ and }g_{0}^{-1}\varphi(t)\cdot x=g\cdot x\}.$
[…]
$\lim_{t\to\infty}g_{0}^{-1}\varphi(t)\cdot x=x,$
we get that for all $t$ large enough, say $t\geq T_{0}$, there exists $g\in G$ such that $(t,g)\in\mathcal{D}$. Using the choice function theorem, see Theorem B.7, there is a definable curve $\tilde{\delta}(t):[T_{0},\infty)\to G$ such that $(t,\tilde{\delta}(t))\in\mathcal{D}$ for all $t\geq T_{0}$. In particular,
$\tilde{\delta}(t)\cdot x=g_{0}^{-1}\varphi(t)\cdot x,\,\forall t\geq T_{0},$
and $\tilde{\delta}$ is bounded. Since the curve $\tilde{\delta}(t)$ is bounded, by the monotonicity theorem, see Theorem B.5, we have that $\lim_{t\to\infty}\tilde{\delta}(t)=\delta_{0}$. Note that $\delta_{0}\cdot x=x$. Let $\delta(t):=g_{0}(\tilde{\delta}(t)\delta_{0}^{-1})^{-1}g_{0}^{-1}$ for all $t\geq T_{0}$. Then,
$\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\forall t\geq T_{0}.$
By the monotonicity theorem, $\delta(t)$ is continuous in $[T^{\prime}_{0},\infty)$, for some $T^{\prime}_{0}\geq T_{0}$. We now extend $\delta$ to the interval $[0,T^{\prime}_{0}]$ by putting
$\delta(t):=\delta(T^{\prime}_{0})\varphi(T^{\prime}_{0})\varphi(t)^{-1},\,\forall t\leq T^{\prime}_{0}.$
Then $\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\forall t\geq 0$. Since $\varphi$ is continuous and definable, $\delta:[0,\infty)\to G$ is continuous, definable and bounded. ∎
Lemma 2.12.
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Suppose that $V$ is a finite-dimensional rational representation of $G$ and $v\in V\setminus\{0\}$ is such that $\varphi(t)\cdot v$ is bounded. Then, there exists $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t)\cdot v=g_{0}\cdot v.$
In particular, there exists a bounded, continuous curve $\beta:[0,\infty)\to G$ definable in $\mathscr{S}$, such that $\beta(t)\varphi(t)\cdot v=v$ for all $t\geq 0$.
Proof.
Since $\varphi:[0,\infty)\to G$ is definable, and since $V$ is a rational representation, $\varphi(t)\cdot v$ is a bounded definable curve. By the monotonicity theorem (see Theorem B.5), $\lim_{t\to\infty}\varphi(t)\cdot v=v^{\prime}\in V$. We claim that $v^{\prime}\in G\cdot v$. Denote $S:=\text{zcl}(G\cdot v)\setminus G\cdot v$, and suppose for contradiction that $v^{\prime}\in S$. Then, by Lemma A.1, which is a slight modification of [Kem78, Lemma 1.1(b)], there exists a rational representation $W$ of $G$ and $G$-equivariant polynomial map $P:V\to W$ such that $P(v)\neq 0$ and $P(S)=0$. In particular, $P(v^{\prime})=0$. Then
$\lim_{t\to\infty}\varphi(t)\cdot P(v)=P(\lim_{t\to\infty}\varphi(t)\cdot v)=P(v^{\prime})=0,$
which is a contradiction, as $\varphi$ is non-contracting. Thus, we conclude that there exists $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t)\cdot v=g_{0}\cdot v.$
The existence of the bounded curve $\beta$ is obtained by Lemma 2.11. ∎
Lemma 2.13.
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Let $H_{1}$ and $H_{2}$ be observable groups of $G$ such that the image of $\varphi$ is bounded in $G/H_{i}$ for $i\in\{1,2\}$. Then, the image of $\varphi$ is bounded in $G/(H_{1}\cap H_{2})$.
Proof.
For $i\in\{1,2\}$, let $V_{i}$ be a finite-dimensional rational representation such that $H_{i}=\{g\in G:g\cdot v_{i}=v_{i}\}$ where $v_{i}\in V_{i}$. Consider the direct-sum representation $V_{1}\oplus V_{2}$, and notice that
$H_{1}\cap H_{2}=\{g\in G:g\cdot(v_{1},v_{2})=(v_{1},v_{2})\}.$
Since the image of $\varphi$ is bounded in $G/H_{i}$, $\varphi(t)\cdot v_{i}$ is bounded for all $i\in\{1,2\}$. In particular, $\varphi(t)\cdot(v_{1},v_{2})$ is bounded in $V_{1}\oplus V_{2}$. Then, by Lemma 2.12, there is a bounded continuous curve $\delta:[0,\infty)\to G$ such that $\delta(t)\varphi(t)\cdot(v_{1},v_{2})=(v_{1},v_{2})$ for all $t\geq 0$. Namely, $\delta(t)\varphi(t)\in H_{1}\cap H_{2}$ for all $t\geq 0$, and therefore the image of $\varphi$ is bounded in $G/(H_{1}\cap H_{2})$. ∎
Corollary 2.14.
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve definable in $\mathscr{S}$. Then there exists a unique observable subgroup $\tilde{H}_{\varphi}\subseteq G$ of minimal dimension such that $\varphi$ is bounded in $G/\tilde{H}_{\varphi}$. Moreover, if $H$ is an observable subgroup such that $\varphi$ is bounded in $G/H$, then $\tilde{H}_{\varphi}\subseteq H$.
Proof.
Let $H_{0}$ be a Zariski connected observable group of the smallest dimension such that $\varphi$ is bounded in $G/H_{0}$. If $H$ is any connected observable group such that $\varphi$ is bounded in $G/H$, then by Lemma 2.13, we get that $\varphi$ is bounded in $G/(H_{0}\cap H)$. If $H_{0}\not\subseteq H$, then $H\cap H_{0}$ is a proper subgroup of dimension strictly smaller than $\dim H_{0}$, which leads to a contradiction. ∎
Corollary 2.15.
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Let $\tilde{H}_{\varphi}\subseteq G$ be the minimal observable group such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Denote $H_{\varphi}:=\tilde{H}_{\varphi}^{0}$, the identity component of $\tilde{H}_{\varphi}$.
Then, there exists a bounded, continuous definable curve $\beta:[T_{0},\infty)\to G$ such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$. Moreover, if $V$ is a finite-dimensional rational representation of $G$ and $v\in V$ is such that $\varphi(t)\cdot v$ is bounded, then $\beta(t)\varphi(t)\cdot v=v$ for all $t\geq T_{0}$.
Proof.
By Lemma 2.12, we obtain a continuous definable curve $\tilde{\beta}:[0,\infty)\to G$ such that $\tilde{\beta}(t)\varphi(t)\in\tilde{H}_{\varphi}$ for all $t\geq 0$. Since $\varphi$ is continuous, it is contained in a connected component of $\tilde{H}_{\varphi}$. By choosing a suitable $h_{0}\in\tilde{H}_{\varphi}$ we get that for $\beta(t):=h_{0}\tilde{\beta}(t)$ it holds $\beta(t)\varphi(t)\in\tilde{H}^{0}_{\varphi}=H_{\varphi}$, for all $t\geq 0$. Now, if $V$ is a finite-dimensional rational representation such that $\varphi(t)\cdot v$ is bounded, then by Lemma 2.12, we conclude that $\varphi$ is bounded modulo the observable group $H:=\{g\in G:g\cdot v=v\}$. Hence, by minimality of $\tilde{H}_{\varphi}$, we get that $\beta(t)\varphi(t)\in H_{\varphi}\subseteq H$. ∎
The last property in Theorem 2.5 remaining to prove is that the identity component of the observable group of smallest dimension is generated by one-parameter unipotent subgroups. This is given by the following.
Lemma 2.16.
Let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}$-definable, non-contracting, continuous curve, and let $\tilde{H}_{\varphi}$ be the smallest observable subgroup such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Then $H_{\varphi}=\tilde{H}_{\varphi}^{0}$ is generated by one-parameter unipotent subgroups.
Proof.
Let $\psi(t):=\beta(t)\varphi(t)$ for all $t\geq 0$, where $\beta$ is a correcting curve of $\varphi$. Let $L\subseteq\tilde{H}_{\varphi}$ be the Zariski closure of the subgroup generated by all one-parameter unipotent subgroups of $\tilde{H}_{\varphi}$. Then, $L$ is observable in $G$, see Remark 2.4. We will now show that the image of $\varphi$ in $\tilde{H}_{\varphi}/L$ is bounded, which implies that $\tilde{H}_{\varphi}=L$, and proves the claim, see Remark 2.7.
Since $L$ is a normal subgroup of $\tilde{H}_{\varphi}$, we get that $\tilde{H}_{\varphi}/L$ is an affine algebraic group. Moreover, since $L$ contains the unipotent radical of $\tilde{H}_{\varphi}$, we get that $\tilde{H}_{\varphi}/L$ is reductive. Then, there is an almost direct product decomposition $\tilde{H}_{\varphi}/L=ZD$, where $Z$ is the center, which is a torus, and $D$ is the derived group, which is a normal semi-simple subgroup. Notice that each simple factor of $D$ must be compact since a non-compact simple group is generated by unipotent one-parameter subgroups. Namely, $D$ is compact. Let $q:\tilde{H}_{\varphi}\to\tilde{H}_{\varphi}/L$ be the natural quotient map. To finish the proof, we show below that for every character $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ it holds that $\chi\circ q(\psi(t))=1$ for all $t\geq 0$. This will complete the proof since then $q\circ\psi([0,\infty))\subseteq Z_{a}D$, where $Z_{a}\subseteq Z$ is the maximal anisotropic compact torus subgroup of $Z$.
Let $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ be a character, and consider the representation of $\tilde{H}_{\varphi}$ on $\mathbb{R}^{2}$ defined by
(2.4) $h\cdot(x,y):=\left(\chi(q(h))x,\chi(q(h))^{-1}y\right),\,\forall(x,y)\in\mathbb{R}^{2},\,\forall h\in\tilde{H}_{\varphi}.$
Since $\tilde{H}_{\varphi}\subseteq G$ is observable, the above representation extends to a representation of $G$, see [Observable_groups_over_fields, Theorem 9]. Now, pick $x_{0},y_{0}\in\mathbb{R}\setminus\{0\}$. Because $\psi$ is non-contracting, we get that $\psi(t)\cdot(x_{0},0)$ and $\psi(t)\cdot(0,y_{0})$ are bounded away from 00. In view of (2.4), we deduce that $\psi(t)\cdot(x_{0},y_{0})$ is bounded. So, by Corollary 2.15, $\psi(t)\cdot(x_{0},y_{0})=(x_{0},y_{0})$. Thus, $\chi\circ q(\psi(t))=1$ for all $t\geq 0$. ∎
2.4. The non-contracting property under homomorphisms
In this subsection, we show that the homomorphic image of a definable non/contracting curve is non/contracting, Lemma 2.17, and also we show that if a definable curve in a homomorphic image is non/contracting, then it has a definable non/contracting lift, Lemma 2.21. Finally, we establish that the non/contracting property is intrinsic for curves in groups whose radical is unipotent, Proposition 2.22.
Lemma 2.17.
Let $G_{1}$ and $G_{2}$ be the real points of affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{1}$ be a continuous, non-contracting curve in $G_{1}$ definable in $\mathscr{S}$. Then $f\circ\varphi$ is a continuous, non-contracting curve in $G_{2}$ definable in $\mathscr{S}$.
Proof.
Let $\rho:G_{2}\to\text{GL}(V)$ be a rational representation of $G_{2}$. Let $v\in V$. Then $\rho\circ f:G_{1}\to\text{GL}(V)$ is a rational representation of $G_{1}$. Let $v\in V$. If $\rho(f\circ\varphi(t))v\to 0$ as $t\to\infty$, then $(\rho\circ f)(\varphi(t))v\to 0$ as $t\to\infty$. Since $\varphi$ is non-contracting, we get $v=0$. ∎
Lemma 2.18.
Suppose that $U\subseteq G$ is a normal, unipotent algebraic subgroup. Let $q:G\to G/U$ be the natural map. Then, a $\mathscr{S}$-definable curve $\varphi:[0,\infty)\to G$ is non-contracting in $G$ if and only if $q\circ\varphi$ is non-contracting in $G/U$.
Proof.
By Lemma 2.17, it suffices to prove that if $q\circ\varphi$ is non-contracting, then $\varphi$ is non-contracting. Suppose for contradiction that $q\circ\varphi$ is non-contracting and $\varphi$ is contracting. Let $V$ be a rational representation of $G$ of the smallest dimension such that there exists a $v\in V\setminus\{0\}$ with $\lim_{t\to\infty}\varphi(t)\cdot v=0$. Since $U$ is a unipotent group, the subspace of $U$-fixed points $V^{U}:=\{x\in V:u\cdot x=x,\forall u\in U\}$ is of positive dimension. Since $U$ is a normal subgroup, $V^{U}$ is $G$-invariant. Then $G$ acts on $V/V^{U}$, and the natural quotient map $\pi:V\to V/V^{U}$ is $G$-equivariant. Now,
$\lim_{t\to\infty}\varphi(t)\cdot\pi(v)=\lim_{t\to\infty}\pi(\varphi(t)\cdot v)=0.$
Since $\dim(V/V^{U})<\dim(V)$, we get that $\pi(v)=0$. Namely, $v\in V^{U}$, and again, since $V$ is of smallest dimension with a $\varphi(t)$-contracting nonzero vector, we get $V=V^{U}$. Since $U$ acts trivially on $V$, the action of $G$ on $V$ factors through the action of $G/U$. Therefore, $\varphi(t)\cdot v=(q\circ\varphi(t))\cdot v$. Since $\varphi(t)\cdot v\to 0$ as $t\to\infty$, and $q\circ\varphi$ is non-contracting, we get a contradiction. ∎
Lemma 2.19.
Let $G_{1}$ and $G_{2}$ be the real points of algebraic groups over $\mathbb{R}$. Suppose that $f:G_{1}\to G_{2}$ is a homomorphism of algebraic groups, and $\varphi:[0,\infty)\to f(G_{1})$ is a curve definable in $\mathscr{S}$. Then there exists a definable lift $\psi:[0,\infty)\to G_{1}$ of $\varphi$. Namely, $\psi$ is definable in $\mathscr{S}$, and $f\circ\psi=\varphi$.
Proof.
Consider the definable set
$\mathcal{D}:=\{(t,g):g\in G_{1},~{}t\geq 0,~{}f(g)=\varphi(t)\}.$
By the choice function theorem (see Theorem B.7), there exists a definable curve $\psi:[0,\infty)\to G_{1}$ such that $(t,\psi(t))\in\mathcal{D}$ for all $t\geq 0$. ∎
Lemma 2.20.
Let $G_{1}$ and $G_{2}$ be the $\mathbb{R}$-points of semi-simple connected algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a surjective homomorphism with finite kernel. Let $\psi:[0,\infty)\to G_{1}$ be a definable curve. Suppose that $f\circ\psi$ is non-contracting for $G_{2}$, then $\psi$ is non-contracting for $G_{1}$.
Proof.
Let $V$ be a finite-dimensional irreducible rational representation of $G_{1}$. By Schur’s lemma, the center of $G_{1}$ acts via a character on $V\otimes\mathbb{C}$. Then, $\ker f$, which is central in $G_{1}$, acts trivially on the $n$-fold tensor product $W=\otimes^{n}(V\otimes\mathbb{C})$, where $n$ is the order of $\ker f$. Therefore, the action of $G_{1}$ on $W$ factors through an action of $G_{1}/(\ker f)\cong G_{2}$ on $W$. Let $v\in V$. If $\psi(t)\cdot v\to 0$, then $f\circ\psi(t)\cdot(v\otimes\cdots\otimes v)\to 0$ as $t\to\infty$. But, since $f\circ\psi$ is non-contracting, we get that $v=0$. ∎
Lemma 2.21.
Let $G_{1}$ and $G_{2}$ be the real points of affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a surjective homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{2}$ be a non-contracting curve in $G_{2}$ definable in $\mathscr{S}$. Then, $\varphi$ has a $\mathscr{S}$-definable lift $\psi:[0,\infty)\to G_{1}$ which is non-contracting in $G_{1}$.
Proof.
In this proof, all subgroups considered are the open subgroups of $\mathbb{R}$-points of algebraic groups defined over $\mathbb{R}$.
Since the hull is generated by one-parameter unipotent subgroups, by modifying $\varphi$ with a bounded correcting curve and replacing $G_{2}$ by $(G_{2})_{u}$, we may assume without loss of generality that the radical of $G_{2}$ is unipotent, and we denote it by $U_{2}$. Let $\bar{G}_{2}=G_{2}/U_{2}$ and $q_{2}:G_{2}\to\bar{G}_{2}$ be the quotient homomorphism. Let $U_{1}$ be the unipotent radical of $G_{1}$. Then, there exists a reductive subgroup $M_{1}$ of $G_{1}$ such that $G_{1}=M_{1}U_{1}$ and $M_{1}\cap U_{1}=\{e\}$, see [RA72, Page 11]. Since $f$ is surjective, $f(U_{1})$ is a unipotent normal subgroup of $G_{2}$, and hence $f(U_{1})\subseteq U_{2}$. In fact, $f(U_{1})=U_{2}$. This can be proved as follows. $f^{-1}(U_{2})=(M_{1}\cap f^{-1}(U_{2}))U_{1}$. Since $f^{-1}(U_{2})$ is a normal subgroup of $G_{1}$, $M_{1}\cap f^{-1}(U_{2})$ is reductive. So $f(M_{1}\cap f^{-1}(U_{2}))$ being a reductive subgroup of $U_{2}$, it is trivial. Since $f$ is surjective, we get $f(U_{1})=U_{2}$. Therefore, $q_{2}\circ f:M_{1}\to\bar{G}_{2}$ is surjective. Since $\bar{G}_{2}$ is semisimple, $q_{2}\circ f([M_{1},M_{1}])=\bar{G}_{2}$. Let $L_{2}=[M_{1},M_{1}]\cap\ker(q_{2}\circ f)$. Since $L_{2}$ is a normal subgroup of the semisimple group $[M_{1},M_{1}]$, there exists a semisimple normal subgroup $L_{1}$ of $[M_{1},M_{1}]$ such that $[M_{1},M_{1}]=L_{1}L_{2}$, where $L_{1}\cap L_{2}$ is a finite central subgroup of $L_{1}$. Therefore, $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is a surjective homomorphism whose kernel is $L_{1}\cap L_{2}$. Also, since $\bar{G}_{2}$ is generated by unipotent one-parameter subgroups, it has no compact simple normal subgroup. Therefore, $L_{1}$ also does not have a compact normal subgroup, and $L_{1}$ is generated by unipotent one-parameter subgroups.
Let $\tilde{G}_{1}=L_{1}U_{1}$. Since $q_{2}\circ f(L_{1})=G_{2}/U_{2}$ and $f(U_{1})=U_{2}$, we have $f:\tilde{G}_{1}\to G_{2}$ is a surjective map. Therefore, by Lemma 2.19, there exists a definable $\psi:[0,\infty)\to\tilde{G}_{1}$ such that $f\circ\psi=\varphi$. Let $q_{1}:\tilde{G}_{1}\to\tilde{G}_{1}/U_{1}$ be the natural quotient map. Since $q_{1}:L_{1}\to\tilde{G}_{1}/U_{1}$ is surjective (it is an isomorphism), let $\tilde{\psi}:[0,\infty)\to L_{1}$ be a definable curve such that $q_{1}\circ\tilde{\psi}=q_{1}\circ\psi$. Since $U_{1}\subseteq\ker(q_{2}\circ f)$, $q_{2}\circ f$ factors through $q_{1}$. Therefore, we have $(q_{2}\circ f)\circ\tilde{\psi}=(q_{2}\circ f)\circ\psi=q_{2}\circ\varphi$. Now $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is surjective with finite central kernel, and $q_{2}\circ\varphi$ is definable and non-contracting in $\bar{G}_{2}$. Therefore, by Lemma 2.20, $\tilde{\psi}$ is non-contracting in $L_{1}$. Therefore, $q_{1}\circ\psi=q_{1}\circ\tilde{\psi}$ is non-contracting in $\tilde{G}_{1}/U_{1}$. Therefore, by Lemma 2.18, $\psi$ is non-contracting in $\tilde{G}_{1}$. Finally, by Lemma 2.17 we get that $\psi$ is non-contracting in $G_{1}$. ∎
The following observation extends Lemma 2.20.
Proposition 2.22.
Suppose that $G_{1}$ is a real algebraic group generated by unipotent one-parameter subgroups. Let $f:G_{1}\to G_{2}$ be a homomorphism of real algebraic groups with finite kernel. Then a definable curve $\varphi:[0,\infty)\to G_{1}$ is non-contracting in $G_{1}$ if and only if $f\circ\varphi$ is non-contracting in $G_{2}$.
Proof.
If $\varphi$ is non-contracting in $G_{1}$, then by Lemma 2.17, it follows that $f\circ\varphi$ is non-contracting in $G_{2}$.
Now suppose that $f\circ\varphi$ is non-contracting in $G_{2}$. Since $G_{1}$ is generated by unipotent one-parameter subgroups, $f(G_{1})$ is a subgroup of $G_{2}$ generated by unipotent one-parameter subgroups. Therefore, the solvable radical of $f(G_{1})$ is unipotent. Hence, $f(G_{1})$ is an observable subgroup of $G_{2}$, see [Observable_group_grosshans, Corollary 2.8]. Therefore, given any finite-dimensional rational representation of $V$ of $f(G_{1})$, $V$ is a subrepresentation of a finite-dimensional representation $W$ of $G_{2}$ restricted to $f(G_{1})$. Let $v\in V$. Suppose that $f\circ\varphi(t)\cdot v\to 0$. Since $v\in V\subseteq W$ and $f\circ\varphi$ is non-contracting in $G_{2}$, we conclude that $v=0$. This proves that $f\circ\varphi$ is non-contracting for $f(G_{1})$. Therefore, replacing $G_{2}$ by $f(G_{1})$, we may assume that $f$ is surjective. Therefore, by Lemma 2.21 there exists a lift $\psi:[0,\infty)\to G_{1}$ of $f\circ\varphi$ which is definable and non-contracting in $G_{1}$. By the monotonicty theorem (see Theorem B.5), there is a $T_{0}\geq 0$ such that $\varphi$ and $\psi$ are continuous on $[T_{0},\infty)$. Let $z=\varphi(T_{0})\psi(T_{0})^{-1}\in\ker f$. So, $t\mapsto z\psi(t)$ and $t\mapsto\varphi(t)$ are two lifts of $f\circ\varphi$ from $[T_{0},\infty)$ to $G_{1}$ from the point $\varphi(T_{0})\in G_{1}$. Since $\ker f$ is finite, $f$ is a covering map, we get $\varphi(t)=z\psi(t)$ for all $t\in[0,\infty)$. Since $\psi$ is non-contracting, so is $\varphi$. ∎
3. $(C,\alpha)$-good property
In our setting, a property more fundamental than the $(C,\alpha)$-good property (Theorem 1.12) is the following.
Definition 3.1.
Let $\delta\in(0,1]$. A family $\tilde{\mathscr{F}}$ of real functions defined on $[T_{0},\infty)$ is called $\delta$-good if there exists a constant $M(\delta)$ depending only on $\delta$ such that
(3.1) $\frac{\|f\|_{I}}{\|f\|_{I_{\delta}}}\leq M(\delta),$
for all $f\in\tilde{\mathscr{F}}$ and all bounded sub-intervals $I_{\delta}\subseteq I\subseteq[T_{0},\infty]$ satisfying $|I_{\delta}|=\delta|I|$.
Such inequalities are well-known in the literature as Remez-type inequalities; see [Remez_ineq].
We will prove the following result, from which Theorem 1.12 will be deduced. For the rest of this section, we fix a polynomially bounded o-minimal structure $\mathscr{S}_{\mathrm{pbd}}$.
Theorem 3.2.
Let $\mathcal{V}$ be a finite-dimensional vector space of real functions defined on $[0,\infty)$ definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that
$\lim_{t\to\infty}f(t)\neq 0,\ \forall f\in\mathscr{F}\setminus\{0\}.$
Then, there exists $T_{0}>0$ such that the family of functions $\tilde{\mathscr{F}}$ of the nonzero functions in $\mathscr{F}$ restricted to $[T_{0},\infty)$ is $\delta$-good, for all $\delta\in(0,1]$.
Remark 3.3.
The $T_{0}$ in Theorem 3.2 can be chosen as a $T_{0}$, which is obtained by the outcome of Lemma 3.8.
We now proceed to prove Theorem 1.12 by assuming Theorem 3.2, and the rest of the section will be dedicated to proving Theorem 3.2.
3.1. Proving Theorem 1.12 via $\delta$-goodness
We first note the following corollary of Theorem 3.2.
Corollary 3.4.
Let $\mathcal{V}$ be a finite-dimensional vector space of real functions on $[0,\infty)$ which are definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\mathscr{F}\subseteq\mathcal{V}$ is a closed definable cone such that for all $f\in\mathscr{F}\setminus\{0\}$,
$\lim_{t\to\infty}f(t)\neq 0.$
Let $T_{0}\geq 1$ such that the outcome of Theorem 3.2 holds. Then, there exist $\mathrm{m},r>0$ such that:
(3.2) $\|f\|_{I}\leq\mathrm{m}x^{r}\|f\|_{I^{\prime}},$
for all $x\geq 1$, $f\in\mathscr{F}\setminus\{0\}$, and bounded intervals $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$.
Proof.
Consider the definable set (see Lemma B.3):
$S:=\left\{(x,y):\forall I^{\prime}\subseteq I\subseteq[T_{0},\infty),~{}\frac{|I|}{|I^{\prime}|}\leq x,~{}\frac{\|f\|_{I}}{\|f\|_{I^{\prime}}}\leq y,~{}\forall 0\neq f\in\mathscr{F}\right\}.$
By Theorem 3.2, the projection of $S$ to the first coordinate includes $[1,\infty).$ Then, by the definable choice function theorem (see Theorem B.7), there exists a function $\phi:[1,\infty)\to\mathbb{R}^{2}$ definable in the same polynomially bounded o-minimal structure, such that
$(x,\phi(x))\in S,~{}\forall x\geq 1.$
Thus, for all $x>1$, $f\in\mathscr{F}\setminus\{0\}$, and intervals $I^{\prime}\subseteq I\subseteq[T_{0},\infty)$ such that $\frac{|I|}{|I^{\prime}|}\leq x$, it holds that
$\|f\|_{I}\leq\phi(x)\|f\|_{I^{\prime}}.$
Since $\phi$ is polynomially bounded, there exist $r>0$ and $T_{1}\geq 1$ such that $\frac{\phi(x)}{x^{r}}\leq c_{0}$ for all $x\geq T_{1}$, for some $c_{0}>0$. Finally, since $\frac{\phi(x)}{x^{r}}$ is bounded in $[1,T_{1}]$, the result follows. ∎
Our proof below is inspired by [Klein_Marg_flows_and_dioph, proof of Proposition 3.2].
Proof of Theorem 1.12 assuming Theorem 3.2.
We recall that the number of connected components of the family of definable sub-level sets
$\{t\in I:|f(t)|\leq\epsilon\},~{}\epsilon>0,~{}I\subseteq[T_{0},\infty),~{}f\in\mathscr{F}\setminus\{0\},$
is bounded uniformly, say by $K$, see Theorem B.6. Now fix $I\subseteq[T_{0},\infty),~{}\epsilon>0$ and an $f\in\mathscr{F}\setminus\{0\}$. Let
$I^{\prime}\subseteq\{t\in I:|f(t)|\leq\epsilon\}$
be an interval of maximum length. Then
$L\leq K|I^{\prime}|\text{, where }L:=|\{t\in I:|f(t)|\leq\epsilon\}|.$
Therefore, ${|I|}/{|I^{\prime}|}\leq{|I|}/{(L/K)}$. By Corollary 3.4, we get
$\|f\|_{I}\leq\mathrm{m}\left(\frac{|I|}{L/K}\right)^{r}\|f\|_{I^{\prime}}\leq\mathrm{m}\left(\frac{|I|}{L/K}\right)^{r}\epsilon.$
Therefore,
|$|\{t\in I:|f(t)|\leq\epsilon\}|=L\leq\mathrm{m}^{\frac{1}{r}}K\left(\frac{\epsilon}{\|f\|_{I}}\right)^{\frac{1}{r}}|I|.$
Thus, $f$ is $(\mathrm{m}^{\frac{1}{r}}K,1/r)$-good on $[T_{0},\infty)$, see (1.6). ∎
3.2. Proving Theorem 3.2
The space of polynomials of bounded degrees is the prototypical example of $\delta$-good functions. We record the following well-known result, as it will play a role in the proof of Theorem 3.2.
Proposition 3.5 ([Dani1993LimitDO, Lemma 4.1]).
Fix $n\in\mathbb{N}$. Then, for every $\delta\in(0,1]$, nested intervals $I_{\delta}\subseteq I$ with $|I_{\delta}|=\delta|I|>0$, and every nonzero polynomial $c_{0}+c_{1}x+\cdots+c_{n}x^{n}$, it holds that:
(3.3) $\frac{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I}}{\|c_{0}+c_{1}x+\cdots+c_{n}x^{n}\|_{I_{\delta}}}\leq(n+1)\frac{n^{n}}{\delta^{n}}.$
We will be using the following convenient fact. Recall by the monotonicity theorem (Theorem B.5) that the derivative of a definable function $f(t)$ on an unbounded interval exists for all large $t$.
Lemma 3.6.
[Miller94, Proposition 3.1] Let $f:[0,\infty)\to\mathbb{R}$ be a function definable in $\mathscr{S}_{\mathrm{pbd}}$ such that $f(t)\sim t^{r}$ as $t\to\infty$, where $r\neq 0$. Then, $f^{\prime}(t)\sim rt^{r-1},\text{ as }t\to\infty$.
Lemma 3.7.
Let $F_{i}:[0,\infty)\to\mathbb{R}$ be functions definable in $\mathscr{S}_{\mathrm{pbd}}$, for $i=0,1,\ldots,N$, and assume that none of the $F_{i}$ are eventually zero. Suppose that the degrees $\deg(F_{0}),\ldots,\deg(F_{N})$ are all distinct. Then, there exists a $T_{0}>0$ such that for any interval $I\subseteq[T_{0},\infty)$ of positive length, it holds that
$\|c_{0}F_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}>0$
for all $(c_{0},c_{1},\ldots,c_{N})\neq 0$.
Proof.
By o-minimality (Theorem B.6), the definable set $F_{0}^{-1}(0)$ has finitely many connected components. Since $F_{0}$ is not eventually zero, for some $T_{0}>0$, $F_{0}(t)\neq 0$ for all $t\geq T_{0}$. Thus for all $I\subseteq[T_{0},\infty)$ we have
$\left\|c_{0}F_{0}+c_{1}F_{1}+\cdots+c_{N}F_{N}\right\|_{I}>0\iff\left\|c_{0}+c_{1}\frac{F_{1}}{F_{0}}+\cdots+c_{N}\frac{F_{N}}{F_{0}}\right\|_{I}>0.$
Since the degrees of $F_{0},F_{1},...,F_{N}$ are distinct, the degrees of
$1,F_{1}/F_{0},...,F_{N}/F_{0},$
are distinct. Thus, we are reduced to proving the statement under the assumption that $F_{0}(t)=1$ for all $t\geq T_{0}$.
Pick $T_{1}\geq T_{0}$ such that the functions $F_{1},\ldots,F_{N}$ are differentiable in $[T_{1},\infty)$, see Theorem B.5. By Lemma 3.6, $\deg(F_{i}^{\prime})=\deg(F_{i})-1$ and $F_{i}^{\prime}$ is not eventually zero for $1\leq i\leq N$. So, the degrees of $F_{1}^{\prime},F_{2}^{\prime},\cdots,F_{N}^{\prime}$ are all distinct. Now, for any interval $I\subseteq[T_{1},\infty)$ of positive length,
if $\text{if }\|c_{0}+c_{1}F_{1}+...+c_{N}F_{N}\|_{I}=0\text{, then }\|c_{1}F_{1}^{\prime}+...+c_{N}F_{N}^{\prime}\|_{I}=0.$
In view of the above observations, the result is straightforward to deduce by induction. ∎
Next, for definable functions $F_{0}(t),F_{1}(t),...,F_{N}(t)$, we consider the Wronskian matrix $W(F_{0},...,F_{N})(t)$, which is defined to be the $(N+1)\times(N+1)$ matrix whose $k$-th row is $(F^{(k)}_{0}(t),F_{1}^{(k)}(t),\ldots,F_{N}^{(k)}(t))$ for $0\leq k\leq N$. The Wronskian is well defined for all $t$ large enough due to Theorem B.5.
Lemma 3.8.
Let the notation be as in Lemma 3.7. Then, there exists a $T_{0}>0$ such that the functions $F_{0},...,F_{N}$ are $(N+1)$-time continuously differentiable and $W(F_{0},...,F_{N})(t)$ is non-singular for all $t\geq T_{0}$.
Proof.
Since $\det(W(F_{0},...,F_{N})(t))$ is a definable function, either
$|\det(W(F_{0},...,F_{N})(t))|>0,$
for all large $t$ or $\det(W(F_{0},...,F_{N})(t))=0$ for all large $t$. We will prove the result by induction. Since $F_{0}$ is not eventually zero, the result follows for $N=0$. Now suppose $N\in\mathbb{N}$ and $T_{0}>0$ such that $\det(W(F_{0},...,F_{N-1})(t))\neq 0$ for all $t\geq T_{0}$. Now, suppose $\det(W(F_{0},...,F_{N})(t))=0$ for all $t\geq T_{1}\geq T_{0}$. Then, by a classical result of Bôcher [Bochner_certain_cases_Wronksian_implies_dep, Theorem II], $F_{0},\ldots,F_{N}$ are linearly dependent on $[T_{1},\infty)$. This contradicts Lemma 3.7. ∎
Lemma 3.9.
Let $\mathcal{V}$ be a real finite-dimensional vector space spanned by functions definable in $\mathscr{S}_{\mathrm{pbd}}$, and suppose that the zero function is the only function in $\mathcal{V}$ which is eventually constantly zero. Then, there exists a basis $\{F_{1},...,F_{N}\}$ of $\mathcal{V}$ such that:
(3.4) $\deg(F_{1})<\cdots<\deg(F_{N}).$
Proof.
We argue by induction. The claim is trivial for $N=1$. Now choose an arbitrary basis $\{H_{1},...,H_{N}\}$ of $\mathcal{V}$, and suppose without loss of generality that $\deg(H_{i})\leq\deg(H_{N})$ for all $i$. Moreover, we claim that there is no loss of generality in assuming that $\deg(H_{i})<\deg(H_{N})$, for all $i<N$. Indeed, for each $i<N$ such that $\deg(H_{i})=\deg(H_{N})$, we can find $c_{i}\neq 0$ such that $\deg(H_{i}-c_{i}H_{N})<\deg(H_{N})$. By replacing $H_{i}$ with $H_{i}-c_{i}H_{N}$, we again get a basis for $\mathcal{V}$. Note that the vector space $\mathcal{U}:=\text{Span}_{\mathbb{R}}\{H_{1},...,H_{N-1}\}$ is $(N-1)$-dimensional, and $\deg(f)<\deg(H_{N})$ for all $f\in\mathcal{U}$. By induction, there is a basis $\{F_{1},...,F_{N-1}\}$ for $\mathcal{U}$ such that $\deg(F_{i})<\deg(F_{i+1})$ for all $i$. Then, $\{F_{1},...,F_{N-1},F_{N}\}$ with $F_{N}:=H_{N}$ is the required basis for $\mathcal{V}$. ∎
Remark 3.10.
If $\mathcal{V}$ is a finite-dimensional vector spanned by definable functions, then the subset of functions which are eventually constantly zero forms a finite-dimensional subspace $\mathcal{V}_{0}\subseteq\mathcal{V}$. Then, there exists a uniform $T_{0}>0$ such that for all $f\in\mathcal{V}_{0}$, it holds that $f(t)=0$, for all $t\geq T_{0}$. Thus, by restricting the functions in $\mathcal{V}$ to $[T_{0},\infty)$, we may assume $\mathcal{V}_{0}=\{0\}$.
Proof of Theorem 3.2.
Let $\mathcal{V}$ be a finite-dimensional vector space of functions $f:[0,\infty)\to\mathbb{R}$ definable in a polynomially bounded o-minimal structure. By Remark 3.10, there is no loss in generality in assuming that only the zero function is eventually constantly zero. Using Lemma 3.9, we may further assume, by possibly enlarging $\mathcal{V}$, that $\mathcal{V}$ has a basis of the form $\{F_{1}^{-},...,F_{n}^{-},F_{0}^{+},F_{1}^{+},...,F_{m}^{+}\}$ such that
$\deg(F_{1}^{-})<\cdots<\deg(F_{n}^{-})<0\leq\deg(F^{+}_{0})<\cdots<\deg(F_{m}^{+}),$
where $m,n\in\mathbb{N}$. Let $N=m+n+1$. We will denote
(3.5) $r_{i}=\deg(F^{+}_{i}),\quad-\kappa_{j}:=\deg(F^{-}_{j}),$
where $r_{i},\kappa_{j}\geq 0$. In addition, we assume for convenience that
$\lim_{t\to\infty}F^{+}_{i}(t)/t^{r_{i}}=1\text{ and }\lim_{t\to\infty}F^{-}_{j}(t)/t^{-\kappa_{j}}=1.$
For any $\underline{c}:=(c_{1}^{-},\ldots,c_{n}^{-},c_{0}^{+},\ldots,c^{+}_{m})\in\mathbb{R}^{N}$, we denote
$f^{-}_{\underline{c}}:=c^{-}_{1}F^{-}_{1}+\cdots+c^{-}_{n}F^{-}_{n},\ f^{+}_{\underline{c}}:=c^{+}_{0}F_{0}^{+}+\cdots+c^{+}_{m}F^{+}_{m}\text{, and }f_{\underline{c}}=f^{-}_{\underline{c}}+f^{+}_{\underline{c}}.$
Let $\mathscr{F}\subseteq\mathcal{V}$ be a closed definable cone such that
$\lim_{t\to\infty}f(t)\neq 0,\,\forall f\in\mathscr{F}\setminus\{0\}.$
Let ~:=$\tilde{\mathscr{F}}:=\{\underline{c}\in\mathbb{R}^{N}:f_{\underline{c}}\in\mathscr{F}\}\cong\mathscr{F}$. Then $\tilde{\mathscr{F}}$ is a closed cone in $\mathbb{R}^{N}$.
Fix a $T_{0}\geq 1$ for which the outcome of Lemma 3.8 holds, and suppose for contradiction that there exists a $\delta\in(0,1)$ such that $\mathscr{F}\setminus\{0\}$ is not $\delta$-good on $[T_{0},\infty)$. Consider the following definable subset (see Lemma B.3):
(3.6) $\mathcal{A}:=\left\{(s,\underline{c},a,l,\alpha):\begin{array}[]{c}s\geq 1,~{}\underline{c}\in\tilde{\mathscr{F}}\setminus\{0\},~{}a>T_{0},~{}l>0,\\ a\leq\alpha\leq a+l-\delta l,~{}\frac{\|f_{\underline{c}}\|_{[a,a+l]}}{\|f_{\underline{c}}\|_{[\alpha,\alpha+\delta l]}}>s\\ \end{array}\right\}.$
By the assumption for contradiction, the projection of $\mathcal{A}$ to the first coordinate equals $[1,\infty)$. By the choice function theorem (Theorem B.7), there exists a polynomially bounded definable curve
$\phi(s):=(\underline{c}(s),a(s),l(s),\alpha(s))\text{ for $s\geq 1$},$
such that $(s,\phi(s))\in\mathcal{A},~{}\forall s\geq 1.$ In particular, there exists an $C_{0}\neq 0$ and an $\kappa\in\mathbb{R}$ such that
(3.7) $\frac{a(s)}{l(s)}\sim C_{0}s^{\kappa},\text{ as $s\to\infty$}.$
The remainder of the proof is organized by dividing the argument into cases based on the sign of $\kappa$ and on whether $l(s)$ and $a(s)$ are bounded or unbounded. In each case, we derive a contradiction. To guide the reader, we begin with a brief overview.
If $\kappa\leq 0$, this implies that the length of the interval $I(s)=[a(s),a(s)+l(s)]$ grows faster than its starting point. The case where $l(s)$ is bounded is easier to handle, since in that case $I(s)$ remains a bounded interval. When $l(s)$ is unbounded, we exploit the fact that polynomially bounded definable functions $f(t)$ are asymptotically equivalent to functions of the form $ct^{r}$.
On the other hand, if $\kappa>0$, the intervals $I(s)$ are short relative to their left endpoints. In this case, we can apply Taylor approximation, and use the $\delta$-goodness of polynomials (see Proposition 3.5) to reach a contradiction.
Before proceeding to the case analysis, we record the following observation, which will be useful throughout the proof. Let $\|\cdot\|_{1}$ denote the $L^{1}$-norm on $\mathbb{R}^{N}$, and define
(3.8) $\hat{\underline{c}}(s):=\frac{\underline{c}(s)}{\|\underline{c}(s)\|_{1}},\ s\geq 1.$
The function $s\mapsto\hat{\underline{c}}(s)$ is a bounded definable function with $\|\hat{\underline{c}}(s)\|_{1}=1$. It follows that
$\lim_{s\to\infty}\hat{\underline{c}}(s)=\mathbf{v}$
for some $\mathbf{v}\in\mathbb{R}^{N}$ satisfying $\|\mathbf{v}\|_{1}=1.$ Since $\underline{c}(s)\in\tilde{\mathscr{F}}\setminus\{0\}$ and $\tilde{\mathscr{F}}$ is a closed cone, we conclude that $\mathbf{v}\in\tilde{\mathscr{F}}\setminus\{0\}$.
Case 1: $\kappa\leq 0$. There are two sub-cases: $\{l(s):s\geq 1\}$ is bounded or not.
Case 1.1. $l(s)$ is bounded in $s$. Then $a(s)$ is also bounded, and by o-minimality $l(s)$ and $a(s)$ converge as $s\to\infty$. We observe that
$\lim_{s\to\infty}l(s)\neq 0.$
Because, since $a(s)\geq T_{0}\geq 1,\forall s\geq 1$, if $\lim_{s\to\infty}l(s)=0$, then $\lim_{s\to\infty}\frac{a(s)}{l(s)}=\infty$, which contradicts the assumption that $\kappa\leq 0$. We also note that the end-points of the nested intervals
$[\alpha(s),\alpha(s)+\delta l(s)]\subseteq[a(s),a(s)+l(s)],$
converge to the end-points of intervals $I_{\delta}\subseteq I\subseteq[T_{0},\infty)$ of positive length, where $\frac{|I_{\delta}|}{|I|}=\delta$. Hence $\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}$ is uniformly bounded in $s$. Now,
$\frac{\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}$ $=\frac{\|f_{\underline{c}(s)}/\lVert\underline{c}(s)\rVert_{1}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}/\lVert\underline{c}(s)\rVert_{1}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}$
$=\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}>s,$
so that
$\frac{\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}}{s}\geq\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}.$
By taking $s\to\infty$, we get
$0=\|f_{\textbf{v}}\|_{I_{\delta}}.$
This is a contradiction to Lemma 3.7.
Case 1.2. $l(s)$ is unbounded in $s$. In this case, we will achieve a contradiction by proving that:
$\lim_{s\to\infty}\hat{}\underline{c}(s)=(v_{1}^{-},...,v_{n}^{-},0,...,0),$
where $(v_{1}^{-},...,v_{n}^{-})\neq 0.$ This leads to a contradiction, because
$f_{(v_{1}^{-},...,v_{n}^{-},0,...,0)}\in\mathscr{F}\setminus\{0\}\text{ and }\lim_{t\to\infty}f_{(v_{1}^{-},...,v_{n}^{-},0,...,0)}(t)=0,$
is contrary to our assumption on $\mathscr{F}$.
Remark 3.11.
This is the only place in the proof of Theorem 1.12 where we use the non-decaying function assumption for the closed cone $\mathscr{F}$, and it highlights the necessity of the non-contracting assumption for obtaining the $C,\alpha$-good property.
We now proceed with our proof. Since $l(s)$ is an unbounded definable function, we have that $l(s)\to\infty$. For $s\geq 1$, we define
(3.9) $C^{+}(s):=\sum_{i=0}^{m}|c^{+}_{i}(s)|l(s)^{r_{i}},\,C^{-}(s):=\sum_{j=1}^{n}|c^{-}_{j}(s)|\text{, and }C(s):=C^{+}(s)+C^{-}(s).$
Since $s\mapsto\frac{(c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}$ is definable and bounded, we get that:
(3.10) $(w^{+}_{0},...,w^{+}_{m}):=\lim_{s\to\infty}\frac{(c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}\in[0,1]^{m+1}.$
Claim.
There exists an interval $J\subset[0,1]$ of positive length such that
$w_{0}^{+}t^{r_{0}}+\cdots+w_{m}^{+}t^{r_{m}}=0,~{}\forall t\in J.$
Here, $r_{i}$ are as in (3.5). In particular, $(w^{+}_{0},...,w^{+}_{m})=0$.
Proof of Claim.
We consider
$\varphi_{{\underline{c}}(s)}(t):=\frac{f_{\underline{c}(s)}(a(s)+tl(s))}{C(s)},\ \forall t\in[0,1].$
Recall that, by our choice function for the definable set (3.6),
(3.11) $s<\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}=\frac{\|\varphi_{{\underline{c}}(s)}\|_{[0,1]}}{\|\varphi_{{\underline{c}}(s)}\|_{I_{\delta}(s)}},$
where $I_{\delta}(s):=\left[\frac{\alpha(s)-a(s)}{l(s)},\frac{\alpha(s)-a(s)+\delta l(s)}{l(s)}\right]\subseteq[0,1]$ having length $|I_{\delta}(s)|=\delta$ for all $s\geq 1.$ Since $\frac{a(s)}{l(s)}$ is a bounded definable function, we have
$\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}\in[0,\infty).$
Also, since the end-points of $I_{\delta}(s)$ are bounded, they converge to end-points of a sub-interval $I_{\delta}$ of $[0,1]$ of length $\delta$. We will prove that:
(3.12) $\limsup_{s\to\infty}\lVert\varphi_{{\underline{c}}(s)}\rVert_{I_{\delta}(s)}=0,$
and that for any $t_{0}\in(0,1)$:
(3.13) $\lim_{s\to\infty}\sup_{t\in[t_{0},1]}\lvert\varphi_{\underline{c}(s)}(t)-w^{+}_{0}\left(x_{0}+t\right)^{r_{0}}+\cdots+w^{+}_{m}\left(x_{0}+t\right)^{r_{m}}\rvert=0.$
We note that the claim readily follows from (3.12) and (3.13). To prove (3.12), it is sufficient to show that
$\sup_{s\to\infty}\|\varphi_{\underline{c}(s)}\|_{[0,1]}<\infty,$
see (3.11). For each $0\leq i\leq m$, there exists a $\nu_{i}>0$ such that for all $t\in[0,1]$, we have as $l\to\infty$,
$\frac{F^{+}_{i}(a+tl)}{l^{r_{i}}}$ $=\frac{\left(a+lt\right)^{r_{i}}}{l^{r_{i}}}+O\left(\frac{\left(a+lt\right)^{r_{i}-\nu_{i}}}{l^{r_{i}}}\right)$
(3.14) $=\left(\frac{a}{l}+t\right)^{r_{i}}+O\left(l^{-\nu_{i}}\left(\frac{a}{l}+t\right)^{r_{i}-\nu_{i}}\right).$
We can pick $S_{0}\geq 1$ such that $l(s)\geq 1$ for all $s\geq S_{0}$ and that the following two statements hold:
  1. (1)
    $\sup_{t\in[0,1]}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}$ is uniformly bounded in $s\geq S_{0}$ and $0\leq i\leq m$.
    Indeed, because $r_{i}\geq 0$, $\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}$, and $t\in[0,1]$.
  2. (2)
    $\sup_{t\in[0,1]}l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}$ is uniformly bounded in $s\geq S_{0}$ and $0\leq i\leq m$.
    This is clear if $r_{i}-\nu_{i}\geq 0$. Now suppose that $r_{i}-\nu_{i}<0$. Then, for all $t\in[0,1]$,
    $l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}\leq$ $l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}\right)^{r_{i}-\nu_{i}}$
    $=$ $\frac{a(s)^{r_{i}-\nu_{i}}}{l(s)^{r_{i}}}=O(1),$
    since $\frac{a(s)}{l(s)}=O(s^{\kappa})$ with $\kappa\leq 0$, and since $a(s)\geq T_{0}\geq 1$.
By (2), as $s\to\infty$, uniformly in $t\in[0,1]$,
(3.15) $f^{+}_{\underline{c}(s)}(a(s)+tl(s))=\sum_{i=0}^{m}(c_{i}^{+}(s)l(s)^{r_{i}})\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}+O(C^{+}(s)).$
In addition, we observe that as $s\to\infty$:
(3.16) $\sup_{t\in[0,1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)).$
Here, $C^{+},C^{-}$ are given by (3.9). Now:
$|\varphi_{{\underline{c}}(s)}(t)|$ $=\frac{|f_{\underline{c}(s)}(a(s)+tl(s))|}{C(s)}$
$=\frac{|f^{+}_{\underline{c}(s)}(a(s)+tl(s))+f_{\underline{c}(s)}^{-}(a(s)+tl(s))|}{C(s)}=O(1),~{}\text{as $s\to\infty$,}$
uniformly in $t\in[0,1]$. This proves (3.12).
Next, we prove (3.13). We have $\lim_{s\to\infty}l(s)=\infty$, $a(s)/l(s)>0$, and $\lim_{s\to\infty}a(s)/l(s)=x_{0}$. Fix $t_{0}\in(0,1)$. Then $t_{0}\leq\lvert a(s)/l(s)+t\rvert\leq x_{0}+2$ for all $t\in[t_{0},1]$, for all large $s$. So, for any $0\leq i\leq m$, by (3.14), as $s\to\infty$,
$\sup_{t\in[t_{0},1]}\left|\frac{F^{+}_{i}(a(s)+tl(s))}{l(s)^{r_{i}}}-(x_{0}+t)^{r_{i}}\right|=O(l(s)^{-\nu_{i}}).$
Let $\nu=\min\{\nu_{0},\ldots,\nu_{m}\}>0$. Then, as $s\to\infty$, uniformly in $t\in[t_{0},1]$,
(3.17) $f^{+}_{\underline{c}(s)}(a(s)+tl(s))=\sum_{i=0}^{m}(c_{i}^{+}(s)l(s)^{r_{i}})(x_{0}+t)^{r_{i}}+O(C^{+}(s)l(s)^{-\nu}).$
Also,
(3.18) $\sup_{t\in[t_{0},1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)l(s)^{-\kappa_{n}}),$
where $\kappa_{n}$ is as in (3.5). By (3.17) and (3.18), for $t\in[t_{0},1]$, we have
||| $\varphi_{\underline{c}(s)}(t)-(w^{+}_{0}(x_{0}+t)^{r_{0}}+\cdots+w^{+}_{m}(x_{0}+t)^{r_{m}})|$
$\leq\sum_{i=0}^{m}\left|\frac{c_{i}^{+}(s)l(s)^{r_{i}}}{C(s)}-w_{i}^{+}\right|(x_{0}+t)^{r_{i}}+O(l(s)^{-\kappa_{n}}+l(s)^{-\nu}),$
and the last expression decays to 00 uniformly for $t\in[t_{0},1]$, as $s\to\infty$. ∎
Using our claim and by recalling (3.10), we get that
$\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}$
$\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s)l(s)^{r_{0}},...,c_{m}^{+}(s)l(s)^{r_{m}})\|_{1}}$
$=(w^{-}_{1},...,w^{-}_{n},w^{+}_{0},...,w^{+}_{m})=(w^{-}_{1},...,w^{-}_{n},0,...,0).$
Therefore,
(3.19) $\|(w^{-}_{1},...,w^{-}_{n})\|_{1}=1.$
We now again consider ^$\hat{}\underline{c}(s)$, as in (3.8). We have
$\textbf{v}=(v^{-}_{1},...,v^{-}_{n},v^{+}_{0},...,v^{+}_{m})$ $:=\lim_{s\to\infty}\hat{}\underline{c}(s)$
$=\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}},$
where $\|\textbf{v}\|_{1}=1$ and $\textbf{v}\in\tilde{\mathscr{F}}\setminus\{0\}$. By (3.19), and since $l(s)\to\infty$ and $r_{i}\geq 0$:
11 11 $=\|(w_{1}^{-},...,w_{n}^{-})\|_{1}$
$=\lim_{s\to\infty}\frac{\|(c_{1}^{-}(s),...,c_{n}^{-}(s))\|_{1}}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})\|_{1}}$
$\leq\lim_{s\to\infty}\frac{\|(c_{1}^{-}(s),...,c_{n}^{-}(s))\|_{1}}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}}$
$=\|(v_{1}^{-},...,v_{n}^{-})\|_{1}\leq 1.$
Therefore, $\|(v_{1}^{-},...,v_{n}^{-})\|_{1}=1$. As a consequence, we get $v^{+}_{0}=\cdots=v^{+}_{m}=0.$
Case 2: $\kappa>0$. Here, we consider two cases corresponding to $a(s)$ being bounded or $a(s)\to\infty$.
Case 2.1. $a(s)$ is bounded. We have that $\lim_{s\to\infty}a(s)=x_{0}\geq T_{0}$. Since $\lim_{s\to\infty}\frac{a(s)}{l(s)}=\infty$, it follows that $\lim_{s\to\infty}l(s)=0$. We again consider ^$\hat{}\underline{c}(s)$, as in (3.8), and recall $\lim_{s\to\infty}\hat{\underline{c}}(s)=\textbf{v}$ with $\textbf{v}\in\tilde{\mathscr{F}}$ such that $\|\textbf{v}\|_{1}=1.$
By our choice of $T_{0}$, the basis functions $F^{-}_{i},F^{+}_{j}$ are $(N+1)$-times continuously differentiable, where $N:=n+m+1$. For $t\geq T_{0}$, let $W(t)$ denote the Wronskian matrix $W(F_{-n}(t),\ldots,F_{m}(t))$, whose $k$-th row is $(F_{-n}^{(k)}(t),\ldots,F_{m}^{(k)}(t))$, for $k=0,\dots,N$. The Taylor polynomial of order $N$ of the function $h\mapsto f_{\hat{\underline{c}(s)}}(a(s)+h)$ centered at $h=0$ is expressed by:
(3.20) $Q_{N,a(s)}(h)=\hat{\underline{c}}(s)W(a(s))^{\top}\cdot(h^{0},\ldots,h^{N}),$
where $\cdot$ denotes the dot product on $\mathbb{R}^{N+1}$. By Taylor’s theorem (mean value form for the remainder), for all $h\in[0,l(s)]$, there exists a $\xi\in a(s)+[0,l(s)]$ such that
$f_{\hat{\underline{c}}(s)}(a(s)+h)-Q_{N,a(s)}(h)=\frac{f_{\hat{\underline{c}}(s)}^{(N+1)}(\xi)}{(N+1)!}h^{N+1}.$
Since $a(s)$, $\hat{\underline{c}}(s)$, and $l(s)$ are bounded, we obtain that $|f_{\hat{\underline{c}}(s)}^{(N+1)}(\xi)|$ is uniformly bounded.
Denote $I(s):=[0,l(s)]$, and let $I_{\delta}(s):=[0,\delta l(s)]$. Thus, we conclude:
$\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I(s)}=\|Q_{N,a(s)}\|_{I(s)}+O(l(s)^{N+1}),$
and similarly,
(3.21) $\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}=\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1}).$
Using Proposition 3.5,
$\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I(s)}=$ $\|Q_{N,a(s)}\|_{I(s)}+O(l(s)^{N+1})$
$\leq$ $M_{N}(\delta)\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1})$
$\leq$ $M_{N}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}+O(l(s)^{N+1})$
$=$ $M_{N}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}\left[1+O\left(\frac{l(s)^{N+1}}{\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}}\right)\right].$
Thus,
$\frac{\|f_{{\underline{c}}(s)}\|_{a(s)+I(s)}}{\|f_{{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}}\ll 1+O\left(\frac{l(s)^{N+1}}{\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}}\right).$
We will now show that
(3.22) $\lim_{s\to\infty}\frac{l(s)^{N+1}}{\|Q_{N,a(s)}\|_{I_{\delta}(s)}}=0.$
Before proceeding with the proof, note that by combining (3.22) with (3.21), we obtain that $\frac{l(s)^{N+1}}{\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}}$ is bounded for all large $s$. This is a contradiction, since we are assuming that $s\leq\frac{\|f_{{\underline{c}}(s)}\|_{I(s)}}{\|f_{{\underline{c}}(s)}\|_{I_{\delta}(s)}}$ for all $s\geq 1$.
We proceed to prove (3.22). Observe that for a polynomial $q(x)=c_{0}+c_{1}x+...+c_{k}x^{k}$ of degree bounded by $k$ and an interval $J\subseteq[0,1]$, it follows by Proposition 3.5,
(3.23) $\|q\|_{J}\geq(k+1)^{-1}k^{-k}\|q\|_{[0,1]}|J|^{k}\gg\|(c_{0},c_{1},...,c_{k})\|_{1}|J|^{k}.$
Here, we have also used that $\|q\|_{[0,1]}$ and $\|\underline{c}\|_{1}$ are two norms on the $(k+1)$-dimensional space of polynomials of degree at most $k$, and we note that the implied constant depends only on $k$. By (3.20), $Q_{N,a(s)}$ is a polynomial of degree bounded by $N$ whose coefficients are given by $\hat{\underline{c}}(s)W(a(s))^{\top}$. We have that $W(a(s))^{\top}\to W(x_{0})^{\top}$ as $s\to\infty$, and we recall Lemma 3.8 which implies that $W(x_{0})^{\top}$ is non-singular. Therefore,
$\lVert\hat{\underline{c}}(s)W(a(s))^{\top}\rVert\gg\lVert\hat{\underline{c}}(s)\rVert_{1}\lVert(W(a(s))^{\top})^{-1}\rVert^{-1}=\lVert(W(x_{0})^{\top})^{-1}\rVert^{-1}>0$
for all $s\gg 1$. Then, by (3.23) applied to $J=[0,\delta l(s)]$, we get
(3.24) $\|Q_{N,a(s)}\|_{I_{\delta}(s)}\gg\lVert(W(x_{0})^{\top})^{-1}\rVert^{-1}\delta^{N}l(s)^{N}.$
Since $\delta$ is fixed, and $l(s)\to 0$, we get (3.22).
Case 2.2. $a(s)$ is unbounded. Here $a(s)\asymp s^{\theta}$, for $\theta>0$. We recall that for each $\mathrm{n}\in\mathbb{N}$, there is a $T_{\mathrm{n}}$ such that all basis functions $F^{-}_{i},F_{j}^{+}$ will be continuously differentiable $\mathrm{n}$ times in the ray $[T_{\mathrm{n}},\infty)$ for all $i,j$ (see Theorem B.5). By Lemma 3.6, it holds that
$\left|\frac{d^{p}}{dt^{p}}F^{-}_{i}(x)\right|\ll_{p}x^{-\kappa_{i}-p},\text{ and }\left|\frac{d^{q}}{dt^{q}}F^{+}_{j}(x)\right|\ll_{q}x^{r_{j}-q},$
where $\kappa_{i},r_{j}$ are as in (3.5), $p,q$ are non-negative integers, and the implied constants depend on $p,q$. Then, for all sufficiently large integers $\mathrm{n}$ such that $r_{i}-(\mathrm{n}+1)<0$ for all $0\leq i\leq m$, it holds that
$\sup_{t\in[a(s),a(s)+l(s)]}\left\lvert\frac{d^{\mathrm{n}+1}}{dt^{\mathrm{n}+1}}f_{\hat{\underline{c}}(s)}(t)\right\rvert=O_{\mathrm{n}}(a(s)^{r_{m}-(\mathrm{n}+1)}),$
where the implied constant depends on $\mathrm{n}$. By Taylor’s theorem, for the Taylor polynomial $P_{\mathrm{n},a(s)}(t)$ of $f_{\hat{\underline{c}}(s)}(t)$ of degree $\mathrm{n}$ centered around $a(s)$, we have
$\sup_{t\in[a(s),a(s)+l(s)]}|f_{\hat{\underline{c}}(s)}(t)-P_{\mathrm{n},a(s)}(t)|$ $=O_{\mathrm{n}}(a(s)^{r_{m}-(\mathrm{n}+1)}l(s)^{\mathrm{n}+1})$
$=O_{\mathrm{n}}\left(a(s)^{r_{m}}\left(\frac{l(s)}{a(s)}\right)^{\mathrm{n}+1}\right)$
(3.25) $=O_{\mathrm{n}}\left(s^{r_{m}\theta}s^{-(\mathrm{n}+1)\kappa}\right),$
where the last line follows by the assumption that $a(s)/l(s)\sim s^{\kappa}$ with $\kappa>0$. Due to Lemma 3.7, the function $\Psi(s):=\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}$ is positive. Since $\Psi$ is definable in a polynomially bounded o-minimal structure, there exists an $\eta>0$ such that
(3.26) $\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}\gg s^{-\eta},~{}\text{as }s\to\infty.$
We fix $\mathrm{n}$ large enough such that $\nu:=r_{m}\theta-(\mathrm{n}+1)\kappa<-\eta.$ We denote $J(s)=[a(s),a(s)+l(s)]$ and we let $J_{\delta}(s)=[\alpha(s),\alpha(s)+\delta l(s)]$. Then:
[…]
$=M$ (${}_{\mathrm{n}}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}\left[1+O\left(\frac{s^{\nu}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}\right)\right]$
[…]
Namely, $\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$ is bounded for all large $s$, contradicting the assumption $s<\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$. ∎
3.3. Relative time near varieties
The following proposition is an analog of [Dani1993LimitDO, Proposition 4.2], and it is key for the linearization technique.
Proposition 3.12.
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous, unbounded curve definable in $\mathscr{S}_{\mathrm{pbd}}$ such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\in\mathbb{R}^{m}\setminus\{0\}$. Fix a nonzero polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$, and denote
$\mathcal{A}:=\{x\in\mathbb{R}^{m}:Q(x)=0\}.$
Then, there exists a $T_{0}>0$ such that the following holds: for an $\epsilon>0$ and a compact subset $\emph{\text{K}}_{1}\subseteq\mathcal{A}$, there exists a compact subset $\emph{\text{K}}_{2}$ with
$\emph{\text{K}}_{1}\subseteq\emph{\text{K}}_{2}\subseteq\mathcal{A}$
such that for every compact neighborhood $\Phi$ of $K_{2}$ in $\mathbb{R}^{m}$, there exists a compact neighborhood $\Psi$ of $K_{1}$ in $\mathbb{R}^{m}$, with $\Psi\subseteq\mathring{\Phi}$, where $\mathring{\Phi}$ denotes the interior of $\Phi$, such that:
(3.27) |$|\{t\in[a,b]:\psi(t)v\in\Psi\}|\leq\epsilon(b-a),$
for all $v\in\mathbb{R}^{m}$ and $[a,b]\subseteq[T_{0},\infty)$, for which:
  • $\psi(t)v\in\Phi,~{}\forall t\in[a,b]$, and
  • $\psi(b)v\in\Phi\setminus\mathring{\Phi}$.
Remark 3.13.
Compared to [Dani1993LimitDO, Proposition 4.2], the statement our proposition is different in the assumption that the curve $\psi(t)v$ remains in $\Phi$ throughout the whole interval $[a,b]$, and exits $\mathring{\Phi}$ at $t=b$. This is a stronger requirement than the one in [Dani1993LimitDO, Proposition 4.2] which only asks for $\psi(t_{0})v\notin\mathring{\Phi}$ for some $t_{0}\in[a,b]$.
Remark 3.14.
Notice that for any affine variety $\mathcal{A}\subseteq\mathbb{R}^{m}$ there exists a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$ such that $\mathcal{A}$ is the zero set of $Q$. Indeed, by the Hilbert basis theorem, $\mathcal{A}$ is the zero set of finitely many polynomials $f_{1},...,f_{r}$, and we let $Q:=f_{1}^{2}+\cdots+f_{r}^{2}$.
Our proof of Proposition 3.12 builds on the following versions of the modified $(C,\alpha)$-good property.
Proposition 3.15.
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve definable in $\mathscr{S}_{\mathrm{pbd}}$, and let $Q:\mathbb{R}^{m}\to\mathbb{R}$ be a polynomial. Then there exist constants $\nu,C,\alpha,T_{0}>0$ such that for any $v\in\mathbb{R}^{m}$ exactly one of the following properties hold:
  1. (1)
    $Q(\psi(t)v)=0,\forall t\geq T_{0}$;
  2. (2)
    $\Theta_{v}(t):=t^{\nu}Q(\psi(t)v)$ is $(C,\alpha)$-good on $[T_{0},\infty)$.
Proof.
Consider the vector space
$\mathcal{V}:=\text{Span}_{\mathbb{R}}\{Q(\psi(t)v),\ t\geq 0:v\in\mathbb{R}^{m}\}.$
It is straightforward to verify that $\mathcal{V}$ is finite-dimensional, and that all functions in $\mathcal{V}$ are definable. In view of Remark 3.10, there is a $T_{0}\geq 0$ such that if $f\in\mathcal{V}$ is eventually zero, then $f(t)=0,\forall t\in[T_{0},\infty)$. We consider $\mathscr{F}$ to be the space of functions of $\mathcal{V}$ restricted to $[T_{0},\infty).$ By Lemma 3.9, there is a basis $\{f_{0},f_{1},...,f_{n}\}$ for $\mathscr{F}$ such that $\deg(f_{0})<\deg(f_{1})<...<\deg(f_{n})$. Let $\nu>0$ such that $\deg(t^{\nu}f_{0})\geq 0$. Then, by Theorem 1.12, we pick $T_{1}>0$ such that any nonzero function
$f\in t^{\nu}\otimes\mathscr{F}=\text{Span}_{\mathbb{R}}\{t^{\nu}Q(\psi(t)v),\ t\geq T_{1}:v\in\mathbb{R}^{m}\}$
is $(C,\alpha)$-good. ∎
We now note the following corollary.
Corollary 3.16.
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve, definable in $\mathscr{S}_{\mathrm{pbd}}$. Fix a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$. Then there exist $T_{0},C,\alpha>0$ such that for any $v\in\mathbb{R}^{m}$ and any interval $[x,y]\subseteq[T_{0},\infty)$, if
$\sup_{t\in[x,y]}|Q(\psi(t)v)|=|Q(\psi(y)v)|,$
then
(3.28) |$\left|\{t\in[x,y]:|Q(\psi(t)v)|\leq\eta\}\right|\leq C\left(\frac{\eta}{|Q(\psi(y)v)|}\right)^{\alpha}(y-x),$
for all $\eta>0$; we may say that $t\mapsto\lvert Q(\psi(t)v)\rvert$ is right-max-$(C,\alpha)$-good in $[T_{0},\infty)$.
Proof.
Fix $\nu,C,\alpha,T_{0}>0$ as in Proposition 3.15. Let $[x,y]\subseteq[T_{0},\infty)$ and let $v\in\mathbb{R}^{m}$ be a vector such that
$|Q(\psi(y)v)|=\delta>0,$
and such that
$|Q(\psi(t)v)|\leq\delta,\ \forall t\in[x,y].$
Then,
$\sup_{t\in[x,y]}|t^{\nu}Q(\psi(t)v)|=y^{\nu}\delta.$
Now, by the $(C,\alpha)$-good property of $|t^{\nu}Q(\psi(t)v)|$, we get:
|$\left|\{t\in[x,y]:|Q(\psi(t)v)|\leq\eta\}\right|$ =|$=\left|\{t\in[x,y]:|t^{\nu}Q(\psi(t)v)|\leq t^{\nu}\eta\}\right|$
|$\leq\left|\{t\in[x,y]:|t^{\nu}Q(\psi(t)v)|\leq y^{\nu}\eta\}\right|$
(3.29) $\leq C\left(\frac{y^{\nu}\eta}{y^{\nu}\delta}\right)^{\alpha}(y-x)=C\left(\frac{\eta}{\delta}\right)^{\alpha}(y-x).$
Proof for Proposition 3.12.
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous unbounded curve such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\neq 0$. Pick $T_{0},C,\alpha>0$ such that the $(C,\alpha)$-good property holds for the family
$\mathscr{F}_{\text{norm}}:=\{\|\psi(t)v\|,\ t\geq T_{0}:v\in\mathbb{R}^{m}\setminus\{0\}\},$
see Proposition 1.13, and such that the right-max-$(C,\alpha)$-property holds for the functions in
$\mathscr{F}_{\text{variety}}:=\{\lvert Q(\psi(t)v)\rvert,\ t\geq T_{0}:v\in\mathbb{R}^{m}\},$
which satisfy the requirements of Corollary 3.16.
Let $\epsilon>0$. Given a compact set $\text{K}_{1}\subseteq\mathcal{A}$, let $R_{1}>0$ be such that $\text{K}_{1}\subseteq B_{R_{1}}(0)$. We choose $R_{2}>R_{1}$ is such that $C(R_{1}/R_{2})^{\alpha}\leq\epsilon$, and we denote $\text{K}_{2}:=\overline{\text{B}_{R_{2}}(0)}\cap\mathcal{A}$; see (3.32) for the use of this choice. We define for $\epsilon^{\prime}>0$:
$\mathcal{A}^{\epsilon^{\prime}}:=\{x\in\mathbb{R}^{m}:|Q(x)|<\epsilon^{\prime}\}.$
Let $\Phi$ be a compact neighborhood of $\text{K}_{2}$. Then, we can pick an $\epsilon_{1}>0$ such that
(3.30) $B_{R_{2}}(0)\cap\mathcal{A}^{\epsilon_{1}}\subseteq\mathring{\Phi}.$
We denote $\epsilon_{2}:=\frac{R_{1}}{R_{2}}\epsilon_{1}$, and we note that $K_{1}\subseteq B_{R_{1}}(0)\cap\mathcal{A}^{\epsilon_{2}}$ and we observe that $C(\epsilon_{2}/\epsilon_{1})^{\alpha}\leq\epsilon$; see (3.33) for the use of this choice. We consider
(3.31) $\Psi:=\overline{B_{R_{1}}(0)\cap\mathcal{A}^{\epsilon_{2}}}.$
Let $v\in\mathbb{R}^{m}$ and an interval $[a,b]\subseteq[T_{0},\infty)$ such that:
  • $\psi(t)v\in\Phi,~{}\forall t\in[a,b]$, and
  • $\psi(b)v\in\Phi\setminus\mathring{\Phi}$.
Then by (3.30), $\|\psi(b)v\|\geq R_{2}$ or $|Q(\psi(b)v)|\geq\epsilon_{1}.$
Suppose that $\|\psi(b)v\|\geq R_{2}$. By (3.31) and by Proposition 1.13,
|$|\{t\in[a,b]|\psi(t)v\in\Psi\}|\leq$ |$|\{t\in[a,b]|\|\psi(t)v\|\leq R_{1}\}|$
(3.32) $\leq$ $C\left(\frac{R_{1}}{R_{2}}\right)^{\alpha}(b-a)\leq\epsilon(b-a).$
Suppose now that $|Q(\psi(b)v)|\geq\epsilon_{1}$. Since $t\mapsto|Q(\psi(t)v)|$ is continuous and definable, and since a definable map takes a particular value only finitely many times, there is a decomposition
$\{t\in[a,b]:|Q(\psi(t)v)|\leq\epsilon_{1}\}=[a_{1},b_{1}]\sqcup\cdots\sqcup[a_{k},b_{k}],$
for some $k\in\mathbb{N}$, such that for all $1\leq i\leq k$:
$|Q(\psi(b_{i})v)|=\epsilon_{1}\text{ and }|Q(\psi(t)v)|\leq\epsilon_{1},\forall t\in[a_{i},b_{i}].$
Since $\Psi\subseteq\mathcal{A}^{\epsilon_{2}}$, we have
$\{t\in[a,b]:\psi(t)v\in\Psi\}\subseteq\{t\in[a,b]:|Q(\psi(t)v)|\leq\epsilon_{2}\},$
and since $\epsilon_{2}<\epsilon_{1}$, we have
$\{t\in[a,b]:|Q(\psi(t)v)|\leq\epsilon_{2}\}=\bigsqcup_{i=1}^{k}\{t\in[a_{i},b_{i}]:|Q(\psi(t)v)|\leq\epsilon_{2}\}.$
We conclude by (3.28) of Corollary 3.16 that
(3.33) |$|\{t\in[a_{i},b_{i}]:|Q(\psi(t)v)|\leq\epsilon_{2}\}|\leq C\left(\frac{\epsilon_{2}}{\epsilon_{1}}\right)^{\alpha}(b_{i}-a_{i})\leq\epsilon(b_{i}-a_{i}),$
for each $1\leq i\leq k$. This proves that
|$|\{t\in[a,b]|\psi(t)v\in\Psi\}|\leq\epsilon(b-a).$
Thus, (3.27) holds in all cases. ∎
4. Non-escape of mass
Let $X^{*}:=X\cup\{\infty\}$ be the one-point compactification of
$X:=\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z}).$
In what follows, $\mathscr{S}_{\mathrm{pbd}}$ is a fixed polynomially bounded o-minimal structure. Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous and unbounded curve. For $x\in X$ and $T>0$ let $\mu_{T,\varphi,x}$ be the measure defined in (1.1). By the Banach-Alaoglu theorem, any weak-star limit $\mu$ as $T\to\infty$ of the measures $\mu_{T,\varphi,x}$ is a probability measure on $X^{*}$. We now state and prove the non-divergence property in the setup of SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$. Later in the section, we extend the result to the case of finite volume quotient spaces of Lie groups. In the general case, when the Lie group is of higher rank, the non-divergence result builds on the non-divergence in SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ by using Margulis’ arithmeticity theorem see [Zimmer:book84, page 3].
Proposition 4.1.
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Then, there exists $T_{0}>0$ such that given a compact set $F\subseteq\text{SL}(n,\mathbb{R})$ and $\epsilon>0$, there exists a compact set $K\subseteq\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that for any $g\in F$ and $T\geq T_{0}$, we have
|$\lvert\{t\in[T_{0},T]:\varphi(t)g\text{SL}(n,\mathbb{Z})/\text{SL}(n,\mathbb{Z})\in K\}\rvert\geq(1-\epsilon)(T-T_{0}).$
We need the following preliminaries for the proof. In the discussion that follows, we will refer to rank-$k$ discrete subgroups of $\mathbb{R}^{n}$ as $k$-lattices. For a $k$-lattice $\Lambda\subseteq\mathbb{Z}^{n}$, let $\|\Lambda\|$ be the volume of a fundamental domain $F\subseteq\text{Span}_{\mathbb{R}}\{\Lambda\}$ of $\Lambda$ with respect to the usual measure on the subspace $\text{Span}_{\mathbb{R}}\{\Lambda\}\subseteq\mathbb{R}^{n}$ (obtained by restricting the usual Euclidean inner product). Recall that if $\{\textbf{v}_{1},...,\textbf{v}_{k}\}$ forms a $\mathbb{Z}$-basis for $\Lambda$, then
$\|\Lambda\|=\|\textbf{v}_{1}\wedge\cdots\wedge\textbf{v}_{k}\|.$
The norm above is defined by choosing the inner product for which the pure wedges of $k$-positively oriented tuples of the canonical basis vectors $e_{i},i=1,...,n$ are orthonormal. Consider the representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
(4.1) $g\cdot(\textbf{v}_{1}\wedge\textbf{v}_{2}...\wedge\textbf{v}_{k}):=g\textbf{v}_{1}\wedge g\textbf{v}_{2}\wedge\cdots\wedge g\textbf{v}_{k},~{}g\in\text{SL}(n,\mathbb{R}).$
For a fixed $k<n$, the action of $\text{SL}(n,\mathbb{R})$ is transitive on the space of $k$-lattices of rank $k<n$. We denote $\mathbb{Z}^{k}:=\text{Span}_{\mathbb{Z}}\{e_{1},...,e_{k}\}$, and we observe that for a unimodular $n$-lattice $L=g\mathbb{Z}^{n}\leq\mathbb{R}^{n}$, where $g\in\text{SL}(n,\mathbb{R})$, the collection of primitive $k$-sublattices of $L$ is given by
$g\gamma\mathbb{Z}^{k},\gamma\in\text{SL}(n,\mathbb{Z}).$
The following powerful theorem on quantitative non-divergence due to Kleinbock [Kleinbockclay] will be needed.
Theorem 4.2.
Let $B\subseteq\mathbb{R}$ be an interval, let $C\in\mathbb{R},\alpha>0$, $10<\rho<$ 1 and let $h:B\to\text{SL}(n,\mathbb{R})$ be a continuous map. Assume that for every $\gamma\in\text{SL}(n,\mathbb{Z})$, and for every $1\leq k<n$ both of the following conditions are satisfied
  1. (1)
    the function $x\to\|h(x)\gamma\cdot e_{1}\wedge\cdots\wedge e_{k}\|$ is $(C,\alpha)$-good on $B$, and
  2. (2)
    $\sup_{x\in B}\|h(x)\gamma\cdot e_{1}\wedge\cdots\wedge e_{k}\|\geq\rho^{k}$.
Then, for every positive number $\rho$ such that $\epsilon<\rho$, we have
(4.2) |$|\{x\in B:\lambda_{1}(h(x)\mathbb{Z}^{n})\leq\epsilon\}|\leq Cn2^{n}\left(\frac{\epsilon}{\rho}\right)^{\alpha}|B|,$
where $\lambda_{1}(\cdot)$ is the function that outputs the length of the shortest nonzero vector of an Euclidean lattice.
Proof of Proposition 4.1.
We identify SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ with the space of unimodular lattices
$\mathcal{L}_{n}:=\{L:=\text{Span}_{\mathbb{Z}}\{v_{1},...,v_{n}\}:\det(v_{i,j})=1\}.$
Consider
$\mathcal{B}_{\delta}:=\{L\in\mathcal{L}_{n}:(L\setminus\{0\})\cap B_{\delta}(0)\neq\emptyset\}.$
Here $B_{\delta}(0)\subseteq\mathbb{R}^{n}$ is the ball of radius $\delta$ centered at the origin. By Mahler’s criterion, $\mathcal{L}_{n}\setminus\mathcal{B}_{\delta}$ is compact for all $\delta>0$ and thus $\mathcal{B}_{\delta}$ is a neighborhood of $\infty$. Fix $g\mathbb{Z}^{n}\in\mathcal{L}_{n}$ for $g\in\text{SL}(n,\mathbb{R})$, and consider the measures
[…]
Let $\mu$ be a weak-* limit of the measures $\mu_{T}$ as $T\to\infty$. In order to show that $\mu(\infty)=0$, it is enough to prove that for every $\epsilon>0$, there is $\delta$ such that
(4.4) $\limsup_{T\to\infty}\frac{|\{t\in[0,T]:\lambda_{1}(\varphi(t)g\mathbb{Z}^{n})\leq\delta\}|}{T}\leq\epsilon.$
This will be concluded by Theorem 4.2 as follows. We first verify the conditions. For $k\in\{1,2,..,n-1\}$, consider the representation (4.1). By Proposition 1.13, there exists $T_{0},C,\alpha>0$ such that for any fixed $k\in\{1,2,..,n-1\}$, and $\gamma\in\text{SL}(n,\mathbb{Z})$ it holds that
(4.5) $\Theta(t)=\|\varphi(t)g\gamma\cdot(e_{1}\wedge\cdots\wedge e_{k})\|$
is $(C,\alpha)$-good in $[T_{0},\infty)$. Let $F\subseteq\text{SL}(n,\mathbb{R})$ be a given compact subset. Then, the number $\rho$ defined by
$\rho:=(\inf\{\|\varphi(T_{0})g\gamma\cdot(e_{1}\wedge\cdots\wedge e_{k})\|:g\in F,\,\gamma\in\text{SL}(n,\mathbb{Z})\})^{\frac{1}{k}}$
is positive. Thus, by Theorem 4.2, for any $g\in F$,
(4.6) |$|\{t\in[T_{0},T]:\lambda_{1}(\varphi(t)g\mathbb{Z}^{n})\leq\delta\}|\leq C\left(\frac{\delta}{\rho}\right)^{\alpha}(T-T_{0}).$
4.1. Non-divergence for homogeneous space of Lie groups
For the following statement, we recall that for a Lie group $G$, $G_{u}$ denotes the Lie-subgroup generated by the one-parameter unipotent subgroups of $G$.
Proposition 4.3.
Let $G$ be a Lie subgroup of $\text{SL}(n,\mathbb{R})$ and let $\Gamma$ be a lattice in $G$. Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting curve. Assume further that $\varphi([0,\infty))\subseteq G_{u}$. Then, there exists a $T_{0}>0$ such that given $\epsilon>0$ and a compact set $F\subseteq G$, there exists a compact set $K\subseteq G/\Gamma$ such that for all $g\in F$ and all $T>T_{0}$,
(4.7) |$\lvert\{t\in[T_{0},T]:\varphi(t)g\Gamma\in K\}\rvert\geq(1-\epsilon)(T-T_{0}).$
Remark 4.4.
We note that the assumption that $\varphi([0,\infty))\subseteq G_{u}$ is non restrictive for our purposes in view of the following. In Proposition 5.13, we show that the hull of a non-contracting curve, definable in a polynomially bounded o-minimal structure and taking values in a Lie subgroup $G\subseteq\text{SL}(n,\mathbb{R})$, is contained in $G_{u}$. Thus, by multiplying the original curve with a correcting curve, the assumption $\varphi([0,\infty))\subseteq G_{u}$ will be satisfied.
Proof.
We follow the arguments as in [Dani1993LimitDO, Theorem 6.1] to reduce the problem to the case when $G$ is a semisimple group and $\Gamma$ is an irreducible lattice in $G$. We note that if $G/\Gamma$ is compact, there is nothing to prove. So, we now assume that $G/\Gamma$ is not compact.
Let $M$ be the smallest closed normal subgroup of $G$ such that $\bar{G}=G/M$ is semisimple with the trivial center and no compact factors. Then $M/(M\cap\Gamma)$ is compact, see [RA72, Theorem 8.24] and [Shah1996LimitDO, Proof of Theorem 2.2]. Let $q:G\to\bar{G}$ denote the quotient homomorphism. Then $q(\Gamma)$ is a lattice in $\bar{G}$, and the natural quotient map $\bar{q}:G/\Gamma\to\bar{G}/q(\Gamma)$ is proper. We note that if $\tilde{G}$ and $\tilde{M}$ denote the Zariski closures of $G$ and $M$ in $\text{SL}(n,\mathbb{R})$, respectively, then $\bar{G}=G/M=\tilde{G}^{0}/\tilde{M}^{0}$. Therefore, $q|_{G_{u}}:G_{u}\to\bar{G}$ is a rational homomorphism of real algebraic groups. Thus $q\circ\varphi$ is a curve definable in $\mathscr{S}_{\mathrm{pbd}}$. By Lemma 2.17, $q\circ\varphi:[0,\infty)\to\bar{G}$ is a non-contracting. Since $\bar{q}$ is proper, proving the theorem for $\bar{G}$ in place of $G$ is sufficient.
Therefore, we assume that $G$ is semisimple with the trivial center and no compact factors. Then, by [RA72, Theorem 5.22], there exist closed normal subgroups $G_{1},\ldots,G_{k}$ of $G$ for some $k$ such that $G$ equals the direct product $G_{1}\times\cdots\times G_{k}$, such that $\Gamma_{i}:=G_{i}\cap\Gamma$ is an irreducible lattice in $G_{i}$, for all $i$. Moreover, $\Gamma_{1}\times\cdots\times\Gamma_{k}$ is a normal subgroup of finite index in $\Gamma$. Hence, $G/\Gamma$ is finitely covered by $G_{1}/\Gamma_{1}\times\cdots\times G_{k}/\Gamma_{k}$. Let $q_{i}:G\to G_{i}$ denote the natural factor map, and observe that $q_{i}\circ\varphi$ is a non-contracting curve in $G_{i}$ definable in $\mathscr{S}_{\mathrm{pbd}}$ by Lemma 2.17. So, proving the theorem for $G_{i}/\Gamma_{i}$ separately for each $i$ will imply the theorem for $G/\Gamma$. Therefore, without loss of generality, we may assume that $\Gamma$ is an irreducible lattice in $G$, which is semisimple with the trivial center and no compact factors.
First, suppose that the real rank of $G$ is $1$. It is straightforward to adapt the proof of [Dani:rank-1, Proposition 1.2] to conclude (4.7). For this purpose, it is enough to use the following property of the map $\varphi$ in place of [Dani:rank-1, Lemmas 2.5 and 2.7]: Let $V=\wedge^{d}\mathfrak{g}$, where $d$ is the dimension of the maximal unipotent subgroup of $G$ and $\mathfrak{g}$ the Lie algebra of $G$. Now, $G$ acts on $V$ via $\wedge^{d}\text{Ad}$. Then there exists $T>0$, $C>0$ and $\alpha>0$ such that for any $g\in G$, the map $t\mapsto\lVert\varphi(t)g\cdot p\rVert$ has the $(C,\alpha)$-good property on $[T_{0},\infty)$ by Proposition 1.13.
Now we suppose that the real rank of $G$ is at least $2$. Then, by the arithmeticity theorem due to G. A. Margulis, $\Gamma$ is arithmetic; that is, see [Zimmer:book84, page 3]: There exists an $m\in\mathbb{N}$, an algebraic semisimple group $H\subseteq\text{SL}(m,\mathbb{R})$ defined over $\mathbb{Q}$ and a surjective algebraic homomorphism
$\rho:H^{0}\to G,$
such that $\ker\rho$ is compact and $\rho(H^{0}(\mathbb{Z}))$ is commensurable with $\Gamma$, where $H^{0}(\mathbb{Z})=H^{0}\cap\text{SL}(m,\mathbb{Z})$. Since $G$ has no compact factors, the map $\rho$ restricted to $H_{u}$ is a finite cover of $G$. By Lemma 2.21, we can lift $\varphi$ to a $\mathscr{S}_{\mathrm{pbd}}$-definable curve $\psi:[0,\infty)\to H_{u}\subseteq H^{0}\subseteq\text{SL}(m,\mathbb{R})$ which is non-contracting in $H_{u}$, such that $\rho\circ\psi=\varphi$. By Proposition 2.22, $\psi$ is non-contracting in $\text{SL}(m,\mathbb{R})$. Denote $Y:=\text{SL}(m,\mathbb{R})/\text{SL}(m,\mathbb{Z})$. As in Proposition 4.1, let $T_{0}>0$ be such that for any compact set $F_{1}\subseteq\text{SL}(m,\mathbb{R})$, there exists a compact set $K_{1}\subseteq Y$ such that for any $T\geq T_{0}$ and $h\in F_{1}$
(4.8) |$\lvert\{t\in[T_{0},T]:\psi(t)h\text{SL}(m,\mathbb{Z})\in K_{1}\}\rvert\geq(1-\epsilon)(T-T_{0}).$
Therefore, since $H^{0}/H^{0}(\mathbb{Z})\hookrightarrow Y$ is a proper $H^{0}$-equivariant continuous injection, the conclusion of the proposition holds for $H^{0}$ in place of $G$, $H^{0}(\mathbb{Z})$ in place of $\Gamma$, and $\psi$ in place of $\varphi$. And, since $\rho\circ\psi=\varphi$, the conclusion of the proposition holds for $\rho(H^{0}(\mathbb{Z}))$ in place of $\Gamma$, and hence it holds for $\Gamma$ due to the commensurability. ∎
As an immediate consequence of Proposition 4.3 we obtain the following.
Theorem 4.5.
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Let $G\subseteq\text{SL}(n,\mathbb{R})$ be a Lie subgroup, $\Gamma$ be a lattice in $G$, and suppose that $\varphi([0,\infty))\subseteq G_{u}$. Then, any weak-* limit $\mu$ of $\mu_{T,\varphi,x}$ as $T\to\infty$ is a probability measure on $G/\Gamma$, for all $x\in G/\Gamma$.
5. Unipotent invariance
We say that a measure $\mu$ on $G/\Gamma$ is invariant under $g\in G$ if $g_{*}\mu=\mu$, where $g_{*}$ denotes the pushforward by the left-translation map by $g$ on $G/\Gamma$. We say that a measure is invariant under a subgroup, if it is invariant under every element of the subgroup.
For the rest of the section, $\mathscr{S}_{\mathrm{pbd}}$ is a fixed polynomially bounded o-minimal structure.
Proposition 5.1.
Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup. Consider a continuous, unbounded, $\mathscr{S}_{\mathrm{pbd}}$-definable curve $\varphi:[0,\infty)\to G$ such that $\det(\varphi(t))$ is constant in $t$. Suppose further that for all $v\in\mathbb{R}^{n}\setminus\{0\}$, $\lim_{t\to\infty}\varphi(t)v\neq 0$. Then there exists a nontrivial one-parameter unipotent group $P\subseteq G$ such that if $\Gamma\subseteq G$ is a discrete subgroup, $g\in G$ and $\mu$ is weak-$\ast$ limit of the measures
[…]
then $\mu$ is invariant under $P$.
The key idea behind the proof is that o-minimal curves exhibit a kind of “tangency at infinity.” Specifically, one can construct definable change-of-speed functions $h(t)$ that diverge to infinity and such that
$\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}=\rho,$
where $\rho$ is a nontrivial element. Under our assumptions, these limiting elements $\rho$ generate a one-parameter unipotent subgroup, and moreover, it holds that $h^{\prime}(t)=1+o(t^{-\nu})$. The latter derivative asymptotic ensures that any limiting measure of $\mu_{T,\varphi,g\Gamma}$ as $T\to\infty$ is invariant under the subgroup generated by $\rho$; see Lemma 5.6 below.
This idea was previously applied by Shah in [Shah1994LimitDO] in the context of averages along polynomial curves. In the polynomial case, one can construct suitable change-of-speed maps using Taylor expansions. In the o-minimal setting, Peterzil and Steinhorn [PS99] showed that the collection of all possible limits of the form $\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}$, where $h$ is a definable function diverging to $\infty$, forms a one-dimensional torsion-free group. We refer to this group as the Peterzil-Steinhorn subgroup, which we define in a simplified form in Definition 5.3 to suit our needs.
Poulios, in his thesis [Poulios_thes], further studied Peterzil-Steinhorn groups and showed that the group associated with an unbounded curve definable in a polynomially bounded structure is either unipotent or $\mathbb{R}$-diagonalizable. Importantly, Poulios establishes a convenient criterion for when the Peterzil-Steinhorn group is unipotent.
We now recall the required notions and results from [Poulios_thes]. For the following, for $r<1$, consider
(5.1) $h_{r,s}(t):=\left(t^{1-r}+(1-r)s\right)^{\frac{1}{1-r}}=t+st^{r}+o(t^{r}),\text{ where }s\in\mathbb{R},$
and let
(5.2) $h_{1,s}(t):=st,\text{ where }s>0.$
Proposition 5.2.
[Poulios_thes, Theorem 3.2.4, Theorem 3.3.1, Corollary 3.4.3] Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup, and let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded curve. Then, there exists a unique $r\leq 1$ such that
(5.3) $M_{\varphi}:=\lim_{t\to\infty}t^{r}\cdot\varphi^{\prime}(t)\cdot\varphi(t)^{-1}$
with $M_{\varphi}\in\mathfrak{g}\neq\{0\}$, where $\mathfrak{g}$ represents the Lie-algebra of $G$.
It holds that $M_{\varphi}$ is nilpotent $\iff$ $r<1$, and $M_{\varphi}$ is $\mathbb{R}$-diagonalizable $\iff r=1$. Moreover:
  1. (1)
    If $r<1$, consider $\rho(s):=\exp(sM_{\varphi}),~{}s\in\mathbb{R}$. Then:
    (5.4) $\rho(s):=\lim_{t\to\infty}\varphi(h_{r,s}(t))\varphi(t)^{-1},\ s\in\mathbb{R}.$
  2. (2)
    If $r=1$, let $d(s):=\exp(\log(s)M_{\varphi})$, then:
    $d(s):=\lim_{t\to\infty}\varphi(h_{1,s}(t))\varphi(t)^{-1},\ s>0.$
Definition 5.3.
For an unbounded curve $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ definable in a polynomially bounded o-minimal structure, we say that $r\leq 1$ is the P.S. order of $\varphi$, if $r$ satisfies (5.3). We call the one-parameter subgroup $P$ generated by $\exp(M_{\varphi})$ as the P.S. group of $\varphi$.
Remark 5.4.
In the notations of Proposition 5.2, if $\varphi$ is a restriction of a one-parameter group to $[0,\infty)$, then $r=0$, $h_{r,s}(t)=t+s$, and $\rho(s)=\varphi(s)$ for all $s\in\mathbb{R}$; that is, $\varphi$ is a restriction of a unipotent one-parameter subgroup.
5.1. Unipotent P.S. groups and limiting unipotent invariance
We now show that when the P.S. group is unipotent, any limiting measure of measures of the form (1.1) will be invariant under the P.S. group. This proof is found in [Shah1994LimitDO] and in [Peterzil2018OminimalFO], and we give it here also for completeness. The following elementary calculus lemma will be needed.
Lemma 5.5.
Let $h:(0,\infty)\to\mathbb{R}$ be a differentiable function such that $\lim_{t\to\infty}h^{\prime}(t)=1$. Then, for any bounded continuous function $F:[0,\infty)\to\mathbb{R}$, we have
[…]
For the proof below, notice that for $r<1$, it holds that $h^{\prime}_{s,r}(t)=1+O(t^{r-1})$.
Lemma 5.6.
Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup and let $\Gamma\subseteq G$ be a discrete subgroup. Suppose that $\varphi:[0,\infty)\to G$ is a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded curve such that its P.S. group is unipotent (namely, (5.3) holds for $r<1$). Let $\rho(s)=\exp(sM_{\varphi})$, as in (5.4). Then
[…]
for all $f\in C_{c}(G/\Gamma)$, $x\in G/\Gamma$ and $s\in\mathbb{R}$.
Proof.
For $f\in C_{c}(G/\Gamma)$ and an $\epsilon>0$, by uniform continuity of $f$, there is an $T_{\epsilon}$ such for all $t\geq T_{\epsilon}$:
$|f(\rho(s)\varphi(t)x)-f(\varphi(h_{r,s}(t)x)|$ $=|f(\rho(s)\varphi(t)x)-f([\varphi(h_{r,s}(t))\varphi(t)^{-1}]\varphi(t)x)|$
$\leq\epsilon/2,$
Then,
[…]
$=O\left(\frac{T_{\epsilon}}{T}\right)+O(\epsilon)+\frac{1}{T}\int_{T_{\epsilon}}^{T}\left(f(\varphi(h_{r,s}(t)x)-f(\varphi(t)x)\right)dt.$
The last term converges to zero by Lemma 5.5. ∎
If a curve is contained in a unipotent group, as in [Peterzil2018OminimalFO], then the P.S. group is automatically unipotent. In an overview, to treat our more general setup, we will distinguish between two type of curves to those that are essentially diagonal (see the definition below), and those that are or not. We will show that if a curve is not essentially diagonal, then its P.S. group is unipotent.
Definition 5.7.
We call a curve $\{\varphi(t)\}_{t\geq 0}\subseteq\text{GL}(n,\mathbb{R})$ essentially diagonal if there exists a decomposition
(5.7) $\varphi(t)=\sigma(t)b(t)C,$
where $b(t)\in\text{GL}(n,\mathbb{R})$ is a diagonal matrix for all large $t$, $C\in\text{GL}(n,\mathbb{R})$ is a constant matrix, and $\sigma(t)\in\text{SL}(n,\mathbb{R})$ is convergent as $t\to\infty$.
The following is our key observation.
Proposition 5.8.
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded, continuous curve. Suppose that $\det(\varphi(t))$ is constant for all large enough $t$. Then, the P.S. group of $\varphi$ is unipotent if and only if the curve $\varphi$ is not essentially diagonal.
We will prove the above Proposition 5.8 in the subsection below. Before we proceed, we consider the following lemma which proves Proposition 5.1 by assuming Proposition 5.8 in combination with Lemma 5.6.
Lemma 5.9.
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded, continuous curve. Suppose that the determinant $\det(\varphi(t))$ is constant in $t$ and suppose that $\lim_{t\to\infty}\varphi(t)v\neq 0,$ for all $v\in\mathbb{R}^{n}\setminus\{0\}$. Then, $\varphi$ is not essentially diagonal and in particular, the P.S. group of $\varphi$ is unipotent.
Proof.
Suppose for contradiction that the P.S. group is not unipotent. Then, by Proposition 5.8, $\varphi$ is essentially diagonal. Namely $\varphi(t)=\sigma(t)b(t)C$, where $C$, $\sigma(t)\in\text{SL}(n,\mathbb{R})$ are bounded and $b(t)$ is a diagonal matrix with $\det(b(t))=c_{0}$ for all $t$. Since we assume that $\varphi$ is unbounded, it follows that $b(t)$ is unbounded, which in turn implies that there is an index $1\leq i\leq n$ such that the corresponding entry function $b_{ii}(t)$ on the diagonal of $b(t)$ decays to zero. In particular, for $v:=C^{-1}e_{i}$ it holds that $\lim_{t\to\infty}\varphi(t)v=0$, which is a contradiction. ∎
5.1.1. Proving Proposition 5.8
We will call the $(i,j)$ index of an upper-triangular matrix $b\in\text{SL}(n,\mathbb{R})$ of the form
(5.8) $b=\begin{bmatrix}f_{1,1}&0&\cdots&0&0&\cdots&0&0&\cdots&0\\ 0&\ddots&\ddots&\vdots&\vdots&&\vdots&\vdots&&\vdots\\ \vdots&\ddots&\ddots&0&0&\cdots&0&0&\cdots&0\\ 0&\cdots&0&f_{i,i}&0&\cdots&0&\boldsymbol{f_{i,j}}&\cdots&f_{i,n}\\ \vdots&&\vdots&\ddots&\ddots&&&\vdots&&\vdots\\ \vdots&&\vdots&&\ddots&\ddots&&\vdots&&\vdots\\ \vdots&&\vdots&&&\ddots&\ddots&\vdots&&\vdots\\ \vdots&&\vdots&&&&0&f_{j,j}&\cdots&f_{j,n}\\ \vdots&&\vdots&&&&&\ddots&\ddots&\vdots\\ 0&\cdots&0&\cdots&\cdots&\cdots&\cdots&\cdots&0&f_{n,n}\end{bmatrix},$
where $f_{i,j}\neq 0$, the first nonzero off-diagonal entry. More precisely, $(i,j)$ is the first nonzero entry among the off-diagonal entries according to the following lexicographic order on $\mathbb{N}^{2}$:
(5.9) $(i,j)\prec(k,l)\iff i<k\text{ or }(i=k\text{ and }j<l).$
Let $B\leq\text{SL}(n,\mathbb{R})$ be the subgroup of upper-triangular matrices. Let $b:[0,\infty)\to B$ be a definable curve. Recall that each definable function $f(t)$ is either zero for all large enough $t$ or $|f(t)|>0$ for all large $t$. Thus, for all large enough $t$, there exists a unique first nonzero off-diagonal entry in $b(t)$, or $b(t)$ is diagonal for all large $t$. We will refer to this entry as the first nonzero off-diagonal entry of the curve $b(t)$.
Lemma 5.10.
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable curve. Then there exists a definable curve $b:[0,\infty)\to B$ and a constant matrix $C\in\text{SL}(n,\mathbb{R})$ such that
$\varphi(t)=\sigma(t)b(t)C,$
where $\sigma(t)\in\text{SL}(n,\mathbb{R})$ converges as $t\to\infty$. Furthermore, for sufficiently large $t$, the curve $b(t)$ satisfies one of the following:
  • $b(t)$ is diagonal, or
  • the first nonzero off-diagonal entry $f_{i,j}$ of $b(t)$ satisfies:
    1. (1)
      $\deg f_{i,i}\neq\deg f_{i,j}$, and
    2. (2)
      $\deg f_{i,j}>\deg f_{j,j}$.
Proof.
Using the KAN decomposition, we first write $\varphi(t)\in\text{SL}(n,\mathbb{R})$, as $\varphi(t)=k(t)p(t)$, where $k(t)\in\text{SO}(n,\mathbb{R})$ and $p(t)\in B$ an upper-triangular matrix. In view of the Gram-Schimdt process, $k(t)$ and $p(t)$ are definable. Since $k(t)$ is bounded and definable, $\lim_{t\to\infty}k(t)$ exists. For all $t$ large, $p(t)$ is either diagonal or $p(t)$ takes the form of (5.8).
We employ the following algorithm to achieve the outcome described in the lemma. At each step, we perform either a column or a row operation on $p(t)$:
  1. (1)
    All off-diagonal entries are eventually 00. Then we are done.
  2. (2)
    We pick the first index $(i,j)$ (with $i<j$) such that the $(i,j)$-th entry is not eventually 00, and proceed to the next step.
  3. (3)
    Suppose that $\deg(f_{i,j})=\deg(f_{i,i})$. Then there is an $c\in\mathbb{R}$ such that $\deg(f_{i,j}-cf_{i,i})<\deg(f_{i,i})$. For this step, subtract from the $j$-th column the $i$-th column multiplied by $c$. This amounts to multiplying $p(t)$ by a constant unipotent matrix from the right. The obtained matrix is the same besides the $i,j$-th entry which is replaced with $f_{i,j}-cf_{i,i}$. There are two possibilities now:
    1. (a)
      $f_{i,i}-cf_{i,j}$ is eventually zero. If all off-diagonal entries are eventually 00, then we are done. Otherwise, the first eventually nonzero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
    2. (b)
      $f_{i,i}-cf_{i,j}$ is not eventually zero: Then the first requirement of Lemma 5.10 is satisfied. We continue then with the following step.
  4. (4)
    If $\deg(f_{i,j})>\deg(f_{j,j})$, then the second requirement of Lemma 5.10 is satisfied, and we are done, otherwise we proceed to the next step.
  5. (5)
    Now suppose $\deg(f_{i,j})\leq\deg(f_{j,j})$. Then subtract from the $i$-th row the $j$-th row multiplied by $\frac{f_{i,j}}{f_{j,j}}$. This amounts to multiplying from the left by a unipotent matrix, which converges as $t\to\infty$. This is so because
    $\deg\left(\frac{f_{i,j}}{f_{j,j}}\right)=\deg(f_{i,j})-\deg(f_{j,j})\leq 0.$
    Then, either all the off-diagonal entries in the resulting matrix are eventually 00, and we are done; or the first eventually nonzero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
The algorithm ends with finitely many steps with either a diagonal matrix or a definable curve $b(t)$ satisfying the requirements of the lemma. ∎
Proof for Proposition 5.8.
First, note that it is enough to prove the statement for $\varphi\subseteq\text{SL}(n,\mathbb{R})$ since the P.S. group is unchanged if we multiply the curve by a constant invertible matrix from the right. In the notations of Lemma 5.10:
$\varphi(t)=\sigma(t)b(t)C,$
where $\lim_{t\to\infty}\sigma(t)=g\in\text{SL}(n,\mathbb{R}).$ Let $r\leq 1$ be such that (5.3) holds for $\{b(t)\}$. Let $h_{s,r}(t)$ be is as in (5.1)–(5.2). We note that
(5.10) $\lim_{t\to\infty}\varphi(h_{s,r}(t))\varphi(t)^{-1}=g\left[\lim_{t\to\infty}b(h_{s,r}(t))b(t)^{-1}\right]g^{-1}.$
Therefore, by Proposition 5.2, the $r$ corresponding to (5.3) for $\varphi(t)$ and $b(t)$ are the same. Now, $b(t)$ is either eventually diagonal or upper-triangular and satisfies the conditions of the Lemma 5.10 for the first (eventually) nonzero off-diagonal entry. If $b(t)$ is eventually diagonal, then for all $s>0$, $\lim_{t\to\infty}b(h_{1,s}(t))b^{-1}(t)$ converges to a diagonal matrix. Namely, the P.S. group is diagonal in this case.
Suppose that $b(t)$ is not eventually diagonal. Let $(i,j)$ be the first nonzero off-diagonal entry in $b(t)$ (for all large enough $t$). We observe that $(i,j)$-th entry in the matrix $b^{\prime}(t)b(t)^{-1}$ is
(5.11) $\frac{-f_{i,i}^{\prime}f_{i,j}+f_{i,j}^{\prime}f_{i,i}}{f_{i,i}f_{j,j}}=\left(\frac{f_{i,j}}{f_{i,i}}\right)^{\prime}\cdot\frac{f_{i,i}}{f_{j,j}}.$
Since $r_{i,i}:=\deg f_{i,i}\neq r_{i,j}:=\deg f_{i,j}$, we have that
$\deg\left(\frac{f_{i,j}}{f_{i,i}}\right)^{\prime}={r_{i,j}-r_{i,i}-1}.$
Thus,
$\deg\left(\left(\frac{f_{i,j}}{f_{i,i}}\right)^{\prime}\cdot\frac{f_{i,i}}{f_{j,j}}\right)=r_{i,j}-r_{j,j}-1.$
Recall by Lemma 5.10 that $r_{i,j}-r_{j,j}>0$. Thus the $(i,j)$-th entry in $tb^{\prime}(t)b(t)^{-1}$ is unbounded. Hence $r\neq 1$. ∎
5.2. Unipotent invariance in quotients
For our purposes, it will not suffice only to establish unipotent invariance for our limiting measures, but we will also require unipotent invariance in certain quotients. We will now describe more precisely our motivation for the results in this section and how they play a role in the proof of Theorem 1.7 in Section 6.
Consider a non-contracting curve $\varphi:[0,\infty)\to G:=\mathbf{G}(\mathbb{R})$ definable in $\mathscr{S}_{\mathrm{pbd}}$, and consider the measures $\mu_{T,\varphi,x}$; see (1.1). By Theorem 4.5, there exists a subsequence $\mu_{T_{i},\varphi,x}$ which converges to a probability measure $\mu$, as $i\to\infty$. By Proposition 5.1, $\mu$ is invariant under a nontrivial unipotent one-parameter subgroup. Let $W$ be the subgroup generated by all unipotent one-parameter subgroups that preserve $\mu$. We employ Ratner’s measure classification theorem for the $W$-invariant ergodic measures and the linearization technique to prove that $\varphi(t)$ is contained in a certain algebraic subgroup $N$, such that $W\trianglelefteq N$ and such that $H_{\varphi}\subseteq N$, where $H_{\varphi}$ is the hull of $\varphi$. Our goal will be to prove that $H_{\varphi}\subseteq W$. To achieve this goal, we would like to show that if $\varphi(t)$ is not bounded modulo $W$, then the image of $\varphi$ in $N/W$, which is definable, is, in fact, non-contracting, and hence its corresponding P.S. group in $N/W$ is unipotent. The latter outcome implies that $\mu$ is invariant under a unipotent one-parameter subgroup of $N$, which is not contained in $W$. This contradicts our choice of $W$ as the group generated by all unipotents preserving $\mu$. Before proceeding with the relevant result, we have the following remark which emphasizes that the requirement that $H_{\varphi}\subseteq N$ is important.
Remark 5.11.
Consider $\varphi(t):=\begin{bmatrix}t&t^{2}\\ 0&1/t\end{bmatrix}$ for $t\geq 1$. Then $\varphi$ is non-contracting in $\text{SL}(2,\mathbb{R})$, and its P.S. group is $U:=\{u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}:s\in\mathbb{R}\}$. Now consider the group $B\leq\text{SL}(2,\mathbb{R})$ of upper triangular matrices. Then, $B/U$ is naturally identified with the group $A\leq\text{SL}(2,\mathbb{R})$ of diagonal matrices. In particular, $\varphi(t)$ modulo $U$ is unbounded, but it is a diagonal curve. Due to the next result, this is explained by the fact that $H_{\varphi}$ is not contained in $B$.
Proposition 5.12.
Let $\psi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be an unbounded, non-contracting curve definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\psi([0,\infty))\subseteq H_{\psi}$, where $H_{\psi}$ is the hull of $\psi$. Let $N\subseteq\text{SL}(n,\mathbb{R})$ be a Zariski closed subgroup such that $H_{\psi}\subseteq N$, and let $H\trianglelefteq N$ be a closed normal subgroup.
Assume that the image of $\psi$ in $N/H_{u}$ is unbounded. Let $q:N\to N/H$ be the natural quotient map. Then, there exists $r<1$ such that
(5.12) $\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1}=p(s),\,\forall s\in\mathbb{R},$
where $\{p(s):s\in\mathbb{R}\}$ is a nontrivial one-parameter subgroup of $N/H$. Here, $h_{s,r}(t)$ is given by (5.1). Importantly, there exists a one-parameter unipotent subgroup $\{u(s):s\in\mathbb{R}\}\subseteq N$ such that $p(s)=q(u(s))$ for all $s\in\mathbb{R}$.
Proof of Proposition 5.12.
Let $\tilde{H}_{u}:=\text{zcl}(H_{u})$, where $H_{u}$ is the subgroup generated by all one-parameter unipotent subgroups contained in $H$. Since $H\trianglelefteq N$, it follows that $H_{u}\trianglelefteq N$, and by Zariski density $\tilde{H}_{u}\trianglelefteq N$. Now, $H_{u}$ is of finite index in $\tilde{H}_{u}$ (see Remark 2.7), therefore, since the image of $\psi$ is unbounded in $N/H_{u}$ it follows that the image of $\psi$ is unbounded in $N/\tilde{H}_{u}$.
We realize $N/\tilde{H}_{u}$ as a real algebraic subgroup of $\text{SL}(m,\mathbb{R})$ for some $m$. For each $t\in[0,\infty)$, let $\bar{\psi}(t)$ denote the image of $\psi(t)$ in $\text{SL}(m,\mathbb{R})$. We claim that $\bar{\psi}$ is non-contracting for $\text{SL}(m,\mathbb{R})$. Indeed, because any finite-dimensional linear representation of $\text{SL}(m,\mathbb{R})$ is an algebraic representation of $\tilde{H}_{\psi}\subseteq N$, and $\psi$ is non-contracting for $\tilde{H}_{\psi}$, where $\tilde{H}_{\psi}$ is the Zariski closure of the hull of $\psi$, see Theorem 2.5, (3).
Therefore by Lemma 5.9, there exists $r<1$ such that
(5.13) $\lim_{t\to\infty}\bar{\psi}(h_{s,r}(t))\bar{\psi}(t)^{-1}=:\bar{p}(s),\,\forall s\in\mathbb{R},$
where $s\mapsto\overline{p}(s)$ is a nontrivial one-parameter unipotent subgroup of $\text{SL}(m,\mathbb{R})$ contained in $N/\tilde{H}_{u}$. Let $\vartheta:N\to N/H_{u}$ denote the quotient homomorphism. Since the quotient map $N/H_{u}\to N/\tilde{H}_{u}$ is a homomorphism with a finite kernel, for each $s\in\mathbb{R}$, the following limit exists:
(5.14) $\tilde{p}(s):=\lim_{t\to\infty}\vartheta(\psi(h_{s,r}(t)))\vartheta(\psi(t))^{-1}.$
In view of Jordan decomposition, there exists a unipotent one-parameter subgroup $\{u(s):s\in\mathbb{R}\}\subseteq N$ such that for all $s\in\mathbb{R}$, the image of $u(s)$ in $N/\tilde{H}_{u}$ is $\bar{p}(s)$, and hence $\tilde{p}(s)=\vartheta(u(s))$. Since $\{\bar{p}(s):s\in\mathbb{R}\}$ is a unipotent group, we obtain that $\{u(s):s\in\mathbb{R}\}$ is a nontrivial unipotent one-parameter subgroup of $N$ not contained in $H_{u}$, and hence it is not contained in $H$. Thus, $\{q(u(s)):s\in\mathbb{R}\}$ is a nontrivial subgroup of $N/H$.
Denote $\eta:N/H_{u}\to N/H$ the quotient homomorphism, and note that $\eta\circ\vartheta=q$. Then, for all $s\in\mathbb{R}$, $p(s):=\eta(\tilde{p}(s))=q(u(s))$, and by (5.14),
$p(s)=\lim_{t\to\infty}\eta\circ\vartheta(\psi(h_{s,r}(t))\psi(t)^{-1})=\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1}.$
5.3. Inclusion of the hull in a Lie subgroup
Proposition 5.13.
Let $G\subseteq\text{SL}(n,\mathbb{R})$ be a connected unimodular closed subgroup, and let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting, continuous curve such that $\varphi([0,\infty))\subseteq G$. Then, $H_{\varphi}\subseteq G_{u}$, and in particular, any correcting curve of $\varphi$ takes values in $G_{u}$.
The result does not follow from Theorem 2.5, because we do not assume that $G$ is an observable subgroup of $\text{SL}(n,\mathbb{R})$.
If $G$ is not algebraic, the hull $H_{\varphi}$ may be contained in the Zariski closure of $G$, but not in $G$. The unimodularity of $G$ ensures that $H_{\varphi}\subseteq G$.
The unimodularity assumption is essential even in the algebraic subgroup case. For example, in Remark 5.11, the curve is non-contracting in $\text{SL}(2,\mathbb{R})$ and its image is contained in the non-unimodular Borel subgroup $B$ of $\text{SL}(2,\mathbb{R})$. But its hull is not contained in $B$.
The following is an immediate corollary of Proposition 5.13.
Corollary 5.14.
Let $G_{1}\subseteq G_{2}$ be real points of real algebraic groups. Let $\varphi:[0,\infty)\to G_{2}$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting, continuous curve. Suppose that $\varphi([0,\infty))\subseteq G_{1}$ and $G_{1}$ is unimodular. Then $\varphi:[0,\infty)\to G_{1}$ is non-contracting.
The proof will be based on Proposition 5.12 and the following key lemma.
Lemma 5.15.
Let $\mathfrak{g}$ be a Lie subalgebra of $\mathfrak{gl}(n,R)$. Let $\mathfrak{g}_{u}$ be the subalgebra generated by nilpotent matrices in $\mathfrak{g}$. Then for any $X\in\mathfrak{g}$,
$\text{tr}(\text{ad}X)=\text{tr}(\text{ad}X|_{\mathfrak{g}_{u}}),$
where $\text{tr}(A)$ denotes the trace of $A$.
So, for the connected Lie subgroup $G$ corresponding to $\mathfrak{g}$, we have
(5.15) $\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=\det(\text{Ad}(g)),\ \forall g\in G.$
Proof.
We will be using repeatedly the following fact – If $T$ is a linear map on a vector space $V$ preserving a subspace $W$, and $T_{V/W}$ is the factored map on $V/W$, then
(5.16) $\text{tr}(T|_{V})=\text{tr}(T|_{W})+\text{tr}(T_{V/W}),$
where $T|_{W}$ is the restriction of $T$ to $W$.
[…]
(5.17) $\text{tr}(\text{ad}X|_{\mathfrak{g}})=\text{tr}(\text{ad}X|_{[\mathfrak{g},\mathfrak{g}]}).$
Let $\tilde{\mathfrak{g}}$ denote the Lie algebra of the Zariski closure of the Lie group associated with $\mathfrak{g}$ in $\text{GL}(n,\mathbb{R})$. By [Chevalley-LieII, Chapter II, Theorem 13],
$[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=[\mathfrak{g},\mathfrak{g}]\subseteq\mathfrak{g}.$
Let $\mathfrak{u}$ denote the radical of the $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$. Then $\mathfrak{u}$ consists of nilpotent matrices; we can derive this from [RA72, (2.5) A structure theorem]. So, $\mathfrak{u}\subseteq\mathfrak{g}_{u}$.
Since $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$ and $\mathfrak{u}$ are ideals in $\tilde{\mathfrak{g}}$ and $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]/\mathfrak{u}$ is semisimple, the trace of the derivation on it corresponding to $\text{ad}X$ is 00. Therefore,
(5.18) $\text{tr}(\text{ad}X|_{[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]})=\text{tr}(\text{ad}X|_{\mathfrak{u}}).$
Let $\mathfrak{v}$ denote the radical of $\mathfrak{g}_{u}$. Then, $\mathfrak{u}\subseteq\mathfrak{v}$. Since $\mathfrak{g}_{u}$ is a Lie subalgebra generated by nilpotent matrices, $\mathfrak{v}$ consists of nilpotent matrices. Also, $\mathfrak{g}_{u}$, and hence $\mathfrak{v}$ are ideals of $\mathfrak{g}$. So, by definition of $\tilde{\mathfrak{g}}$, $\mathfrak{v}$ is an ideal of $\tilde{\mathfrak{g}}$. So, $\mathfrak{v}\subseteq\mathfrak{u}$. Hence $\mathfrak{v}=\mathfrak{u}$.
As above, since $\mathfrak{g}_{u}/\mathfrak{v}$ is semisimple, we conclude that
(5.19) $\text{tr}(\text{ad}X|_{\mathfrak{g}_{u}})=\text{tr}(\text{ad}X|_{\mathfrak{v}}).$
Now, (5.16) follows from (5.17)-(5.19). ∎
Proof of Proposition 5.13.
Consider the exterior product of the Adjoint representation of $\text{SL}(n,\mathbb{R})$ on $\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{sl}(n,\mathbb{R})$, and fix $p\in\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{g}_{u}\setminus\{0\}$. Let
$N$ :=$:=\{g\in\text{SL}(n,\mathbb{R}):g\cdot p=p\}$
=$=\{g\in\text{SL}(n,\mathbb{R}):\text{Ad}(g)\mathfrak{g}_{u}=\mathfrak{g}_{u},\,\det(\text{Ad}(g)|_{\mathfrak{g_{u}}})=1\}$
So, $N$ normalizes $G_{u}$. Since $G$ is unimodular, by (5.15), $G\subseteq N$.
We have $\varphi:[0,\infty)\to G$. Let $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a correcting curve for $\varphi$, and denote $\psi(t):=\beta(t)\varphi(t)$. Now $N$ is observable in $\text{SL}(n,\mathbb{R})$ (Definition 2.2), and $\varphi([0,\infty))\subseteq G\subseteq N$. So $H_{\varphi}\subseteq N$ by the minimality of the hull, Theorem 2.5,(1). Now, by Remark 2.8, we have that $\psi([0,\infty))\subseteq H_{\psi}=H_{\varphi}$.
Assume for contradiction that $\psi$ is not bounded modulo $G_{u}$. Let $q:N\to N/G_{u}$ be the quotient map. Then, by Proposition 5.12, there exists a one-parameter unipotent subgroup of $\{u(s)\}_{s\in\mathbb{R}}\subseteq N$ such that
$q(u(s))=\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1},\,\forall s\in\mathbb{R},$
and $q\circ u$ is a nontrivial unipotent one-parameter subgroup of $N/G_{u}$. Let $\beta_{\infty}=\lim_{t\to\infty}\beta(t)$. Then, for all $s\in\mathbb{R}$,
$q(\beta_{\infty}^{-1}u(s)\beta_{\infty})=\lim_{t\to\infty}q(\varphi(h_{s,r}(t)))q(\varphi(t))^{-1},$
and since $\varphi([0,\infty))\subseteq G\subseteq N$, we get that $\{q(\beta_{\infty}^{-1}u(s)\beta_{\infty}):s\in\mathbb{R}\}$ is a nontrivial one-parameter subgroup of $G/G_{u}$. It then follows that $\{\beta_{\infty}^{-1}u(s)\beta_{\infty}:s\in\mathbb{R}\}$ is a nontrivial one-parameter unipotent subgroup contained in $G$ not contained in $G_{u}$, a contradiction. Finally, since $\varphi$ is bounded modulo $G_{u}$, and since $\text{zcl}(G_{u})$ is observable, we get that $H_{\varphi}\subseteq G_{u}$ by the minimality of the hull, Theorem 2.5,(1). ∎
6. Avoidance of singular sets
Suppose that $\Gamma\subseteq\text{SL}(n,\mathbb{R})$ is a discrete subgroup, and $\mu$ is a probability measure on $X:=\text{SL}(n,\mathbb{R})/\Gamma$ which is invariant under a nontrivial one-parameter unipotent subgroup of $\text{SL}(n,\mathbb{R})$. Then, by the celebrated Ratner’s theorem (see [Ratner91a]), every $W$-ergodic component of $\mu$ is homogeneous.
The measure $\mu$ that we will consider is a weak-* limit of the measures $\mu_{T,\varphi,x}$ (see (1.1)), for a non-contracting, definable polynomially bounded curve $\varphi$. By Theorem 4.5 and Proposition 5.1, $\mu$ is a probability measure, and it is invariant under a nontrivial unipotent subgroup.
As an immediate consequence of Ratner’s theorem (see Theorem 6.3) if $\mu$ is not $\text{SL}(n,\mathbb{R})$-invariant, then it is strictly positive on a Ratner’s singular set, which is the image of a certain strictly lower dimensional algebraic subvariety of $\text{SL}(n,\mathbb{R})$ on $X=\text{SL}(n,\mathbb{R})/\Gamma$. To analyze this situation, we will consider a finite-dimensional representation $V$ of $\text{SL}(n,\mathbb{R})$ and represent the singular set via an algebraic subvariety, say $\mathcal{A}$, in $V$. The trajectory $t\mapsto\varphi(t)x$ defining $\mu$ will be represented by countably many simultaneous trajectories $t\mapsto\varphi(t)\gamma\cdot p$ ($\forall\gamma\in\Gamma$) for a well chosen $p\in V$ with $\Gamma\cdot p$ discrete. We will show that $\mu$ cannot be positive on the singular set unless for some $\gamma\in\Gamma$, $\varphi(t)\gamma\cdot p$ converges to a point of $\mathcal{A}$ as $t\to\infty$. This leads to group-theoretic constraints.
Studying the dynamics of $\varphi(t)x$-trajectories in thin neighborhoods of a singular set in $X$ via simultaneously analyzing countably many trajectories of $\varphi$ in a thin neighborhood of an algebraic variety in a vector space is known as the linearization technique. Showing that an unbounded trajectory of $\varphi$ avoids remaining too close to an algebraic variety for too long is called the avoidance of singular sets - this is achieved via the $(C,\alpha)$-good property. The remainder of the section is devoted to employing these techniques in our situation.
6.1. Description of Ratner’s singular sets
Definition 6.1.
Let $\mathcal{H}$ be the class of all closed connected subgroups $H$ of $\text{SL}(n,\mathbb{R})$ such that $H/H\cap\Gamma$ admits an $H$-invariant probability measure and the closed subgroup $H_{u}\subseteq H$ which is generated by all unipotent one-parameter subgroups of $H$ acts ergodically on $H/H\cap\Gamma$ concerning the $H$-invariant probability measure.
Theorem 6.2 ([Ratner91a, Theorem 1.1]).
The collection $\mathcal{H}$ is countable.
We note that if $\Gamma=\text{SL}(n,\mathbb{Z})$, then $\text{zcl}(H\cap\Gamma)$ is defined over $\mathbb{Q}$ and $H=\text{zcl}(H\cap\Gamma)(\mathbb{R})^{0}$ [Shah1991, (2.3) and (3.2)] for all $H\in\mathcal{H}$, so $\mathcal{H}$ is countable.
Let $H\in\mathcal{H}$. Define
$N(H,W)=$$\{g\in\text{SL}(n,\mathbb{R}):W\subseteq gHg^{-1}\},\text{ and}$
$S(H,W)=$ $\bigcup_{H^{\prime}\in\mathcal{H},H^{\prime}\subsetneq H}N(H^{\prime},W).$
Note that $N(W)N(H,W)N(H)=N(H,W)$, where $N(F)$ denotes the normalizer of $F$ in $\text{SL}(n,\mathbb{R})$. Also, for all $\gamma\in\Gamma$, we have $\gamma H\gamma^{-1}\in\mathcal{H}$,
$N(H,W)\gamma=N(\gamma^{-1}H\gamma,W)\text{ and }S(H,W)\gamma=S(\gamma^{-1}H\gamma,W).$
Let $\pi:\text{SL}(n,\mathbb{R})\to X=\text{SL}(n,\mathbb{R})/\Gamma$ denote the quotient map. Consider the refined singular set
(6.1) $T_{H}(W)=\pi(N(H,W)\setminus S(H,W))=\pi(N(H,W))\setminus\pi(S(H,W)).$
for the second equality see [Mozes1995OnTS, Lemma 2.4]. Also, $T_{H}(W)=T_{\gamma H\gamma^{-1}}(W)$ for all $\gamma\in\Gamma$.
Theorem 6.3.
(Ratner [Ratner91a]) [Mozes1995OnTS, Theorem 2.2] Let $\nu$ be a $W$-invariant Borel probability measure on $X$. For each $H\in\mathcal{H}$, let $\nu_{H}$ denote the restriction of $\nu$ on $T_{H}(W)$, which is $W$-invariant. Then $\nu$ decomposes as:
$\nu=\sum_{H\in\mathcal{H}^{*}}\nu_{H},$
where $\mathcal{H}^{*}\subseteq\mathcal{H}$ is a countable set of subgroups, each of which is a representative of a distinct $\Gamma$-conjugacy class.
6.2. O-minimal curves in representations – a dichotomy theorem
For $H\in\mathcal{H}$, consider
(6.2) $V_{H}:=\bigwedge^{\dim(H)}\mathfrak{sl}(n,\mathbb{R})\text{ and }p_{H}\in\bigwedge^{\dim(H)}\mathfrak{h}\setminus\{0\},$
where $\mathfrak{h}$ denotes the Lie algebra of $H$. Consider the action of $\text{SL}(n,\mathbb{R})$ on $V_{H}$ induced by its Adjoint representation its Lie algebra; that is, $g\cdot(\bigwedge_{i=1}^{d}x_{i}):=\bigwedge_{i=1}^{d}\text{Ad}_{g}(x_{i})$ and extended linearly. We denote by $\eta_{H}:\text{SL}(n,\mathbb{R})\to V_{H}$ the orbit map
(6.3) $\eta_{H}(g)=g\cdot p_{H},\,\forall g\in\text{SL}(n,\mathbb{R}).$
The notation $\eta_{H}$ is used only for denoting preimages under the orbit map. We note some basic observations.
Theorem 6.4.
Let $H\in\mathcal{H}$. Then:
  1. (1)
    [Dani1993LimitDO, Proof of Proposition 3.2] Let $\mathfrak{w}$ be the Lie algebra of $W$, and consider the linear subspace
    (6.4) $A_{H}:=\{v\in V_{H}:v\wedge w=0,\forall w\in\mathfrak{w}\}.$
    Then $\eta_{H}^{-1}(A_{H})=N(H,W)$.
  2. (2)
    [Dani1993LimitDO, Theorem 3.4]
    $\Gamma\cdot p_{H}$ is closed, and hence discrete, in $V_{H}$.
For example, if $\Gamma=\text{SL}(n,\mathbb{Z})$, then $\mathfrak{h}$ is defined over $\mathbb{Q}$, so we can pick $p_{H}\in\wedge^{d}\mathfrak{h}(\mathbb{Z})\setminus\{0\}$, and hence, $\Gamma\cdot p_{H}\subseteq V_{H}(\mathbb{Z})$.
Next, we will provide a version of [Mozes1995OnTS, Proposition 3.4] for definable curves. Our proof is based on the linearization technique developed by Dani and Margulis [Dani1993LimitDO, Proof of Theorem 1]. The result can be considered an analogue of Theorem 4.2. In what follows, $\mathscr{S}_{\mathrm{pbd}}$ denotes a polynomially bounded o-minimal structure.
Theorem 6.5.
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Let $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a correcting curve of $\varphi$ and let $H_{\varphi}$ be the hull of $\varphi$ such that $\beta(t)\varphi(t)\subseteq H_{\varphi}$ for all $t$, see Definition 2.6. Fix $H\in\mathcal{H}$. Let $\emph{\text{K}}_{1}\subseteq A_{H}$ be a compact set. Then, given $x\in X$, one of the following holds:
  1. (1)
    There exists $w\in\pi^{-1}(x)\cdot p_{H}\cap\emph{\text{K}}_{1}$ such that
    (6.5) $H_{\varphi}\cdot w=\{w\}.$
  2. (2)
    Given $10<\epsilon<$ 1, there exists a closed set $\mathcal{S}\subseteq\pi(S(H,W))$ such that given a compact set $K\subseteq X\setminus\mathcal{S}$, there exist a compact neighborhood $\Psi$ of $\emph{\text{K}}_{1}$ in $V_{H}$ and a $T_{x}>0$ such that for all $T>T_{x}$,
    (6.6) |$|\{t\in[0,T]:\beta(t)\varphi(t)x\in K\cap\pi(\eta_{H}^{-1}(\Psi))\}|\leq\epsilon T.$
We now recall some observations and results that will be used in the proof. Let $N(H)$ be the normalizer of $H$ in $\text{SL}(n,\mathbb{R})$. Note that,
(6.7) $g\cdot p_{H}=\det(\text{Ad}_{g}|_{\mathfrak{h}})p_{H},\,\forall g\in N(H).$
Set $N_{\Gamma}:=N(H)\cap\Gamma.$ By [Dani1993LimitDO, Lemma 3.1], $N_{\Gamma}\subseteq\{\pm p_{H}\}$. If $N_{\Gamma}=\{\pm p_{H}\}$, let $\tilde{V}_{H}:=V_{H}/\{\pm 1\}$, the space $V_{H}$ modulo the equivalence relation defined by identifying $v\in V_{H}$ with $-v$. If $N_{\Gamma}\cdot p_{H}=\{p_{H}\}$, let $\tilde{V}_{H}:=V_{H}$.
For any $g\in\text{SL}(n,\mathbb{R})$ and $y=\pi(g)\in X$, we treat $\pi^{-1}(y)\cdot\tilde{p}_{H}=g\Gamma\cdot\tilde{p}_{H}$ as the discrete set of representatives of $y$ in $\tilde{V}_{H}$, where $\tilde{p}_{H}$ denotes the image of $p_{H}$ in $\tilde{V}_{H}$. We need the following result on the uniqueness of representatives in thin neighborhoods.
Proposition 6.6.
[Mozes1995OnTS, Proposition 3.2] Let $D$ be a compact subset of $A_{H}$, and consider
$S(D)=\{g\in\eta_{H}^{-1}(D):g\gamma\in\eta_{H}^{-1}(D)\text{ for some }\gamma\in\Gamma\setminus N_{\Gamma}\}\subseteq\text{SL}(n,\mathbb{R}).$
Then:
  1. (1)
    $S(D)\subseteq S(H,W)$.
  2. (2)
    $\pi(S(D))$ is closed in $X$.
  3. (3)
    Suppose that $K\subseteq X\setminus\pi(S(D))$ is compact. Then, there is a compact neighborhood $\Phi$ of $D$ in $V_{H}$ such that any $y\in K$ has at most one representative in $\tilde{\Phi}$, where $\tilde{{\Phi}}$ is the image of ${\Phi}$ in $\tilde{V}_{H}$.
    That is, for all $y\in\pi(\eta_{H}^{-1}({\Phi}))\cap K$, the set $\pi^{-1}(y).\tilde{p}_{H}\cap\tilde{{\Phi}}$ consists of a single element.
Proof of Theorem 6.5.
We assume for simplicity that $N_{\Gamma}\cdot p_{H}=\{p_{H}\}$, namely $\tilde{V}_{H}=V_{H}$. The case for which $N_{\Gamma}\cdot p_{H}=\{\pm p_{H}\}$ can be treated in essentially the same way, and is left to the readers.
By choosing a basis for $V_{H}$, we identify $V_{H}\cong\mathbb{R}^{m}$ for some $m$. We will denote by $\psi(t)\in\text{GL}(m,\mathbb{R})$ the action of $\beta(t)\varphi(t)$ on $V_{H}$. We will apply Proposition 3.12 for $\psi(t)$. Since $A_{H}$ is a linear subspace, $A_{H}=\{Q=0\}$ for a polynomial map $Q:V_{H}\to\mathbb{R}$, see Remark 3.14. Let $K_{1}\subseteq A_{H}$ and let $10<\epsilon<$ 1 be given. Let $\text{K}_{2}\subseteq A_{H}$ be a compact set that is given by Proposition 3.12 with $\text{K}_{1}\subseteq\text{K}_{2}$.
Choices of $\mathcal{S}$ and $\Phi$ for injective property
For applying Proposition 6.6, we use $\text{K}_{2}$ in place of $D$, and let $\mathcal{S}:=\pi(S(\text{K}_{2}))$. Let $K\subseteq X\setminus\mathcal{S}$ be a compact set, and let $\Phi\subseteq V_{H}$ be a compact neighborhood of $\text{K}_{2}$ which satisfies the outcome of Proposition 6.6,(3). As a consequence, for any $t\geq 0$, $w_{1},w_{2}\in\Gamma\cdot p_{H}$, and $g\in\text{SL}(n,\mathbb{R})$, if $g\Gamma\in K$ and $g\cdot w_{i}\in\Phi$ for $i=1,2$, then $w_{1}=w_{2}$; let us call this the injective property of $K$ in $\Phi$.
Choice of $\Psi$ using $(C,\alpha)$-good property
We take $T^{\prime}_{0}>0$ to satisfy the outcomes of Propositions 1.13 and 3.12 for $\psi$. By Proposition 3.12, we choose a compact neighborhood $\Psi$ of $\text{K}_{1}$ in $V_{H}$ such that $\Psi\subseteq\mathring{\Phi}$ and for any $v\in V$ and any closed interval $[a,b]\subseteq[T_{0}^{\prime},\infty)$, if $\psi([a,b])v\subseteq\Phi$ and $\psi(b)v\notin\mathring{\Phi}$, then
(6.8) |$\lvert\{s\in[a,b]:\psi(s)v\in\Psi\}\rvert\leq\epsilon(b-a).$
It may be noted that in view of (3.30) and (3.31), there exist $R_{2}>0$ and $\epsilon_{1}>0$ such that
(6.9) $\Psi\subseteq\{v\in V_{H}:\lVert v\rVert<R_{2}\text{ and }\lvert Q(v)\rvert<\epsilon_{1}\}\subseteq\mathring{\Phi}.$
Modifying $\Psi$ for dichotomy
Let $x\in X$ be given. Since $\pi^{-1}(x)\cdot p_{H}\cap\Psi$ is finite, by replacing $\Psi$ with a smaller compact neighborhood of $\text{K}_{1}$, we may assume that $\pi^{-1}(x)\cdot p_{H}\cap\Psi=\pi^{-1}(x)\cdot p_{H}\cap K_{1}$.
By Theorem 2.5,(3), the following two mutually exclusive cases exhaust all possibilities.
  • Case 1
    There exists $w\in\pi^{-1}(x)\cdot p_{H}\cap\Psi$ such that $H_{\varphi}\cdot w=\{w\}$.
  • Case 2
    For all $w\in\pi^{-1}(x)\cdot p_{H}\cap\Psi$, $\lim_{s\to\infty}\psi(s)w=\infty$.
If Case 1 holds, then by our choice of $\Psi$, the first conclusion of the theorem holds. Therefore, to complete the proof, we assume that Case 2 holds.
Consider
(6.10) $J=\{t\geq T^{\prime}_{0}:\beta(t)\varphi(t)x\in\pi(\eta^{-1}_{H}(\Psi))\cap K\}.$
Then $J$ is closed. Because $\pi(\eta^{-1}_{H}(\Psi))$ is closed in $X$. To show this, suppose $g_{i}\to g$ in $\text{SL}(n,\mathbb{R})$ and $\gamma_{i}\in\Gamma$ are such that $g_{i}\gamma_{i}\cdot p_{H}\in\Psi$. Since $\Psi$ is compact and $\Gamma\cdot p_{H}$ is discrete, we can pick $\gamma\in\Gamma$ such that $\gamma_{i}\cdot p_{H}=\gamma\cdot p_{H}$ for infinitely many $i$’s. Therefore $g\gamma\cdot p_{H}\in\Psi$.
Then, to prove (6.6), it is enough to show that there exists $T_{x}\geq T_{0}^{\prime}$ depending on $x$ such that
(6.11) $\lvert J\cap[T_{0}^{\prime},T]\rvert\leq 2\epsilon T,\ \forall\,T\geq T_{x}.$
Without loss of generality, we may assume that $J$ is unbounded.
Choice of a representative of $x$ in for every $t\in J$
For every $t\in J$, there exists a representative of $x$, say $w\in\pi^{-1}(x)\cdot p_{H}$, such that $\psi(t)w\in\Psi\subseteq\Phi$. Since $\beta(t)\varphi(t)x\in K$, by the injective property of $K$ in $\Phi$, $w$ is unique, and we denote it by $w_{t}$.
Let $t\in J$. We note that
$\lim_{s\to\infty}\psi(s)w_{t}=\infty,$
because otherwise $w_{t}$ is fixed by $H_{\varphi}$ due to (3) of Theorem 2.5, and since $\psi(t)\subseteq H_{\phi}$, we have $w_{t}\in\Psi\cap\pi^{-1}(x)\cdot p_{H}$, contradicting the Case 2.
Therefore, $\{s\geq T_{0}^{\prime}:\psi(s)w_{t}\in\mathring{\Phi}\}$ is a bounded open subset of $[T_{0}^{\prime},\infty)$ containing $t$. Let $U(t)$ be its connected component containing $t$, and denote
$t^{+}:=\sup U(t)<\infty.$
Then
(6.12) $\psi([t,t^{+}))w_{t}\subseteq\mathring{\Phi}\text{ and }\psi(t^{+})w_{t}\in\Phi\setminus\mathring{\Phi}.$
Bound on $|J\cap[t,t^{+}]|$
For any $s\in J\cap[t,t^{+}]$, we have $\beta(s)\varphi(s)x\in K$, $\beta(s)\varphi(s)\cdot w_{s}\in\Psi\subseteq\mathring{\Phi}$, and $\beta(s)\varphi(s)\cdot w_{t}\in\Phi$, and hence by the injectivity of $K$ in $\Phi$, $w_{s}=w_{t}$. Then,
(6.13) $\begin{array}[]{ll}J\cap[t,t^{+}]&=\{s\in[t,t^{+}]\cap J:\psi(s)w_{s}\in\Psi\}\\ &=\{s\in[t,t^{+}]\cap J:\psi(s)w_{t}\in\Psi\}.\end{array}$
By (6.8), (6.12), and (6.13),
(6.14) $\lvert J\cap[t,t^{+}]\rvert\leq\left|\{s\in[t,t^{+}]:\psi(s)w_{t}\in\Psi\}\right|\leq\epsilon(t^{+}-t).$
Choosing a sequence $t_{i}\in J$
For each $i\in\mathbb{N}$, we define $t_{i}\in J$ inductively as follows: Let $t_{1}=\inf J\in J$. Let $i\in\mathbb{N}$ be such that $t_{i}$ is defined. Then we let $t_{i+1}=\inf J\cap(t_{i}^{+},\infty)$. Since $J$ is closed, $t_{i}\in J$.
By our construction, $t_{i}<t_{i}^{+}<t_{i+1}$ for all $i\in\mathbb{N}$. Let $i\in\mathbb{N}$. Since
$J\cap[T_{0}^{\prime},t_{i}^{+}]=\cup_{j=1}^{i}J\cap[t_{j},t_{j}^{+}),$
by (6.14),
(6.15) $\lvert J\cap[T_{0}^{\prime},t_{i}^{+}]\rvert\leq\epsilon(t_{i}^{+}-T_{0}^{\prime}).$
We will bound $|J\cap[t_{i+1},T]|$ using Proposition 1.13. We need the following:
Claim.
As $i\to\infty$, $\lVert w_{t_{i}}\rVert\to\infty$ and $t_{i}\to\infty$.
For $w\in\pi^{-1}(x)\cdot p_{H}$, the maps $s\mapsto\lVert\psi(s)w\rVert$ and $s\mapsto Q(\psi(s)w)$ are definable and for $s\geq T_{0}^{\prime}$. Let
$S_{w}:=\{s\geq T_{0}^{\prime}:\lVert\psi(s)w\rVert=R_{2}\text{ or }\lvert Q(\psi(s)w)\rvert=\epsilon_{1}\}.$
Since $S_{w}$ is a definable subset of $\mathbb{R}$, it has only finitely many connected components, see Theorem B.6. For any $i\in\mathbb{N}$, if $w=w_{t_{i}}$, then $(t_{i},t_{i}^{+})$ contains at least one connected component of $S_{w}$; indeed, since $\psi(t_{i})w_{t_{i}}\in\Psi$ and $\psi(t_{i}^{+})w_{t_{i}}\not\in\mathring{\Phi}$, by (6.9), $t_{i},t_{i}^{+}\not\in S_{w}$ and $(t_{i},t_{i}^{+})\cap S_{w}\neq\emptyset$. Therefore, since $\{t_{i}\}_{i=1}^{\infty}$ is an infinite sequence, the set $\{i\in\mathbb{N}:w=w_{t_{i}}\}$ is finite.
So, $\{w_{t_{i}}:i\in\mathbb{N}\}$ is infinite. Since $\pi^{-1}(x)\cdot p_{H}$ is discrete in $V_{H}$, we get $\lim_{i\to\infty}\|w_{t_{i}}\|=\infty$. Since $\psi(t_{i})w_{t_{i}}\in\Psi$ for all $i$ and the map $s\mapsto\lVert\psi(s)^{-1}\rVert$ is continuous on $[T_{0}^{\prime},\infty)$, we conclude that $t_{i}\to\infty$. This proves the claim.
By Proposition 1.13, and recalling (1.6), given $\epsilon>0$ we can pick $M_{0}>0$ such for any $w\in V_{H}$, if $\lVert\psi(T^{\prime}_{0})w\rVert\geq M_{0}$, then
(6.16) |$\left|\{s\in[T_{0}^{\prime},T]:\psi(s)w\in\Psi\}\right|\leq\epsilon(T-T_{0}^{\prime}).$
Due to the above claim, let $i_{x}\in\mathbb{N}$ be such that $\lVert w_{t_{i}}\rVert\geq\lVert\psi(T_{0}^{\prime})^{-1}\rVert\cdot M_{0}$ for all $i\geq i_{x}$. Set $T_{x}=t_{i_{x}}$. Let $T\geq T_{x}$. Since $t_{j}\to\infty$ as $j\to\infty$, $i:=\max\{j:t_{j}\leq T\}<\infty$. Then $i\geq i_{x}$, so $\lVert\psi(T^{\prime}_{0})w_{t_{i}}\rVert\geq M_{0}$. By (6.16),
(6.17) $\lvert J\cap[t_{i},T]\rvert\leq\lvert\{s\in[t_{i},T]:\psi(s)w_{t_{i}}\in\Psi\}\rvert\leq\epsilon(T-T_{0}^{\prime}).$
By combining (6.15) and (6.17), we obtain (6.11). ∎
6.3. Proof of Theorem 1.7
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\mathcal{G}\subseteq G$ is a connected, closed Lie subgroup with a lattice $\Gamma\subseteq\mathcal{G}$. Let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve for $G$ such that $\varphi\left([0,\infty)\right)\subseteq\mathcal{G}$. Let $H_{\varphi}$ be the hull of $\varphi$ in $G$ and let $\beta$ be a definable correcting curve in $G$, so that $\beta\varphi\subseteq H_{\varphi}$, see Definition 2.6. Then, by Proposition 5.13, $H_{\varphi}\subseteq\mathcal{G}$ and $\beta([0,\infty))\subseteq\mathcal{G}$.
We choose an injective algebraic map $\iota:G\hookrightarrow\text{SL}(n,\mathbb{R})$ for some $n\geq 2$. Note that $\iota(\Gamma)$ is a discrete subgroup of $\text{SL}(n,\mathbb{R})$. Importantly, $\iota\circ\varphi$ is non-contracting in $\text{SL}(n,\mathbb{R})$ by Lemma 2.17, and $\iota(H_{\varphi})=H_{\iota\circ\varphi}$ by Lemma 2.10. By abuse of notations, in the following we replace $\varphi$ with $\iota\circ\varphi$, replace $\Gamma$ with $\iota(\Gamma)$ and replace $G$ with $\iota(G)$.
Let $x_{0}\in\mathcal{G}/\Gamma\hookrightarrow\text{SL}(n,\mathbb{R})/\Gamma$, and fix a limiting measure $\mu$ as $T\to\infty$ of $\mu_{T,\beta\varphi,x_{0}}$ (defined in (1.1)), with respect to the weak-$\ast$ topology on the space of probability measures on $\text{SL}(n,\mathbb{R})/\Gamma$. We fix a sequence $\{T_{i}\}$ such that $T_{i}\to\infty$ and
[…]
By Theorem 4.5, $\mu$ is a probability measure.
Let $W$ be the subgroup of $\text{SL}(n,\mathbb{R})$ generated by all unipotent one-parameter subgroups under which the limiting measure $\mu$ is invariant. By Proposition 5.1, $\dim(W)>0$. Using Theorem 6.3, there exists $H\in\mathcal{H}$ such that
(6.19) $\mu(\pi(N(H,W)))>0\text{ and }\mu(\pi(S(H,W)))=0.$
Here $H\in\mathcal{H}$ is of smallest dimension such that $\mu(\pi(N(H,W)))>0$.
In view of (6.7), we define
(6.20) $N^{1}(H)$ $:=\eta_{H}^{-1}(p_{H})=\{g\in\text{SL}(n,\mathbb{R}):g\cdot p_{H}=p_{H}\}$
=$=\{g\in N(H):\det((\text{Ad}g)|_{\mathfrak{h}})=1\},$
where $p_{H}$ and $\eta_{H}$ are as in (6.3). Since $H\cap\Gamma$ is a lattice in $H$, $H$ is unimodular (see [RA72, Remark 1.9]). So, $H\subseteq N^{1}(H)$.
The remainder of the proof proceeds as follows. In Lemma 6.7, we show that there exists an element $g_{0}\in\text{SL}(n,\mathbb{R})$ such that $g_{0}\Gamma=x_{0}$ and
$\{\beta(t)\varphi(t)g_{0}\}_{t\geq 0}\subseteq g_{0}N^{1}(H).$
We first note that $\Gamma N^{1}(H)=\eta_{H}^{-1}(\Gamma\cdot p_{H})$ is closed, as $\Gamma\cdot p_{H}$ is discrete in $V_{H}$ (see Theorem 6.4, (2)). Consequently, the quotient $N^{1}(H)\Gamma/\Gamma\subseteq X$ is also closed.
By (6.18), it follows that $\mu$ is supported on $g_{0}N^{1}(H)x_{0}$. One may view $g_{0}N^{1}(H)\Gamma/\Gamma$ as foliated by orbits of $g_{0}Hg_{0}^{-1}$, and we will show that $\mu$ is invariant under $g_{0}Hg_{0}^{-1}$. The key step is, to prove that $\mu$ is actually supported on the single orbit of $g_{0}Hg_{0}^{-1}$ passing through $x_{0}$, a fact established in Corollary 6.9. This will then immediately imply our main result, Theorem 1.7, as it shows that
$\overline{H_{\varphi}x_{0}}=g_{0}Hg_{0}^{-1}x_{0},$
and that $\beta\varphi$ is equidistributed on this orbit.
Lemma 6.7.
Fix $H\in\mathcal{H}$ such that (6.19) holds. Then, there exists $g_{0}\in\text{SL}(n,\mathbb{R})$ such that $x_{0}=\pi(g_{0}):=g_{0}\Gamma$ and
(6.21) $\left\{g_{0}^{-1}\beta(t)\varphi(t)g_{0}\right\}_{t\geq 0}\subseteq N^{1}(H).$
Importantly,
(6.22) $\nu:=(g_{0}^{-1})_{*}\mu,$
is supported on $\pi(N^{1}(H))$, $\nu$ is $H$-invariant, and the largest group generated by unipotent one-parameter subgroups preserving $\nu$ is $H_{u}$.
Proof.
In view of (6.1), due to (6.19), there is a compact subset $C\subseteq N(H,W)\setminus S(H,W)$ such that $\mu(\pi(C))=\alpha$ for some $\alpha>0$, and $\pi(C)\cap\pi(S(H,W))=\varnothing$. Consider the compact subset $\text{K}_{1}\subseteq A_{H}$ given by $\text{K}_{1}:=C\cdot p_{H}$. We will now apply Theorem 6.5 to $x_{0}\in X$.
We show that outcome (2) of Theorem 6.5 cannot hold. Let $\epsilon:=\frac{\alpha}{2}$. Denote by $\mathcal{S}\subseteq\pi(S(H,W))$ the closed subset given by outcome (2) of Theorem 6.5. Let $K\subseteq X\setminus\mathcal{S}$ be an open neighborhood of $\pi(C)$, which exists as $\pi(C)$ is compact and $\mathcal{S}$ is closed. Also, fix an arbitrary compact neighborhood $U$ of $\text{K}_{1}$ in $V_{H}$. Then, $K\cap\pi(\eta_{H}^{-1}(U))$ is an open neighborhood of $\pi(C)$. By (6.18), for all $i$ large enough,
|$|\{t\in[0,T_{i}]:\beta(t)\varphi(t)x_{0}\in K\cap\pi(\eta_{H}^{-1}(U))\}|\geq\epsilon T_{i},$
which contradicts outcome (2) of Theorem 6.5.
Therefore, the first outcome (1) of Theorem 6.5 must hold. That is, we can pick $g_{0}\in\pi^{-1}(x_{0})$ such that $g_{0}\cdot p_{H}\in\text{K}_{1}\subseteq A_{H}$ and such that:
(6.23) $\beta(t)\varphi(t)g_{0}\cdot p_{H}=g_{0}\cdot p_{H},\text{ for }t\geq 0.$
Namely, $\beta(t)\varphi(t)g_{0}\in g_{0}N^{1}(H)$ for all $t\geq 0$. Since $\pi(N^{1}(H))$ is closed, we obtain that $\mu$ is supported on a subset of $\pi(g_{0}N^{1}(H))$. In fact, we claim that $\mu(\pi(g_{0}N^{1}(H)))=1$
To see this, observe that since $g_{0}\cdot p_{H}\in A_{H}$, then $g_{0}\in N(H,W)$ by Theorem 6.4,(1). Moreover, $g_{0}N^{1}(H)\subseteq N(H,W)$, since $N(H,W)$ is right-$N^{1}(H)$ invariant.
In particular, we conclude that $\mu(\pi(g_{0}N^{1}(H)\setminus S(H,W)))=1$ (see (6.19)). Therefore, by [Mozes1995OnTS, Lemma 2.4], almost every $W$-ergodic component of $\mu$ is invariant under $gHg^{-1}=g_{0}Hg_{0}^{-1}$ for some $g\in g_{0}N^{1}(H)\setminus S(H,W)$. Therefore $\mu$ is $g_{0}Hg_{0}^{-1}$-invariant. Then, $\nu=(g_{0}^{-1})_{*}\mu$ is $H$-invariant, and especially, $H_{u}$ preserves $\nu$.
Finally, we claim that $H_{u}=g_{0}^{-1}Wg_{0}\subseteq H$. To see this equality, recall that $g_{0}\in N(H,W)$, so that $g_{0}^{-1}Wg_{0}\subseteq H$, and recall that $W$ is chosen to be the subgroup of $\text{SL}(n,\mathbb{R})$ generated by all unipotent one-parameter subgroups preserving $\mu$. ∎
We will use the following notations: $\psi(t):=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$, for $t\geq 0$, $N:=N^{1}(H)$, $\Delta:=N^{1}(H)\cap\Gamma$ and $q:N\to N/H$ denotes the natural quotient map.
Consider $\bar{N}:=q(N)$, $\bar{\Delta}:=q(\Delta)$. Since $H\Delta=\Delta H$ is closed, we have that $\bar{\Delta}\leq\bar{N}$ is discrete (since $\Delta H$ is closed, then $q(\Delta)=\Delta H/H\cong\Delta/\Delta\cap H$). We let $\bar{q}:N/\Delta\to\bar{N}/\bar{\Delta}$ be the natural map. Notice that the fibers of $\bar{q}$ are $H$-orbits. Namely,
$\bar{q}^{-1}(\bar{n}\bar{\Delta})=nH\Delta/\Delta=Hn\Delta/\Delta.$
Since $\nu$ is a probability measure on $N/\Delta\cong\pi(N^{1}(H))$, it follows that $\bar{\nu}:=\bar{q}_{*}\nu$ is a probability measure on $\bar{N}/\bar{\Delta}$.
Lemma 6.8.
Suppose that $\psi(t)$ is unbounded in $N/H_{u}$. Then there exists a unipotent one-parameter subgroup $U\leq N$ such that $U$ is not contained in $H$, the measure $\bar{\nu}$ is invariant under $q(U)$, and $q(U)$ is a nontrivial one-parameter subgroup.
Moreover, $\nu$ is $U$-invartiant.
Proof.
We note that the measure push-forward map $q_{*}$ is continuous, and in particular,
[…]
Thus, we may conclude the statement by applying Proposition 5.12, and Lemma 5.5 (see the proof of Lemma 5.6. The only difference here is that $N/H$ is not a subgroup of $\text{GL}(n,\mathbb{R})$).
By the ergodic decomposition for the $H$-action on $N/\Delta$ (see e.g. [EW11, Theorem 8.20]), for all $f\in C_{c}(N/\Delta)$,
[…]
where $\mu_{H\Delta/\Delta}$ denotes the $H$-invariant probability measure on $H\Delta/\Delta$.
Therefore, since $\bar{\nu}$ is $q(U)$-invariant, $\nu$ is $U$-invariant. ∎
Corollary 6.9.
It holds that $\psi(t)\in H_{u}$ for all $t\geq 0$, and $\nu$ is the $H$-invariant probability on $H\Delta/\Delta\cong H\Gamma/\Gamma$. Moreover, $H$ is the smallest subgroup such that $g_{0}^{-1}H_{\varphi}g_{0}\subseteq H$ and that the orbit $H\Delta/\Delta$ is closed.
Proof.
We first show that $\psi(t)$ is bounded in $N/H_{u}$. Suppose $\psi(t)$ is unbounded in $N/H_{u}$. Then by Lemma 6.8, $\nu$ is invariant by the group generated by $U$ and $H$, which strictly includes $H_{u}$. This is a contradiction, because $H_{u}$ is the maximal subgroup generated by unipotent one-parameter subgroups preserving $\nu$.
Now, since the image of $\psi(t)=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$ is boundned in $\text{SL}(n,\mathbb{R})/H_{u}$, we get that the image of $\beta(t)\varphi(t)$ is bounded in $\text{SL}(n,\mathbb{R})/g_{0}H_{u}g_{0}^{-1}$. Thus, since $\text{zcl}(H_{u})$ is observable (see Lemma 2.3), by minimality of the hull (see Theorem 2.5), $H_{\varphi}\subseteq g_{0}H_{u}g_{0}^{-1}$, and in particular $\beta(t)\varphi(t)\in g_{0}H_{u}g_{0}^{-1}$ for all $t$. Namely, $\psi(t)\in g_{0}^{-1}H_{\varphi}g_{0}\subseteq H_{u}$ for all $t\geq 0$. In particular, $\bar{\nu}$ is the Dirac measure at the identity coset. Therefore $\nu=\mu_{H\Delta/\Delta}$ by (6.24). In particular $\{\psi(t)\Delta\}_{t\geq 0}$ is dense in $H\Delta/\Delta$. In particular $g_{0}^{-1}H_{\varphi}g_{0}\Delta$ is dense in $H\Delta/\Delta$, which finishes the proof.
Theorem 1.7 follows directly from the above corollary.
7. Curves with a largest possible hull
The main goal of this section is to prove Proposition 1.9 on the equidistribution of the curve (1.5). As an overview, the key idea is that the curve (1.5) will be strongly non-contracting, in the sense that it takes nonzero wedges to $\infty$, Lemma 7.2. The following result, of independent interest, implies the hull of (1.5) is semi-simple and acts irreducibly on $\mathbb{R}^{n}$. Finally, a key classification of [shah2023equidistribution, Theorem A.7] will allow to prove that the hull equals to the full group $\text{SL}(n,\mathbb{R})$.
Lemma 7.1.
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous curve definable in an o-minimal structure such that $\varphi$ has the following divergence property:
(7.1) $\lim_{t\to\infty}\varphi(t)\cdot(v_{1}\wedge\dots\wedge v_{k})=\infty,$
for all $11\leq k\leq n-$ 1 and all linearly independent vectors $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$. Then, the hull $H_{\varphi}$ of $\varphi$ is a connected semi-simple with no compact factors, and the action of $H_{\varphi}$ on $\mathbb{R}^{n}$ is irreducible.
Proof.
Since $\text{zcl}(H_{\varphi})$ is observable, we have a finite-dimensional rational representation $V$ of $\text{SL}(n,\mathbb{R})$ and $v\in V$ such that $\text{zcl}(H_{\varphi})$ is the isotropy group of $v$. Note that $\varphi(t)\cdot v,t\geq 0$ is bounded since $\varphi$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$. Since we assume (7.1), it follows by Lemma 2.1 that the orbit $\text{SL}(n,\mathbb{R})\cdot v$ is Zariski closed. Then, Matsushima criterion [Matsushima_crit] implies that $\text{zcl}(H_{\varphi})$ is reductive. Since $H_{\varphi}$ is connected and generated by one-parameter unipotent subgroups, it follows that $H_{\varphi}$ is a connected semi-simple group with no compact factors. Finally, we prove that $H_{\varphi}$ acts on $\mathbb{R}^{n}$ irreducibly. Suppose for contradiction that $\mathbb{R}^{n}=V_{1}\oplus V_{2}$ where both $V_{1}$ and $V_{2}$ are nontrivial sub-spaces invariant under $H_{\varphi}$. Let $v_{1},\dots,v_{k}$ be a basis for $V_{1}$. Here $1\leq k<n$ because $V_{2}$ is not trivial. Since semi-simple groups have no nontrivial characters, we obtain that $H_{\varphi}\cdot(v_{1}\wedge\cdots\wedge v_{k})=v_{1}\wedge\cdots\wedge v_{k}$. Since $\varphi([0,\infty))$ is bounded in $\text{SL}(n,\mathbb{R})/H_{\varphi}$, we get $\varphi(t)\cdot(v_{1}\wedge\cdots\wedge v_{k})$ is bounded as $t\to\infty$. This contradicts (7.1). ∎
Let $\varphi$ be the curve given by (1.5). We now verify that $\varphi$ satisfies a divergence property, which is stronger than (7.1).
Lemma 7.2.
For any $1\leq k\leq n$, and $0\neq v\in\bigwedge^{k}\mathbb{R}^{n+1}$, $\lim_{t\to\infty}\varphi(t)\cdot v\to\infty$ as $t\to\infty$.
Proof.
We denote by $e_{0},e_{1},\dots,e_{n}$ the canonical basis vectors of $\mathbb{R}^{n+1}$, where $e_{i}$ denotes the vectors whose coordinates equal 00 besides the $(i+1)$th position in which $1$ appears. Fix an integer $1\leq k\leq n$ and let $\mathcal{I}$ be the collection of all subsets of $\{0,\ldots,n\}$ of cardinality $k$. Let $\varphi$ be the curve given by (1.5). For any $I:=\{i_{1},\ldots,i_{k}\}\in\mathcal{I}$, where $i_{1}<i_{2}<\cdots<i_{k}$, let $e_{I}=e_{i_{1}}\wedge\cdots\wedge e_{i_{k}}$. There are two cases for $\varphi(t)\cdot e_{I}$:
If $i_{1}=0$, then
$\varphi(t)\cdot e_{I}=$ $f_{0}(t)e_{0}\wedge[f_{i_{2}}(t)e_{0}+h_{i_{2}}(t)^{-1}e_{i_{2}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
(7.2) $=$ $f_{0}(t)\left(\prod_{l=2}^{k}h_{i_{l}}(t)^{-1}\right)e_{I};$
and if $i_{1}\geq 1$, then
$\varphi(t)\cdot e_{I}=$ $[f_{i_{1}}(t)e_{0}+h_{i_{1}}(t)^{-1}e_{i_{1}}]\wedge\cdots\wedge[f_{i_{k}}(t)e_{0}+h_{i_{k}}(t)^{-1}e_{i_{k}}]$
$=$ $\sum_{l=1}^{k}(-1)^{l-1}f_{i_{l}}(t)\left(\prod_{j\in\{1,\ldots,k\}\setminus\{l\}}h_{i_{j}}(t)^{-1}\right)e_{\{0\}\cup I\setminus\{i_{l}\}}$
(7.3) $+\prod_{j\in\{1,\ldots,k\}}h_{i_{j}}(t)^{-1}e_{I}.$
Let $\mathbf{v}\in\bigwedge^{k}\mathbb{R}^{n+1}\setminus\{0\}$. Then $\mathbf{v}=\sum_{I\in\mathcal{I}}c_{I}e_{I}$, where $c_{I}\in\mathbb{R}$. Then,
(7.4) $\varphi(t)\cdot\mathbf{v}=\sum_{I\in\mathcal{I}}C_{I}(t)e_{I},$
where $C_{I}(t)\in\mathbb{R}$. The coefficients of our interest are $C_{I}(t)$ such that $0\in I$. For any $1\leq i_{2}<\cdots<i_{k}\leq n$, and $J=\{i_{2},\ldots,i_{k}\}$,
(7.5) $C_{\{0\}\cup J}(t)=\left(\prod_{j\in J}h_{j}(t)^{-1}\right)\sum_{l\in\{0,1,\ldots,k\}\setminus J}f_{l}(t)c_{J\cup\{l\}}(-1)^{j_{l}},$
where $j_{l}=\min\{j\in\{2,\ldots,k\}:i_{j}>l\}$.
Let $\mathcal{I}_{\mathbf{v}}=\{I\in\mathcal{I}:c_{I}\neq 0\}\neq\varnothing$, as $\mathbf{v}\neq 0$. Let $i_{1}=\min\cup\mathcal{I}_{\mathbf{v}}$. We pick $I\subseteq\mathcal{I}_{\mathbf{v}}$ such that $i_{i}\in I$. Let $J=I\setminus\{i_{1}\}$. Then, $c_{J\cup\{l\}}=0$ for every $l<i_{1}$. Therefore, by (7.5)
(7.6) $C_{\{0\}\cup J}(t)=\left(\prod_{j\in J}h_{j}(t)\right)^{-1}\left(f_{i_{1}}c_{I}+\sum_{l\in\{i_{1}+1,\ldots,k\}\setminus J}f_{l}(t)c_{J\cup\{l\}}(-1)^{j_{l}}\right).$
By condition (1), $\sum_{i=1}^{n}\deg h_{i}=n$ and $\deg h_{1}\geq\cdots\geq\deg h_{n}$. Therefore
(7.7) $\sum_{i=1}^{j}\deg h_{i}\geq j$
for each $1\leq j\leq n$. Since $1\leq k\leq n$ and $\deg h_{n}>0$, we get
(7.8) $\deg h_{i_{2}}+\cdots+\deg h_{i_{k}}\leq\deg h_{1}+\cdots+\deg h_{k-1}\leq n-\deg h_{n}<n,$
and if $i_{1}\geq 1$, then by (7.7),
(7.9) $\deg h_{i_{2}}+\cdots+\deg h_{i_{k}}\leq n-(\deg h_{1}+\cdots+\deg h_{i_{1}})\leq n-i_{1}.$
Since $c_{I}\neq 0$, by our assumptions (2) and (3),
(7.10) $\deg\left(c_{I}f_{i_{1}}+\sum_{l\in\{i_{1}+1,\ldots,k\}\setminus J}(-1)^{j_{l}}c_{J\cup\{l\}}f_{l}(t)\right)\begin{array}[]{ll}\geq n&\text{if $i_{1}=0$}\\ >n-{i_{1}}&\text{if $i_{1}\geq 1$}.\end{array}$
By combining (7.6), (7.8), (7.9) and (7.10), we conclude that $\deg C_{\{0\}\cup J}(t)>0$. Hence, the conclusion of the lemma follows from (7.4). ∎
Lemma 7.3.
Let $\{\rho(s):s\in\mathbb{R}\}$ be the P.S. group of $\varphi$ (Definition 5.3). Then, there exists $v=(v_{1},\ldots,v_{n})\in\mathbb{R}^{n}\setminus\{0\}$, such that
(7.11) $\rho(s):=\begin{bmatrix}1&sv\\ 0&I_{n}\end{bmatrix}\in\text{SL}(n+1,\mathbb{R}),\,\forall s\in\mathbb{R}.$
Proof.
Due to Lemma 7.2, by Lemma 5.9, $\rho(s)$ is unipotent. Moreover, there exists $r<1$ such that for any $s\in\mathbb{R}$,
(7.12) $\varphi(h_{r,s}(t))\varphi(t)^{-1}\to\rho(s),$
where $h_{r,s}$ is as in (5.1). We note that $\varphi(s)$ is contained in the subgroup $F$ of $\text{SL}(n+1,\mathbb{R})$ whose nonzero entries are only on the diagonal and in the top row. Therefore, by (7.12), $\rho(s)$ is a one-parameter unipotent subgroup of $F$. Hence, $\rho(s)$ is of the form given by (7.11). ∎
Proof of Proposition 1.9.
Let $\{e_{0},\ldots,e_{n}\}$ denote the standard basis of $\mathbb{R}^{n+1}$. Let $(v_{1},\ldots,v_{n})\in\mathbb{R}^{n}$ be as in Lemma 7.3. Let $W_{1}$ denote the orthogonal complement of $v=\sum_{i=1}^{n}v_{i}e_{i}$ in the span of $\{e_{1},\ldots,e_{n}\}$. Then the fixed points space of $\rho(s)$ in $\mathbb{R}^{n+1}$ is $W:=\mathbb{R}e_{0}+W_{1}$.
Now let $\beta$ denote the correcting definable curve such that $\beta(t)\varphi(t)\subseteq H_{\varphi}$. Since $\beta(t)\to\beta(\infty)\in\text{SL}(n+1,\mathbb{R})$ as $t\to\infty$, by (7.12), we conclude that $\rho_{1}(s):=\beta(\infty)\rho(s)\beta(\infty)^{-1}\in H_{\varphi}$ for all $s\in\mathbb{R}$. By Lemma 7.2 and Lemma 7.1, $H_{\varphi}$ is a semisimple group acting irreducibly on $\mathbb{R}^{n+1}$. Therefore, by Jacobson-Morosov theorem, there exists a homomorphism $\eta:\text{SL}(2,\mathbb{R})\to H_{\varphi}$ such that $\eta(u(s))=\rho_{1}(s)$, where $u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}\in\text{SL}(2,\mathbb{R})$. As noted earlier, the fixed points space of $\rho_{1}$ on $\mathbb{R}^{n+1}$ is $\beta(\infty)W$, and it has dimension $n$. Expressing the representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{n+1}$ via $\eta$ as a direct sum of irreducible representations of $\text{SL}_{2}(\mathbb{R})$, and noting that the dimension of the $u(s)$-fixed subspace is $n$, we conclude that $\mathbb{R}^{n+1}=V_{1}\oplus V_{2}$, where $V_{1}$ is isomorphic to the standard representation of $\text{SL}(2,\mathbb{R})$ on $\mathbb{R}^{2}$, and $\text{SL}(2,\mathbb{R})$ acts trivially on $V_{2}$. Therefore, by [shah2023equidistribution, Theorem A.7], either $H_{\varphi}=\text{SL}(n+1,\mathbb{R})$ or $n+1=2d$ for some $d\geq 2$ and $H_{\varphi}$ is conjugate to the standard symplectic group $\text{SP}(\mathbb{R}^{2d})$.
Consider the case when $H_{\varphi}\neq\text{SL}(n+1,\mathbb{R})$. Then, there exists a basis $w_{1},w^{\prime}_{1},w_{2},w^{\prime}_{2},\ldots,w_{d},w^{\prime}_{d}$ of $\mathbb{R}^{2d}$ such that $H_{\varphi}$ stabilizes
$\mathbf{v}=w_{1}\wedge w^{\prime}_{1}+w_{2}\wedge w^{\prime}_{2}+\cdots+w_{n}\wedge w^{\prime}_{n}\in\wedge^{2}\mathbb{R}^{2d}.$
Now, by Lemma 7.2, $\varphi(t)$ cannot be bounded in $\text{SL}(2d,\mathbb{R})/H_{\varphi}$, contradictions the assumption that $H_{\varphi}$ is the hull of $\varphi$. Therefore, $H_{\varphi}=\text{SL}(n+1,\mathbb{R})$. This completes the proof of Proposition 1.9. ∎
Proof of Corollary 1.10.
If $G$ is a linear Lie group, the result follows immediately from Proposition 1.9 and Theorem 1.7. If $G$ is not a linear Lie group, for our proof to work, it is sufficient that after the composition with the Adjoint representation of $G$ on its Lie algebra $\mathfrak{g}$, the curve is definable and non-contracting in $\text{GL}(\mathfrak{g})$. This property is valid in our situation. ∎
Appendix A Kempf’s Lemma for real algebraic groups
We note that [Kem78, Lemma 1.1(b)] is stated for reductive algebraic groups over an arbitrary field. We need the same result for any algebraic group defined over $\mathbb{R}$. Here, we verify that the proof given by Kempf is valid for real algebraic groups $G$ using the most naive language.
For finite-dimensional real vector spaces $V$ and $W$, we say that $P:V\to W$ is a polynomial map if, after choosing bases for $V$ and $W$, the map $P$ in those coordinates is a multi-variable polynomial in each coordinate. We say that the map is $G$-equvairant if $g\cdot P(x)=P(g\cdot x)$ for all $g\in G$ and $x\in V$.
Lemma A.1 (Kempf).
Let $G$ be the set of $\mathbb{R}$-points of an algebraic group defined over $\mathbb{R}$. Let $V$ be a rational representation of $G$. Suppose that for $v\in V$, it holds that $S:=\text{zcl}(G\cdot v)\setminus G\cdot v$ is non-empty. Then there exists a rational representation $W$ and an equivariant polynomial map $P:V\to W$ such that $P(S)=\{0\}$ and $P(v)\neq 0$.
Proof.
First, note that since $\text{zcl}(G\cdot v)$ and $G\cdot v$ are $G$-invariant, we get that $S$ is $G$-invariant. Let $\mathbb{R}[V]$ be the coordinate ring on the vector space $V$ and consider $I\subseteq\mathbb{R}[V]$ the polynomial ideal of the polynomials vanishing on $S$. For $d\in\mathbb{N}$, let $I(d)\subseteq I$ be the subset of polynomials in $I$ with degree bounded by $d$. We may choose $d$ such that the ideal generated by $I(d)$ is $I$. Notice that $I(d)$ is a finite-dimensional vector space. Consider the right action of $G$ on $\mathbb{R}[V]$ defined by
(A.1) $(f.g)(x):=f(g\cdot x),~{}g\in G,~{}x\in V,~{}f\in\mathbb{R}[V].$
The action preserves the degree of a polynomial. So, it preserves $I(d)$ because $S$ is $G$-invariant. Let $I(d)^{*}$ be the space of linear functionals on $I(d)$. Using the representation (A.1), we get a left action on $I(d)^{*}$ defined by
(A.2) $g\cdot l(f):=l(f\cdot g),~{}g\in G,~{}l\in I(d)^{*},~{}f\in I(d).$
Now consider the map $P:V\to I(d)^{*}$ defined by
(A.3) $P(x):=\text{ev}_{x},$
where $\text{ev}_{x}$ is the functional mapping $f\in I(d)$ to $f(x)$. It is straightforward to verify that $P$ is an equivariant polynomial map. Finally, we claim that $P(v)=\text{ev}_{v}$ is a nonzero functional. By the definition of $S$ and $P$, $P(x)=0$ for all $x\in S$. It remains to show that $P(v)\neq 0$. For that, recall the closed orbit lemma [Bo91, Section 1.8], which states that $S$ is Zariski closed. This implies that there exists a $p\in I(d)$ such that $p(S)=0$ and $p(v)\neq 0$. Since $P(v)(p)=p(v)$, $P(v)\neq 0$. ∎
Appendix B Basic notions in the theory of o-minimal structures
In order to keep our paper self contained, we will now recall the basic definitions and results on o-minimal structures on the field of real numbers. We refer to [van_den_dries_omin_book] and [Geom_cat_and_o_min_str] for more details.
We begin with the notion of a structure on the ordered field of real numbers, and later introduce the stronger notion of an o-minimal structure.
Definition B.1.
A structure $\mathscr{S}$ on the ordered field of real numbers is a sequence of families of sets
$\mathscr{S}_{d}\subseteq\{\text{subsets of }\mathbb{R}^{d}\},~{}d\in\mathbb{N},$
such that the following requirements are satisfied:
  1. (1)
    For each $n\in\mathbb{N}$, $\mathscr{S}_{n}$ is a Boolean algebra of sets, and $\mathbb{R}^{n}\in\mathscr{S}_{n}$.
  2. (2)
    For each $n\in\mathbb{N}$, $\mathscr{S}_{n}$ contains the diagonals $\{(x_{1},\dots,x_{n}):x_{i}=x_{j}\},$ for all $1\leq i<j\leq n$.
  3. (3)
    For each $n\in\mathbb{N}$, if $A\in\mathscr{S}_{n}$, then $A\times\mathbb{R}\in\mathscr{S}_{n+1}$ and ×$\mathbb{R}\times A\in\mathscr{S}_{n+1}$.
  4. (4)
    For each $n\in\mathbb{N}$, if $A\in\mathscr{S}_{n+1}$, then $\pi(A)\in\mathscr{S}_{n}$, where $\pi:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ is the projection to the first $n$ coordinates.
  5. (5)
    $\mathscr{S}_{3}$ contains the graph of addition $\{(x,y,x+y):x,y\in\mathbb{R}\}$ and the graph of multiplication $\{(x,y,xy):x,y\in\mathbb{R}\}$.
Given a structure $\mathscr{S}$, we say that $A\subseteq\mathbb{R}^{d}$ is definable if $A$ belongs to $\mathscr{S}_{d}$. We say that a function $f:B\to\mathbb{R}^{m}$ is definable, for $B\subseteq\mathbb{R}^{d}$, if its graph $\{(x,f(x):x\in B\}$ belongs to $\mathscr{S}_{d+m}$.
We note the following basic properties, which are proved directly from Definition B.1
Lemma B.2.
Let $\mathscr{S}$ be a structure on the field of real numbers. Then:
  1. (1)
    The restriction of a definable function to a definable set is definable function on that set.
  2. (2)
    If $f:\mathbb{R}^{d}\to\mathbb{R}$ is definable, and $A\subseteq\mathbb{R}^{d}$, is definable, then the subsets:
    $\left\{\mathbf{x}\in A\mid f(\mathbf{x})=0\right\},\ \left\{\mathbf{x}\in A\mid f(\mathbf{x})>0\right\}$
    are definable.
  3. (3)
    If $f:A\to B$ and $g:B\to C$ are definable, then the composition $g\circ f$ is also definable.
  4. (4)
    If $f,g:A\to\mathbb{R}$ are definable, then $f+g,$ $f-g,$ $f\cdot g$ are definable, and $f/g$ on the domain $\{x\in A:g(x)\neq 0\}$ is definable.
The notion of a structures belongs to model theory, and it is most natural to describe definable sets using first order logic formulae, see [Geom_cat_and_o_min_str, Appendix A] for a concise introduction. The advantage of the definition above is in allowing to define structures without resorting to notions of model theory.
The proof of the statement below demonstrates the technique of proving the definability of functions using first-order logic formulae. The claim is useful for us since the type of the function considered there is used in the proof of Theorem 3.2.
Lemma B.3.
Suppose that $f:\mathbb{R}^{2}\to\mathbb{R}$ is a bounded definable function, and let $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ be definable functions such that $a(s)<b(s),~{}\forall s$. Then the function $\Psi:\mathbb{R}\to\mathbb{R}$ defined by
$\Psi(s):=\sup\left\{|f(s,t)|:t\in\left[a(s),b(s)\right]\right\},\ \forall s\in\mathbb{R},$
is definable.
Proof.
To prove the $\Psi$ is definable, we need to verify that the graph of $\Psi$, $\Gamma(\Psi):=\{(s,\Psi(s)):s\in\mathbb{R}\}$ is definable. Consider the definable set:
$A_{1}:=\{(s,y,t):a(s)\leq t\leq b(s)\text{ and }|f(s,t)|>y\},$
and consider the projection $\pi_{3,2}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ to the first coordinates. Then $B_{1}:=(\pi_{2,3}(A_{1}))^{c}$ is definable, and it can be alternatively described by:
$B_{1}:=\{(s,y):(\forall t\in[a(s),b(s)])\ |f(s,t)|\leq y\}.$
In a similar way (while slightly complicated), the set:
$B_{2}:=\{(s,y):(\forall\epsilon>0)(\exists t\in[a(s),b(s)])\ y-\epsilon\leq|f(s,t)|\}$
is definable. The graph of $\Psi$ is then $B_{1}\cap B_{2}$. ∎
B.1. O-minimal structures
The notion of o-minimal structures was introduced in [generalization_tarski_seind_dries], as a generalization of the real semi-algebraic geometry.
Definition B.4.
(o-minimal structures) A structure $\mathscr{S}$ expanding the real field is called o-minimal if $\mathscr{S}_{1}$ consists of all possible sets which are finite unions of points and intervals.
The semi-algebraic sets are definable in every o-minimal structure. Here, the semi-algebraic sets are boolean combinations of sets of the form $\{x\in\mathbb{R}^{d}:P(x)=0\}$, $\{x\in\mathbb{R}^{d}:P(x)>0\}$, where $P\in\mathbb{R}[X_{1},\dots,X_{d}]$. By the Tarski-Seidenberg theorem, the semi-algebraic sets form an o-minimal structure.
We collect the following important results, see [van_den_dries_omin_book] for detailed proofs. For the following, fix an o-minimal structure.
Theorem B.5.
[van_den_dries_omin_book, Ch. 3, Thm. 1.2 and Ch. 7, Prop. 2.5] Suppose that $f:(a,b)\to\mathbb{R}$ is definable, where $-\infty\leq a<b\leq\infty$. Then, there exist $a_{0},a_{1},...,a_{k+1}$ such that $a=:a_{0}<a_{1}<\dots<a_{k}<a_{k+1}:=b$ and such that in each sub-interval $I=(a_{i},a_{i+1})$, $f|_{I}$ is differentiable and $f|_{I}$ is either strictly monotonic, or constant.
A direct consequence of the monotonicity theorem is that a definable function $f:\mathbb{R}\to\mathbb{R}$ either converges as $t\to\infty$ or diverges to $\infty$ or to $\infty$- ∞.
Theorem B.6.
[van_den_dries_omin_book, Ch.3, Cor. 3.7]. If $A\subseteq\mathbb{R}^{m}$ is definable, than $A$ has finitely many connected components. Moreover, the following uniformity holds. For $A\subseteq\mathbb{R}^{d}\times\mathbb{R}^{n}$ and $x_{0}\in\mathbb{R}^{d}$, consider
$A_{x_{0}}:=\{(x_{0},y):(x_{0},y)\in A\}.$
Then, for a definable $A$, there exists an $N\in\mathbb{N}$ such for all $x\in\mathbb{R}^{m}$, the set $A_{x}$ has at most $N$ connected components.
Theorem B.7.
[van_den_dries_omin_book, Ch. 3, Prop. 1.2]. Suppose that $A\subseteq\mathbb{R}^{m+n}$ is definable. Let $\pi_{m+n,m}:\mathbb{R}^{m+n}\to\mathbb{R}^{m}$ be the projection to the first $m$ coordintates. Then, there is a definable function $f:\pi_{m+n,m}(A)\to\mathbb{R}^{n}$ such that the graph of $f$ is contained in $A$.
[…]
[title=Bibliography]